### v1

# Natural solution of SUSY $\mu$ problem from modulus stabilization in modular flavor model

Abstract
We propose to solve the SUSY $\mu$-problem in the framework of modular flavor symmetry. The bare $\mu$-term is forbidden by modular symmetry and the effective $\mu$-term is re-generated after the modulus field is stabilized. We discuss the stabilization mechanism of a single modulus field with the presence of SUSY breaking contributions described in the non-linear realized SUSY scheme with nilpotent goldstino $X_{nl}$ superfield. The presence of $X_{nl}$ contributions can possibly slightly change the original stabilized modulus VEV to a value located in the vicinity of some fixed points. Natural value of $\mu_{eff}$, which is much smaller than the SUSY scale, can either be the consequence of expansion of typical modular forms in terms of the tiny deviation parameter (away from the fixed point $\omega$), or be the joint consequence of suppression from both powers of $q$ [or $(2\Im\tau)^{-1}$] and the asymptotic suppression behavior of typical modular forms away from the fixed point $i\infty$.
1 Introduction
Low energy supersymmetry (SUSY) is well motivated theoretically and regarded as one of the most appealing extensions of the Standard Model. Not only the quadratic divergences related to the Higgs mass can be eliminated, but also successful gauge coupling unification can be realized in the framework of low energy SUSY. Especially, the observed 125 GeV scalar, discovered by both the ATLAS ATLAS and CMS CMS collaborations of LHC, falls miraculously within the narrow $115\sim 135$ GeV ’window’ required by minimal supersymmetric standard model (MSSM) and it can be seen as a triumph of low energy SUSY.
Although the framework of low energy SUSY is attracting, it still has several unsatisfactory features. In addition to the problem of null observation of conclusive signals for superparticles at the LHC, the MSSM is also bothered by the SUSY $\mu$-problem: the $\mu$-parameter is a SUSY-respecting parameter appearing in the superpotential, while it is required to be roughly of order the soft SUSY breaking scale to allow a Higgs VEV of order 174 GeV without too much miraculous cancelation.
There are many proposals to tackle the $\mu$-problem, for example, by singlet extension to next-to-minimal supersymmetric standard model (NMSSM) or by non-renormalizable Lagrangian terms in the superpotential via Kim-Nilles mechanism KN (or in the Kahler potential via Giudice-Masiero mechanism GM ). Most solutions consist of two steps Bae:2019dgg : First, the appearance of $\mu$-parameter is forbidden, usually by some discrete or continuous symmetry; Second, the $\mu$ term is re-generated via some mechanism such as symmetry breaking. One can see in Bae:2019dgg some detailed realizations of the two steps for the $\mu$-problem.
We would like to tackle the $\mu$-problem with a similar but different approach. The symmetry adopted to forbid the $\mu$-parameter is a modular flavor symmetry. Besides, the re-generation of $\mu$ term is realized after modular symmetry breaking with the modulus stabilization mechanism accompanied by typical SUSY breaking contributions. Naturally suppressed $\mu$-parameter can thus be realized.
Modular symmetry is a geometrical symmetry of torus $T^{2}$ and it corresponds to a change of the corresponding cycle basis. Modular symmetry can naturally arise in string theory, for example, (exact) moduli spaces of Calabi-Yau threefolds, like the quintic and fairly generically in large classes of elliptic Calabi-Yau threefolds. Recently, modular flavor symmetry was proposed in Feruglio:2017spp to explain the flavor puzzles of the standard model (plus the massive neutrinos), which no longer necessitates the introduction of the flavon fields that may cause difficulties in flavor model buildings. For reviews, see Feruglio:2019ybq ; Kobayashi:2023zzc ; Ding:2023htn ; Ding:2024ozt . The modular symmetry can constrain stringently the superpotential as the relevant Yukawa couplings are modular forms. The matter superfields and the modular forms in the superpotential transform in some representations of certain finite modular group $\Gamma_{N}\equiv\bar{\Gamma}/\bar{\Gamma}(N),N\leq 5$ (or its double cover $\Gamma_{N}^{\prime}$). In modular flavor symmetry framework, the masses and mixings for the quarks and leptons can be predicted with very few input parameters. Modular flavor models based on $\Gamma_{2}\simeq S_{3}$ Kobayashi:2018vbk ; Okada:2019xqk ; Du:2020ylx , $\Gamma_{3}\simeq A_{4}$ Kobayashi:2018vbk ; Petcov:2018snn ; Criado:2018thu ; Kobayashi:2018scp ; Okada:2018yrn ; Novichkov:2018yse ; Okada:2020ukr ; Petcov:2022fjf ; Nomura:2024ctl ; Pathak:2024sei , $\Gamma_{4}\simeq S_{4}$ Penedo:2018nmg ; Petcov:2018snn ; Novichkov:2018ovf ; Kobayashi:2019xvz ; Wang:2020dbp ; Qu:2021jdy , $\Gamma_{5}\simeq A_{5}$ Petcov:2018snn ; Novichkov:2018nkm ; Ding:2019xna ; Yao:2020zml ; deMedeirosVarzielas:2022ihu ; Abbas:2024bbv and non-holomorphic modular flavor models Qu:2024rns ; Nomura:2024atp ; Ding:2024inn ; Li:2024svh had been adopted in various studies to explain the low energy flavor structures of the Standard Model. Besides, modular flavor symmetries can combine with the GUT framework to increase greatly the prediction ability of the models, such as the combination of $\Gamma_{2},\Gamma_{3}$ with $SU(5)$ deAnda:2018ecu ; Kobayashi:2019rzp ; Du:2020ylx ; Chen:2021zty ; Zhao:2021jxg ; King:2021fhl ; Ding:2021zbg , SO(10) Ding:2021eva ; Ding:2022bzs or flipped SU(5) Charalampous:2021gmf ; Du:2022lij ; King:2024gsd GUT models.
The vacuum expectation value (VEV) of the modulus field can be the only source of flavour symmetry breaking, it can also be the only source of CP symmetry breaking. The VEV of the modulus field $\tau$ will in general break the modular symmetry completely except three inequivalent fixed points $\tau=i\infty,i,e^{i2\pi/3}$ in the fundamental domain. The modular forms, thus the Yukawa coupling, and the fermion mass matrices will be fixed after the modular flavour symmetry is fully or partially broken by VEV of $\tau$. From a top-down perspective, the value of modulus VEV can be determined dynamically with modular-invariant scalar potential in 4d $N=1$ supergravity, which is also crucial to give nonzero masses for modulus field to render it phenomenological acceptable.
Experiences from heterotic string theory on moduli stabilization target:space ; Leedom:2022zdm ; Gonzalo:2018guu indicate that the scalar potential favors minima at the boundary of the fundamental domain or at points inside the fundamental domain that close to the fixed point $\omega=e^{i2\pi/3}$ with negative vacuum energy. In the presence of additional ”matter”-like fields, it is possible to obtain dS vacua with additional uplifting contributions, such as the introduction of an anti-D3 brane that provides a positive uplifting potential, which would induces a SUSY breaking vacuum shift to a tiny positive vacuum energy. The presence of the anti-brane like sources can be described in terms of a nilpotent goldstino superfield KS in a supersymmetric effective field. Although SUSY is non-linearly realized, the couplings of chiral, gauge and gravity superfields coupled to the constrained nilpotent goldstino superfield $X$ can be used to describe spontaneous SUSY breaking in supergravity Burgess:2021juk ; Aparicio:2015psl ; NL:SUGRA1 ; NL:SUGRA2 ; NL:SUGRA3 ; NL:SUGRA4 ; NL:SUGRA5 ; NL:SUGRA6 and the presence of the anti-brane type sources with a supersymmetric action.
This paper is organized as follows. In Sec 2, we briefly review the $\Gamma_{N}$ modular flavor symmetry. In Sec 3, we discuss the solution of SUSY $\mu$-problem with modular symmetry. In Sec 4, we discuss the stabilization of singlet modulus field, with the presence of SUSY breaking contributions described in the non-linear realized SUSY scheme with nilpotent goldstino $X_{nl}$ superfield. In Sec 5, we discuss the re-generation of naturally small $\mu_{eff}$ after modulus stabilization. Sec 6 contains our conclusions.
2 Modular $\Gamma_{N}$ symmetry
The full modular group is the group of linear fraction transformations acting on the complex modulus $\tau$ in the upper half complex plane
$\gamma=\left(\begin{array}[]{cc}a&b\\ c&d\end{array}\right):\tau\longmapsto\gamma\tau=\frac{a\tau+b}{c\tau+d},~{}~{}\text{with}~{}~{}a,b,c,d\in\mathbb{Z},~{}~{}ad-bc=1,~{}~{}~{}{\rm Im}\tau>0\,.$ (1)
It can be generated by two generators $S$ and $T$
$S:\tau\mapsto-\frac{1}{\tau},~{}~{}~{}~{}\quad T:\tau\mapsto\tau+1\,,$ (2)
which satisfy
$S^{4}=1,~{}~{}(ST)^{3}=1,~{}~{}S^{2}T=TS^{2}.$ (3)
The two generators can be represented by the following $SL(2,Z)$ matrices,
$S=\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right),~{}~{}~{}\quad T=\left(\begin{array}[]{cc}1&1\\ 0&1\end{array}\right)\,.$ (4)
Finite modular groups $\Gamma_{N}$ Feruglio:2017spp are defined by $\Gamma_{N}=\overline{\Gamma}/\overline{\Gamma}(N)$, with $\overline{\Gamma}\equiv SL(2,Z)/\{\pm 1\}$ and $\overline{\Gamma}(N)$ defined by
$\overline{\Gamma}(N)=\left\{\begin{array}[]{c}\Gamma(N),~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}N\geq 3\\ \Gamma(N)/\{\pm 1\},~{}~{}~{}N=2\end{array}\right.$ (7)
with $\Gamma(N)$ the principal congruence subgroup of level $N$
$\Gamma(N)=\left\{\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in SL(2,\mathbb{Z}),~{}b=c=0\,(\mathrm{mod~{}N}),a=d=1\,(\mathrm{mod~{}N})\right\}\,.$ (8)
Finite modular groups $\Gamma_{N}$ can be obtained by further imposing the condition $T^{N}=1$ in addition to $S^{2}=(ST)^{3}=1$.
Modular form $f(\tau)$ of weight $2k$ and level $N$ is a holomorphic function of the complex modulus $\tau$ with well-defined transformation properties under the action of ${\Gamma}(N)$
$f\left(\frac{a\tau+b}{c\tau+d}\right)=(c\tau+d)^{2k}f(\tau)~{},~{}~{}~{}\forall~{}\gamma\equiv\left(\begin{array}[]{cc}a&b\\ c&d\end{array}\right)\in{\Gamma}(N)\,.$ (9)
Modular forms of weight $2k$ and level $N$ span a linear space of finite dimension. They can be arranged into some modular multiplets $Y_{\mathbf{r}}^{2k}$ that transform as irreducible representation $\mathbf{r}$ of the finite modular group $\Gamma_{N}$ up to automorphy factor Feruglio:2017spp , i.e.
$Y_{\mathbf{r}}^{(2k)}(\gamma\tau)=(c\tau+d)^{2k}\rho_{\mathbf{r}}(\gamma)Y_{\mathbf{r}}^{(2k)},~{}~{}~{}~{}\forall~{}\gamma\in\overline{\Gamma}\,,$ (10)
where $\gamma$ is the representative element of the coset $\gamma\overline{\Gamma}(N)$ in $\Gamma_{N}$, and $\rho_{\mathbf{r}}(\gamma)$ is the representation matrix of the element $\gamma$ in the irreducible representation $\mathbf{r}$. The actions of modular group on supermultiplets read
$\phi_{i}(\gamma\tau)=(c\tau+d)^{-k_{\phi}}\rho^{\mathbf{r^{\prime}}}_{ij}(\gamma)\phi_{j}~{},~{}~{}~{}\forall\gamma\in\bar{\Gamma}~{},$ (11)
where $k_{\phi}$ denotes the modular weight of $\phi$, and $\rho$ is a unitary representation of the quotient group $\bar{\Gamma}_{N}$.
3 Natural $\mu$-term in modular flavor model
As noted in Bae:2019dgg , to solve the SUSY $\mu$ problem, the appearance of $\mu$-parameter in the superpotential should be forbidden by some symmetry. We propose that the superfields $H_{u}$, $H_{d}$ transform non-trivially under modular transformation so as that ordinary $\mu H_{u}H_{d}$ term is forbidden by modular symmetry.
The relevant Lagrangian can be given as
${\cal L}$ $\supseteq$ $\int d^{4}\theta\left[-3\log(2\Im\tau)+\frac{H_{u}^{\dagger}H_{u}}{(2\Im\tau)^{k_{u}}}+\frac{H_{d}^{\dagger}H_{d}}{(2\Im\tau)^{k_{d}}}+\cdots\right]$ (12)
$+$ $\int d^{2}\theta\left[\alpha\Lambda_{SUSY}\frac{Y_{N}^{k_{r}}(\tau)}{\eta(\tau)^{2k_{3}}}H_{u}H_{d}+\cdots\right],$
with the coefficient for bilinear $H_{u}H_{d}$ term given by
$\mu=\alpha\Lambda_{SUSY}\frac{Y_{N}^{(k_{r})}(\tau)}{\eta(\tau)^{2k_{3}}}~{},$ (13)
for $k_{H_{u}}+k_{H_{d}}=k_{r}-k_{3}+3$ with $62k_{3}\in 24\mathbb{Z}+$ 6 to guarantee the modular invariance of
$e^{G}\equiv e^{K}|W|^{2}~{}.$ (14)
Here, $\alpha\sim{\cal O}(1)$ is a dimensionless parameter and $\Lambda_{SUSY}$ is typically of order the SUSY scale. The Dedekind $\eta$-function is a modular form with modular weight $1/2$ under modular transformation up to a phase, whose transformation under $\gamma$ can be given as
$\eta(\gamma\tau)=\epsilon_{\eta}(\gamma)(c\tau+d)^{1/2}\eta(\tau)~{},~{}~{}\forall\gamma\in\overline{\Gamma}~{}.$ (15)
with $\epsilon_{\eta}(\gamma)$ depending on $\gamma$ (but not on $\tau$) and being given in terms of the Rademacher $\Phi(\gamma)$ function
$\epsilon_{\eta}(\gamma)=\exp\left[\frac{2\pi i}{24}\Phi(\gamma)-i\frac{\pi}{4}\right].$ (16)
that satisfies $\epsilon_{\eta}(\gamma)^{24}=1$ DHoker:2022dxx .
The expansion of $\eta(\tau)$ function is
$\eta(\tau)$ $=$ $q^{1/24}\sum\limits_{n=-\infty}^{\infty}(-1)^{n}q^{n(3n-1)/2},$ (17)
$=$ $q^{1/24}\left[1+\sum\limits_{n=1}^{\infty}(-1)^{n}\left(q^{n(3n-1)/2}+q^{n(3n+1)/2}\right)\right],$
$=$ $q^{1/24}\left[1-q-q^{2}+q^{5}+q^{7}-q^{12}-\cdots\right],~{}~{}~{}~{}~{}q\equiv e^{i2\pi\tau}~{}.$
It is obvious that the Kahler potential for superfields with non-trivial modular weights, for example, the $H_{u}$ and $H_{d}$ superfields, are non-minimal. Therefore, after the moduli fields $\tau$ acquire a VEV by modulus stabilization mechanism, the $H_{u},H_{d}$ fields need to be normalized properly to get their canonical forms, which amounts to the replacements
$H_{u}\rightarrow H_{u}(2\Im\tau)^{k_{u}/2}~{},~{}~{}~{}H_{d}\rightarrow H_{d}(2\Im\tau)^{k_{d}/2}~{}.$ (18)
So, when the modulus field is fixed by modulus stabilization, the effective $\mu$-parameter will be re-generated to be
$\mu_{eff}=\alpha\Lambda_{SUSY}\frac{Y_{N}^{k_{r}}(\tau_{0})}{\eta(\tau_{0})^{2k_{3}}}(2\Im\tau_{0})^{(k_{u}+k_{d})/2}=\alpha\Lambda_{SUSY}\frac{Y_{N}^{k_{r}}(\tau_{0})}{\eta(\tau_{0})^{2k_{3}}}(2\Im\tau_{0})^{(k_{r}-k_{3}+3)/2}.$ (19)
We need to explain why such $\mu_{eff}$ is much smaller than $\Lambda_{SUSY}$ and be of order the soft SUSY breaking scale $m_{soft}$. We indicate that the smallness of $\mu_{eff}$ is the consequence of large suppression factor, either origin from some power of deviation of modulus VEV from typical fixed point or from the non-trivial power (that determined by the choice of modular weights for $H_{u}$ and $H_{d}$) of some factor that related to the modulus VEV. All of the suppression factors are determined by the modular stabilization mechanism, with the stabilized modulus VEVs affected by SUSY breaking effects.
4 Stabilization of modulus field with SUSY breaking effects
To stabilize the modulus field, effective supergravity scalar potentials from weight $-3$ modular form superpotential can be used to determine the minimum. The modulus VEV can always be restricted to the fundamental domain $D_{H}$
${D}_{F}\equiv\left\{\tau\in\mathcal{H}:-\frac{1}{2}\leq\Re{\tau}<\frac{1}{2},\,|\tau|>1\right\}\cup\left\{\tau\in\mathcal{H}:-\frac{1}{2}<\Re\tau\leq 0,\,|\tau|=1\right\}\,,$ (20)
of the modular group, as any choice of the modulus VEV in the upper-half plane can be uniquely related to a $\tau\in D_{F}$ in the fundamental domain via a modular transformation, which is physically equivalent to such a choice. Discussions on the modulus stabilization with single modulus and multiple moduli can be found in Kobayashi:2019xvz ; NPP:2201.02020 ; PLNSR:2304.14437 ; King:2023snq ; King:2024ssx ; Ding:2024neh ; Higaki:2024jdk ; Kobayashi:2023spx ; Funakoshi:2024yxg .
To guarantee the modular invariance and Kahler transformation invariance of
$e^{G}\equiv e^{K}|W|^{2}~{},$ (21)
the superpotential for modulus should take the following form
$W(\tau)=c_{1}^{3}\frac{H(\tau)}{\eta(\tau)^{6}}.$ (22)
for the Kahler potential of the form
$K_{mod}=-3\log(2\Im\tau)~{},$ (23)
in the unit $M_{Pl}=1$ and $c_{1}M_{Pl}$ denotes the typical scale for the superpotential for dimensionless number $c_{1}$.
The modular invariant holomorphic function $H(\tau)$ is given by
$H(\tau)_{m,n}$ $=$ $\left(j(\tau)-1728\right)^{m/2}j(\tau)^{n/3}{\cal P}\left(j(\tau)\right),$ (24)
$\equiv$ $\left(\frac{G_{6}(\tau)}{\eta(\tau)^{12}}\right)^{m}\left(\frac{G_{4}(\tau)}{\eta(\tau)^{12}}\right)^{n}{\cal P}\left(j(\tau)\right)~{},$
where ${\cal P}$ denotes a polynomial with respect to $j(\tau)$. The Klein j-function is defined as a modular form of zero weight
$j(\tau)$$\equiv$ $\frac{3^{6}5^{3}}{\pi^{12}}\frac{G_{4}(\tau)^{3}}{\eta(\tau)^{24}},$ (25)
$\approx$ $744+q^{-1}+196884q+21493760q^{2}+864299970q^{3}+{\cal O}(q^{4}),$
with $j(i\infty)$ diverges, $j(\omega)=0$ and $j(i)=1728$.
We should note that, in such a stabilization mechanism, fixed points $z=i,\omega\equiv e^{i2\pi/3}$ are always the extrema of the scalar potential, although may not be the global minimum. With proper SUSY breaking contributions, possibly from the ”matter”-like sector, the stabilized modulus value, which located at/near the fixed points $z_{FP}=\omega,i$, can slightly deviate from the original global minimum and possibly uplift the vacuum to dS (de Sitter) one. In fact, the classical 10d supergravity Maldacena-Nunez no-go result MN suggests that any constructions that lead to dS vacuum must include non-classical ingredients as D-branes, O-planes, or quantum corrections. To describe the scenario in which SUSY is linearly realized in a gravitational sector and non-linearly realized in the uplifting SUSY breaking sector, (for example, from anti-D branes) the formalism of non-linearly realized SUSY as formulated by KS can be used for such a system. It was proved in Burgess:2021juk with the non-linear SUSY realization scheme that the supergravity form is stable even if the supermultiplets in SUSY breaking sector are split badly enough to allow some of its members to be integrated out while the others are not.
When SUSY is spontaneously broken in supergravity, the goldstino is eaten by the gravitino via the super-Higgs mechanism. When such a breaking happens at low energies compared with the Planck mass, the goldstino couplings can be described by introducing a (constrained) independent superfield in the supergravity effective action. The goldstino provides a way for an arbitrary non-supersymmetric low-energy theory to realize SUSY non-linearly. The emergence of the constraints may also be viewed as an effective field theory limit in which certain mass parameters become very large and the corresponding modes decouple.
Coupling of non-linearly realized SUSY with nilpotent goldstino $X_{nl}^{2}=0$ can be described by
${\cal L}\supseteq\int d^{4}\theta X_{nl}^{*}X_{nl}+\left[\int d^{2}\theta f_{0}X_{nl}+h.c.\right]~{},$ (26)
with the chiral superfield
$X_{nl}=\frac{\psi_{X}\psi_{X}}{2F_{X}}+\sqrt{2}\theta\psi_{X}+\theta^{2}F_{X}~{},$ (27)
is the IR limit of certain microscopic superfield. As noted in Aparicio:2015psl , for the anti-D3 brane in the KKLT scenario, the representation in terms of $X$ is very convenient and enough to capture all the anti-D3 brane degrees of freedoms since it allows to treat the effect of anti-D3 brane in terms of standard supergravity couplings of matter and moduli superfields to the nilpotent goldstino. From the perspective of low energy effective theory, regardless of the UV origin of SUSY breaking, such a non-linear realization scheme is also convenient for modulus stablization and vacuum uplifting discussions.
When coupled to supergravity, the most general form is given by
[…]
with
$K={X}^{*}_{nl}X_{nl}~{},~{}~{}~{}W=f_{0}X+W_{0}~{}.$ (29)
and $f_{0}=-F_{X}$.
To accommodate the modular flavor symmetry, the Kahler potential and superpotential needs to be adapted to
$K$ $\supseteq$ $-3\log(2\Im\tau)+\frac{{X}^{*}_{nl}X_{nl}}{(2\Im\tau)^{k_{X}}}~{},$
$W$ $\supseteq$ $\left(f_{0}X_{nl}\left[\eta(\tau)\right]^{2k_{X}}+c_{1}^{3}\right)\frac{H(\tau)}{\eta^{6}(\tau)}.$ (30)
The scalar potential for $X$ and $\tau$ can be obtained from
$V=e^{K}\left[K^{\alpha\bar{\beta}}\nabla_{\alpha}W\nabla_{\bar{\beta}}W-3W\overline{W}\right],$ (31)
with
$\nabla_{\alpha}W=W_{\alpha}+K_{\alpha}W~{}.$ (32)
From the Kahler potential, we can calculate the inverse Kahler metric
$K^{\bar{\alpha}\beta}$ $=$ $\left(\begin{array}[]{cc}\frac{(2\Im\tau)^{2}}{3}\left[1-\frac{k_{X}}{3}\frac{|X|^{2}}{(2\Im\tau)^{k_{X}}}\right]&\frac{i(2\Im\tau)k_{X}X}{3}\\ -\frac{i(2\Im\tau)k_{X}{X}^{*}}{3}&(2\Im\tau)^{k_{X}}\left[1+\frac{k_{X}^{2}|X|^{2}}{3(2\Im\tau)^{k_{X}}}\right]\end{array}\right)~{}.$ (35)
Here the indices $\alpha\sim(\tau,X_{nl})$ and the nilpotency of $X_{nl}$ is used in the calculation.
The scalar potential can be calculated to be
$V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})$ (36)
$=$ $\frac{e^{\frac{|X_{nl}|^{2}}{(2\Im\tau)^{k_{X}}}}}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{\frac{(2\Im\tau)^{2}}{3}\left(1-\frac{k_{X}}{3}\frac{|X_{nl}|^{2}}{(2\Im\tau)^{k_{X}}}\right)\left|\tilde{W}_{M}\tilde{c}_{M}+H(\tau)\tilde{W}_{T}\right|^{2}\right.$
$+$ $(2\Im\tau)^{k_{X}}\left[1+\frac{k_{X}^{2}|X_{nl}|^{2}}{3(2\Im\tau)^{k_{X}-1}}\right]\left|H(\tau)\tilde{W}_{X}\right|^{2}$
$+$ $2\Re\left[\frac{i(2\Im\tau)k_{X}X_{nl}}{3}(\tilde{W}_{X}H(\tau))^{*}\left(\tilde{W}_{M}\tilde{c}_{M}+H(\tau)\tilde{W}_{T}\right)\right]$
-- $\left.3\left|\tilde{c}_{W}\right|^{2}\left|{H(\tau)}\right|^{2}\right\}~{},$
with
$\tilde{W}_{M}$$\equiv$ $H^{\prime}(\tau)-\frac{3iH(\tau)}{2\pi}\hat{G}_{2}(\tau,\tau^{*})~{},$
$\tilde{W}_{T}$$\equiv$ $k_{X}f_{0}X_{nl}\left[\eta(\tau)\right]^{2k_{X}}\left[\frac{i}{2\pi}\hat{G}_{2}(\tau,\tau^{*})+\frac{i}{2\Im\tau}+\frac{ic_{1}^{3}X_{nl}^{*}}{f_{0}(2\Im\tau)^{k_{X}+1}\eta(\tau)^{2k_{X}}}\right]~{},$
$\tilde{c}_{M}$ $=$ $f_{0}X_{nl}\left[\eta(\tau)\right]^{2k_{X}}+c_{1}^{3}~{},$
$\tilde{W}_{X}$ $=$ $f_{0}\eta(\tau)^{2k_{X}}\left(1+\frac{|X_{nl}|^{2}}{2(\Im\tau)^{k_{X}}}\right)+\frac{{X}^{*}_{nl}c_{1}^{3}}{(2\Im\tau)^{k_{X}}}.$ (37)
and the constraint $X_{nl}^{2}=0$ is used to eliminate some of the high order $X_{nl}$ terms.
We also use the expression for the derivative of Dedekind $\eta$ function, which is given by
$\frac{\eta^{\prime}(\tau)}{\eta(\tau)}=\frac{i}{4\pi}G_{2}~{},$ (38)
where the holomorphic Eisenstein series $G_{2}$, which is a quasi-modular form
$G_{2}(\gamma\tau)=(c\tau+d)^{2}G_{2}(\tau)-2\pi ic(c\tau+d)~{},$ (39)
is related to its weight two non-holomorphic counterpart $\hat{G}_{2}(\tau,\bar{\tau})$ as
$\hat{G}_{2}(\tau,{\tau}^{*})=G_{2}(\tau)-\frac{\pi}{\Im\tau}~{}.$ (40)
The generic complicate form of the scalar potential may obscure the discussions of the stabilization mechanism. So, we would like to concentrate on its typical simplification forms
  • Scenario with $k_{X}=0$:
    The scalar potential can be simplified to be
    $V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})$ (41)
    $=$ $\frac{e^{|X_{nl}|^{2}}}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\tilde{c}_{M}\right|^{2}\right.$
    $+$ $\left.\left|H(\tau)\right|^{2}\left|f_{0}\left(1+{|X_{nl}|^{2}}\right)+{{X}^{*}_{nl}c_{1}^{3}}\right|^{2}-3\left|\tilde{c}_{M}\right|^{2}\left|{H(\tau)}\right|^{2}\right\}~{},$
    $=$ $\frac{1}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}\left(f_{0}^{2}|X_{nl}|^{2}+c_{1}^{6}+2\Re\left[f_{0}X_{nl}c_{1}^{3}\right]\right)\right.$
    $+$ $\left|H(\tau)\right|^{2}\left(f^{2}_{0}+2f_{0}^{2}{|X_{nl}|^{2}}+2\Re\left[f_{0}c_{1}^{3}{X}_{nl}^{*}\right]+|X_{nl}|^{2}c_{1}^{6}\frac{}{}\right)~{}$
    -- $\left.3\left(f_{0}^{2}|X_{nl}|^{2}+c_{1}^{6}+2\Re\left[f_{0}X_{nl}c_{1}^{3}\right]\right)\left|{H(\tau)}\right|^{2}\right\}~{}$
    $+$ $\frac{|X_{nl}|^{2}}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}c_{1}^{3}\right|^{2}+\left|H(\tau)\right|^{2}f^{2}_{0}-3c_{1}^{6}\left|{H(\tau)}\right|^{2}\right\}~{},$
    with
    $\tilde{W}_{M}$$\equiv$ $H^{\prime}(\tau)-\frac{3iH(\tau)}{2\pi}\hat{G}_{2}(\tau,\tau^{*})~{},$
    $\tilde{c}_{M}$ $=$ $f_{0}X_{nl}+c_{1}^{3}~{},$ (42)
    Note the constraints $X_{nl}^{2}=\bar{X}_{nl}^{2}=0$ are imposed for the scalar potential.
    The modulus VEV can be determined by the extrema conditions of the scalar potential
    $\frac{\partial}{\partial\tau}V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})=\frac{\partial}{\partial\tau^{*}}V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})=0~{},$
    and the minimum condition can be determined by positive definiteness of Hessian matrix in terms of
    $\frac{\partial^{2}}{\partial x^{2}}V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})$ $=$ $2\left(\frac{\partial^{2}}{\partial\tau\partial\tau^{*}}+\Re\frac{\partial^{2}}{\partial\tau\partial\tau}\right)V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*}),$
    $\frac{\partial^{2}}{\partial y^{2}}V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})$ $=$ $2\left(\frac{\partial^{2}}{\partial\tau\partial\tau^{*}}-\Re\frac{\partial^{2}}{\partial\tau\partial\tau}\right)V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*}),$
    $\frac{\partial^{2}}{\partial x\partial y}V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})$ $=$ $-2\Im\frac{\partial^{2}}{\partial\tau\partial\tau}V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*}),$ (43)
    with $\tau\equiv x+iy\equiv\rho_{\tau}e^{i\theta_{\tau}}$, as well as those conditions for $X_{nl}$. It was argued in target:space ; Ding:2024neh that, modular invariance and $\tau\leftrightarrow-\tau^{*}$ symmetry of the scalar potential ensures that the first derivative of scalar potential along certain directions at the boundary of the fundamental domain vanishes
    $\left.\frac{\partial}{\partial x}V\right|_{x=\pm 1/2,0}=0~{},~{}~{}~{}\left.\frac{\partial}{\partial\rho_{\tau}}V\right|_{\rho_{\tau}=1}=0~{}.~{}~{}~{}$ (44)
    Then we will determine the extrema condition
    $\frac{\partial}{\partial\rho}V=\frac{\partial}{\partial\theta}V=0~{},$ (45)
    for $X_{nl}\equiv\rho e^{i\theta}$. It can be calculated that the extreme conditions for $X_{nl}$ reduce to the constraints
    $\frac{\partial}{\partial\theta}V=0~{},\Rightarrow\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}-2\left|H(\tau)\right|^{2}=0~{},$ (46)
    and
    $\frac{\partial}{\partial\rho}V=0~{},$$\Rightarrow$ $\left|H(\tau)\right|^{2}f_{0}^{2}=0~{},$ (49)
    $\Rightarrow$ $\left\{\begin{array}[]{l}\left|H(\tau)\right|^{2}=0,~{}~{}~{}~{}~{}~{}{\rm for}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}f_{0}\neq 0~{}.\\ ~{}~{}~{}~{}~{}~{}~{}f_{0}=0~{},~{}~{}~{}~{}{\rm and/or}~{}~{}~{}\left|H(\tau)\right|^{2}\neq 0.\end{array}\right.$
    So, we have the minimum conditions for $X_{nl}$
    $\left|\tilde{W}_{M}\right|^{2}=\left|H(\tau)\right|^{2}=0~{},$ (50)
    in the case $f_{0}\neq 0$. Substituting back into the scalar potential, we can see that the potential energy at the minimum vanishes. So, the presence non-trivial $X_{nl}$ sector, whose contributions can uplift the scalar potential, can naturally lead to a flat Minkowski universe.
    We will concentrate on the constraint $H(\tau)=0$ when $f_{0}\neq 0$. When $f_{0}=0$, the scalar potential for modulus fields will not be affected by $X_{nl}$ as the two sector decouples in the $k_{X}=0$ limit.
    From the expression of $H(\tau)$, the constraint $H(\tau)=0$ requires either the condition $j(\tau)^{n/3}[j(\tau)-1728]^{m/2}=0$ or ${\cal P}(j(\tau))=0$. Obviously, $\tau=i,\omega$ are its solutions for $m\geq 1$ and $n\geq 1$, respectively. As ${\cal P}(j(\tau))$ is an arbitrary polynomial function of $j(\tau)$, from the fundamental theorem of algebra, arbitrary complex value of $j(\tau)$ can be the root of typical chosen polynomial equation ${\cal P}(j(\tau))=0$, leading to no interesting constraint at all. So, we would like to focus on the constraint
    $\tilde{W}_{M}\equiv H^{\prime}(\tau)-\frac{3iH(\tau)}{2\pi}\hat{G}_{2}(\tau,\tau^{*})=0~{},$ (51)
    that satisfies $H(\tau)=0$. Such a constraint reduces to
    $H^{\prime}(\tau)=\left\{j(\tau)^{n/3}[j(\tau)-1728]^{m/2}{\cal P}(j(\tau))\right\}^{\prime}=0~{},$ (52)
    $\Rightarrow$ $j^{\prime}(\tau)\left[\frac{n}{3}j(\tau)^{n/3-1}[j(\tau)-1728]^{m/2}{\cal P}(j(\tau))+\frac{m}{2}j(\tau)^{n/3}[j(\tau)-1728]^{m/2-1}{\cal P}(j(\tau))\right.$
    $\left.+j(\tau)^{n/3}[j(\tau)-1728]^{m/2}\frac{\partial{\cal P}(j(\tau))}{\partial j(\tau)}\right]=0~{},$
    $\Rightarrow$ $H(\tau)j^{\prime}(\tau)\left[\frac{n}{3}\frac{1}{j(\tau)}+\frac{m}{2}\frac{1}{j(\tau)-1728}+\frac{1}{{\cal P}(j(\tau))}\frac{\partial{\cal P}(j(\tau))}{\partial j(\tau)}\right]=0~{},$
    So $H^{\prime}(\tau)=0$ can trivially hold for $H(\tau)=0$.
    On the other hand, $H^{\prime}(\tau)=0$ can also be satisfied by the other two conditions
    $\left[\frac{n}{3}\frac{1}{j(\tau)}+\frac{m}{2}\frac{1}{j(\tau)-1728}+\frac{1}{{\cal P}(j(\tau))}\frac{\partial{\cal P}(j(\tau))}{\partial j(\tau)}\right]=0~{},$ (53)
    and $j^{\prime}(\tau)=0$.
    We have the following discussions for the condition (53)
    • $n\neq 0$ and $m=0$ ($m\neq 0$ and $n=0$) or $mn\neq 0$:
      The condition (53) reduces to
      $\frac{d{\cal P}(j(\tau))}{{\cal P}(j(\tau))}=-\frac{n}{3}\frac{dj(\tau)}{j(\tau)}~{},$ (54)
      which gives
      ${\cal P}(j(\tau))=C_{0}\left[j(\tau)\right]^{-n/3}~{}.$ (55)
      So, unless $n\in 3\mathbb{Z}$ is negative, ${\cal P}(j(\tau))$ is not a polynomial of $j(\tau)$. Similarly, we can deduce that $m\in 2\mathbb{Z}$ should be negative. Such a deduction can also hold for $m\neq 0$ with $n\neq 0$ case, which also requires $m=n=0$. As $m,n$ are defined to be positive integers, this condition requires $m=n=0$.
    • $m=0$ and $n=0$:
      With the constraints that $m=n=0$, the condition (53) reduces to
      $\frac{\partial{\cal P}(j(\tau))}{\partial j(\tau)}=0~{},$ (56)
      which gives
      ${\cal P}(j(\tau))={\rm Constant}.$ (57)
      as a constant. As $H(\tau)=0$, the constant should be zero.
    So, the condition (53) requires $H(\tau)$ to be zero. As the constraint (53) is trivial, we would like to discuss the condition $j^{\prime}(\tau)=0$. From the derivative of the Klein j-function
    $\frac{\partial}{\partial\tau}j(\tau)=-2\pi i\frac{{E_{6}(\tau)}{E_{4}^{2}(\tau)}}{[\eta(\tau)]^{24}}=0~{},$ (58)
    we can see that its solutions are $\omega$ and $i$.
    So, we can conclude that, in the scenario with $k_{X}=0$ for $f_{0}\neq 0$, the extrema of the scalar potential can only be determined by $H(\tau)=0$ (and the extrema conditions from $X_{nl}$ are also satisfied by $z_{0}=\omega,i$). The corresponding vacuum is Minkowski type. As noted previously, the constraint $H(\tau)=0$ predicts that the fixed points $z_{0}=\omega,i$ are the local extrema. Other solutions of $\tau$ from ${\cal P}(j(\tau))=0$ are essentially possible because of the arbitrariness of the polynomial form of ${\cal P}(j(\tau))$. So, other minimum, which is determined by (44) to locate at the boundary of the fundamental domain or at the imaginary axis, is also possible.
    In fig.1, we show the energies of the scalar potential at modulus values taken within the fundamental domain with various parameter choices. It can be seen that the energies are non-negative when $f_{0}\neq 0$.
    Figure 1: The values of $V(X_{nl},\bar{X}_{nl},\tau,{\tau}^{*})$ within the fundamental domain in the scenario with $k_{X}=0$ for various parameter choices when $f_{0}\neq 0$.
  • Scenario with small $X_{nl}$ (when $|X_{nl}|\rightarrow 0$):
    In this case, at small $|X_{nl}|$ region, the $|X_{nl}|^{2}$ terms can be neglected in addition to the . constraint $X_{nl}^{2}=0$. The scalar potential can be simplified to
    $V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})$ (59)
    $=$ $\frac{1}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}\left(c_{1}^{6}+2\Re\left[f_{0}X_{nl}c_{1}^{3}\left[\eta(\tau)\right]^{2k_{X}}\right]\right)\right.$
    $+$ $\left.\frac{(2\Im\tau)^{2}}{3}\left(2\Re\left[k_{X}f_{0}X_{nl}\left[\eta(\tau)\right]^{2k_{X}}\tilde{W}_{E}\right]\right)H(\tau)c_{1}^{3}\tilde{W}_{M}^{*}\right]$
    $+$ $(2\Im\tau)^{k_{X}}\left|H(\tau)\right|^{2}\left(f_{0}^{2}\left|\eta(\tau)\right|^{4k_{X}}+2\Re\left[f_{0}\eta(\tau)^{2k_{X}}\frac{{X}_{nl}c_{1}^{3}}{(2\Im\tau)^{k_{X}}}\right]\right)$
    $+$ $2\Re\left[\frac{i(2\Im\tau)k_{X}X_{nl}}{3}\left(f_{0}[\eta(\tau)]^{2k_{X}}H(\tau)\right)^{*}\tilde{W}_{M}c_{1}^{3}\right]$
    -- $\left.3\left(c_{1}^{6}+2\Re\left[f_{0}X_{nl}c_{1}^{3}\right]\right)\left|{H(\tau)}\right|^{2}\frac{}{}\right\}~{},$
    with
    $\tilde{W}_{M}$$\equiv$ $H^{\prime}(\tau)-\frac{3iH(\tau)}{2\pi}\hat{G}_{2}(\tau,\tau^{*})~{},$
    $\tilde{W}_{E}$$\equiv$ $\frac{i}{2\pi}\hat{G}_{2}(\tau,\tau^{*})+\frac{i}{2\Im\tau}~{}.$ (60)
    It can be seen easily from the linear terms of $X_{nl}$ that the extrema conditions for $X_{nl}$ will not lead to trivial constraints for $\tau$ with arbitrary choices of $k_{X}$. So, other modulus values lying on the boundary of fundamental domain or imaginary axis other than the fixed point can possibly be the minimum. The (otherwise stabilized) modulus VEV (without contributions from $X_{nl}$ sector) can be destabilized by non-vanishing $X_{nl}$. Note that the $|X_{nl}|^{2}$ terms will dominate the scalar potential at very large $|X_{nl}|$ values, ensuring that the potential can be bounded from below. We will discuss the extrema conditions for the most general case in our subsequent studies.
5 Re-generation of naturally small $\mu_{eff}$ after modulus stabilization
It is known that the fixed points $\tau_{FP}=i,\omega,i\infty$ for modular theories based on the $\Gamma_{N}$ invariance can preserve $Z_{2}^{S},Z_{3}^{ST},Z_{N}^{T}$ residual symmetries, respectively Novichkov:2021evw . One can easily seen from the modular invariance properties of scalar potential that the self-dual fixed points should be the local extrema of the scalar potential, since they act non-trivially (with residual symmetries) on the derivatives of the potential and must vanish. Although they may not be the global minima, experiences from modulus stabilization in string theory indicate that the minima should always lie at/near $\tau_{FP}=i,\omega\equiv e^{i2\pi/3}$ for the tree-level scalar potential, as the scalar potential in general diverges for $\tau\rightarrow i\infty$.
It was found in Novichkov:2021evw that the hierarchies among the charged leptons and quarks mass can follow solely from the properties of the modular forms in the vicinity of symmetric points, avoiding the fine-tuning of the constant parameters (or without the need to introducing extra weighton-like scalar fields). Similar ideas, augmented with additional suppression powers of $q$ (or $(2\Im\tau)^{-1}$), can be used to generate the hierarchy between $\Lambda_{SUSY}$ and $\mu_{eff}$.
Studies on target-space modular symmetry in string theory indicates target:space that all extrema lie on the boundary of the fundamental region or on the imaginary axis. We anticipate that, due to the SUSY breaking contributions from $X_{nl}$ sector, the extrema can change with respect to the change of $f_{0}$ so as that the VEV $\tau_{0}$ can be pushed slightly away from the fixed points $\tau_{FP}=i,\omega\equiv e^{i2\pi/3}$. Such deviation can lead to the small parameters within the $\mu_{eff}$.
Besides, although the scalar potential for modulus field $\tau$ in general diverges when $\tau\rightarrow i\infty$, the scalar potential can still possibly lead to the stabilization of $\tau$ to a value with $\Im\tau_{0}$ being medium because of the contributions from $X_{nl}$. Then, for proper choice of $k_{3}$, the effective $\mu$ parameter can naturally be much smaller than the SUSY scale $\Lambda_{SUSY}$ by a large suppression factor, which origins from both powers of $q$ [or $(2\Im\tau)^{-1}$] and typical suppression factor from modular forms at $\tau_{0}$ (with medium $\Im\tau_{0}$).
5.1 Natural $\mu_{eff}$ with modulus Stabilization near fixed points $\omega\equiv e^{i2\pi/3}$
The Higgs superfields $H_{u},H_{d}$ in MSSM should always be singlets under modular flavor symmetry. They can be assigned with some singlet representations and non-trivial modular weights for their superpotential.
For modular $A_{4}$ symmetry, at the point $\omega$, the following modular forms
$Y_{\bf 1}^{(4)}(\omega)=Y_{\bf 1}^{(8)}(\omega)=Y_{\bf 1^{\prime}}^{(8)}(\omega)=Y_{\bf 1}^{(12)}(\omega)=Y_{\bf 1^{\prime}}^{(12)}(\omega)=Y_{\bf 1}^{(16)}(\omega)=0~{},$ (61)
vanish as well as $Y_{\bf 1^{\prime\prime}}^{(16)}(\omega)$ etc. The small deviation of modulus VEV away from $\omega$ by modulus stabilization mechanism can act as the desired suppression parameter for $\mu_{eff}$.
Expressions of modular forms with larger modular weights can be constructed from $Y_{\bf 3}^{(2)}\equiv(Y_{1},Y_{2},Y_{3})$ by
$Y_{\bf 1^{\prime}}^{(8)}$ $=$ $(Y_{1}^{2}+2Y_{2}Y_{3})(Y_{3}^{2}+2Y_{1}Y_{2})~{},~{}~{}~{}Y_{\bf 1}^{(8)}=(Y_{1}^{2}+2Y_{2}Y_{3})^{2}~{},$
$Y_{\bf 1^{\prime}}^{(8)}$ $=$ $(Y_{1}^{2}+2Y_{2}Y_{3})^{3}(Y_{3}^{2}+2Y_{1}Y_{2})~{},~{}~{}~{}Y_{\bf 1}^{(16)}=(Y_{1}^{2}+2Y_{2}Y_{3})^{4}~{}.$ (62)
From the expansion Petcov:2022fjf at the point $z_{0}=\omega+\epsilon$ (with $\epsilon$ depending on the choices of $f_{0}$ satisfying $|\epsilon|\ll 0$) near $\omega$ point
$\frac{Y_{2}(z_{0})}{Y_{1}(z_{0})}$ $\simeq$ $\omega\left(1+2.235i\epsilon-2.95717\epsilon^{2}+\cdots\right)~{},$
$\frac{Y_{3}(z_{0})}{Y_{1}(z_{0})}$ $\simeq$ $-\frac{1}{2}\omega^{2}\left(1+4.47i\epsilon-10.9096\epsilon^{2}+\cdots\right),$ (63)
one can arrive at the expansion for $Y_{\bf 1}^{(16)}(z_{0})$ etc as
$Y_{\bf 1^{\prime}}^{(8)}(\omega+\epsilon)$ $\simeq$ $(10.5823+6.10969i)\epsilon-(39.9459-69.1884i)\epsilon^{2}+\cdots~{},$
$Y_{\bf 1}^{(16)}(\omega+\epsilon)$ $\simeq$ $1325.96\epsilon^{4}+18871.7i\epsilon^{5}+\cdots~{},$ (64)
with $Y_{1}(\omega)\approx 0.948674$.
[…]
$W\supseteq\alpha\Lambda_{SUSY}\frac{Y_{\bf r}^{(k_{r})}(\tau)}{\eta(\tau)^{2k_{3}}}H_{u}H_{d}~{},$ (65)
with $k_{H_{u}}+k_{H_{d}}=k_{r}-k_{3}+3$ for $62k_{3}\in 24\mathbb{Z}+$ 6 and ${\bf r}(H_{u})\otimes{\bf r}(H_{d})\otimes{\bf r}\supset{\bf 1}$. When the modulus field is stabilized close to the point $\omega$, the effective $\mu$-term is given by
$\mu_{eff}$ $=$ $\alpha\Lambda_{SUSY}\eta^{-2k_{3}}(\omega+\epsilon)Y_{\bf r}^{(k_{r})}(\omega+\epsilon)(2\Im\tau)^{(k_{u}+k_{d})/2}~{},$ (66)
$=$ $\alpha\Lambda_{SUSY}\eta^{-2k_{3}}(\omega)Y_{\bf r}^{(k_{r})}(\omega+\epsilon)(\sqrt{3})^{(k_{r}-k_{3}+3)/2}$
$\approx$ $\alpha\Lambda_{SUSY}\left[0.711506\cdot e^{-0.523599}\right]^{-k_{3}/2}Y_{\bf r}^{(k_{r})}(\omega+\epsilon)(\sqrt{3})^{k_{r}+3/2}~{}.$
So, to obtain much suppressed $\mu_{eff}$, we either need very large negative value $k_{3}$ or large $k_{r}$ to get large powers of $\epsilon$. However, it can be calculated that the suppression from $k_{3}$ power is not very efficient to get tiny number. So we choose to adopt large powers of $\epsilon$ for suppressed $\mu_{eff}$ with proper modular weights and representations for $H_{u},H_{d}$.
For example, we can adopt the choice
$(r_{H_{u}},k_{H_{u}})\sim(~{}{\bf 1^{\prime}},~{}6),~{}~{}(r_{H_{d}},k_{H_{d}})\sim(~{}{\bf 1^{\prime\prime}},~{}7)~{},$ (67)
and the choice of modular forms $Y_{\bf 1}^{(16)}$ (or possibly $Y_{\bf 1}^{(4)},Y_{\bf 1^{\prime}}^{(8)},Y_{\bf 1}^{(12)}$ with proper $k_{H_{u}}$ and $k_{H_{d}}$). So, we have
$\mu_{eff}$ $=$ $\frac{\alpha}{\bar{N}_{Y_{\bf 1}^{(16)}}}\eta(\omega)^{-6}Y_{\bf 1}^{(16)}(\omega+\epsilon)(\sqrt{3})^{(k_{u}+k_{d})/2}~{},$ (68)
$\approx$ $\frac{\alpha}{360.275}(0.8e^{-0.131i})^{-6}\Lambda_{SUSY}1325.96\epsilon^{4}(\sqrt{3})^{13/2}~{},$
$\approx$ $4.98\times 10^{2}e^{0.785i}\alpha\Lambda_{SUSY}\epsilon^{4}~{}.$
Here $\bar{N}_{Y_{\bf 1}^{(16)}}=360.275$ denotes the global normalization factor Petcov:2023fwh , which does not depend on the VEV of the modulus $\tau$ and defined by
$(\overline{{\rm N}}^{(K)}_{\rm Y})^{2}=\int_{D}\int(N^{(K)}_{\rm Y}(\tau,\tau^{*}))^{2}\,\frac{dxdy}{y^{2}}\equiv\int_{D}\int\sum_{i}|Y^{(K)}_{{\bf r}i}(\tau)|^{2}(2\,y)^{K}\frac{dx\,dy}{y^{2}}\,,$ (69)
with
$N_{Y}^{(K)}=\left(\sum\limits_{i}\left|Y_{i}^{(K)}\right|^{2}(2{\rm Im}(\tau))^{K}\right)^{1/2}~{},$ (70)
or
[…]
for non-cusp modular forms, within which $T$ is a real constant, the region $D_{T}$ is the lower part of the fundamental domain bounded from above by the line $y=T$, $T>1$, and $|a_{0}|^{2}$ is the constant $q$-independent term in the $q$-expansion of $\sum_{i}|Y^{(K)}_{{\bf r}i}(\tau)|^{2}$.
For $\epsilon\sim 10^{-5}$ and $\Lambda_{SUSY}$ typically of order the reduced Planck scale, we can obtain the desired $\mu_{eff}$ at electroweak scale. Note that the smallness of deviation parameter $\epsilon$ (away from $\omega$) can be the consequence of SUSY breaking that depends on the choice of $f_{0}$, for example, if the parameter choice with $f_{0}=0$ leads to the global minimum at $\omega$. In general, adopting the modular form $Y_{\bf 1}^{(4(k+2))}$ with $(k\in N)$, we can obtain suppression factor $\epsilon^{k+2}$ for tiny $\epsilon(f_{0})$. We show some of the benchmark points that can lead to the suppression factor in Table 5.1.
$m$ $n$ $f_{0}/M_{\mathrm{Pl}}^{2}$ $c_{1}/M_{\mathrm{Pl}}$ $X_{nl}/M_{\mathrm{Pl}}$ $\epsilon$ 00 $3$ $0.2$ $10.0$ $5.194\times 10^{-4}$ 00 $3$ $0.2$ $0.1$ $0.2$ $8.767\times 10^{-4}$ 00 $2$ $0.2$ $0.1$ $0.02$ $4.227\times 10^{-4}$ 00 $3$ $0.2$ $0.1$ $0.02$ $6.916\times 10^{-4}$ $1$ $3$ $0.2$ $0.1$ $0.02$ $9.396\times 10^{-4}$ $1$ $3$ $0.2$ $0.1$ 00 $4.711\times 10^{-4}$ Table 1: Parameter choices for scenario with $k_{X}=0$, which can stabilize the modulus at the point $\omega+\epsilon$ (for $|\epsilon\sim 10^{-4}|$).
5.2 Natural $\mu_{eff}$ with modulus stabilization at medium $\Im\tau$
Although the scalar potential for modulus in general diverges for $\tau\rightarrow i\infty$, the introduction of nilpotent Goldstino $X$ superfield or multiple moduli/matter fields may alter the shape of the scalar potential at intermediate values of $\tau$, making it possible the stabilization of modulus at $\tau_{0}$ with medium $\Im\tau_{0}$ (for example, $\Im\tau_{0}\sim 5-10$). Detailed discussions on such a possibility will be given in our subsequent studies. In this work, we just concentrate on the implications of modulus stabilization at medium $\Im\tau$ on the SUSY $\mu$-problem. Note that the stabilized VEV $\tau_{0}$ (with medium $\Im\tau_{0}$) can be regarded to lie close to the $i\infty$ fixed point.
To obtain the suppressed effective $\mu$-parameter, in addition to the suppression factor from powers of $q^{1/24}$ or $(2\Im\tau)^{-1}$, we would also like to choose the modular function that vanishes at $\Im\tau=i\infty$. For example, the modular functions satisfy
$Y_{\bf 1^{\prime}}^{(4)}(i\infty)=Y_{\bf 1^{\prime}}^{(8)}(i\infty)=Y_{\bf 1^{\prime\prime}}^{(8)}(i\infty)=0~{},$ (72)
as well as the vanishing singlet components from their direct products, such as the modular functions $Y_{\bf 1}^{(12)}(i\infty)$,$Y_{\bf 1^{\prime}}^{(16)}(i\infty)$.
The asymptotic behavior of the previous modular forms near $i\infty$ within the fundamental domain can be calculated from Petcov:2023vws (for $z_{0}$ with $\Im z_{0}$ being sufficient large)
$Y_{\bf 1^{\prime}}^{(4)}(z_{0})$ $=$ $-12p\epsilon+96p^{4}\epsilon^{4}~{},$
$Y_{\bf 1^{\prime}}^{(8)}(z_{0})$ $=$ $-6p\epsilon-1230p^{4}\epsilon^{4}~{},$
$Y_{\bf 1^{\prime\prime}}^{(8)}(z_{0})$ $=$ $144p^{2}\epsilon^{2}-2304p^{5}\epsilon^{5}~{},$ (73)
with
$\epsilon=\exp\left[-\frac{2\pi}{3}\Im z_{0}\right]~{},~{}~{}p=\exp\left[i\frac{2\pi}{3}\Re z_{0}\right],$ (74)
where $\epsilon\ll 1$.
Let us suppose that the modulus could be stabilized at the imaginary axis (within the fundamental domain) with $z_{0}=it_{0}$ for medium large $t_{0}$, when proper SUSY breaking contributions with non-vanishing $f_{0}$ are taken into account.
The effective $\mu$-term from the bilinear term for $H_{u}H_{d}$ in (65), which satisfies $k_{H_{u}}+k_{H_{d}}=k_{r}-k_{3}+3$ for $62k_{3}\in 24\mathbb{Z}+$ 6, is given by
$\mu_{eff}$ $=$ $\alpha\Lambda_{SUSY}\eta(it_{0})^{-2k_{3}}Y_{\bf r}^{(k_{r})}(it_{0})(2t_{0})^{(k_{u}+k_{d})/2}~{},$ (75)
$\approx$ $\alpha\Lambda_{SUSY}\eta(it_{0})^{-2k_{3}}Y_{\bf r}^{(k_{r})}(it_{0})(2t_{0})^{(k_{r}-k_{3}+3)/2}~{},$
$\approx$ $\alpha\Lambda_{SUSY}e^{\frac{k_{3}\pi t_{0}}{6}}(2t_{0})^{-\frac{k_{3}}{2}}Y_{\bf r}^{(k_{r})}(it_{0})(2t_{0})^{\frac{k_{r}+3}{2}}~{}.$
with proper normalization factors for $H_{u},H_{d}$ from the Kahler potential.
For fixed $k_{r}$, we require that the factor
$\left[e^{\frac{\pi t_{0}}{3}}(2t_{0})^{-1}\right]^{\frac{k_{3}}{2}}\ll 1.$ (76)
As the factor
$e^{\frac{\pi t_{0}}{3}}(2t_{0})^{-1}>1~{},$ (77)
for all $t_{0}>0$, the value of $k_{3}$ needs to be negative.
We can choose the following modular weights and representations
$(r_{H_{u}},k_{H_{u}})\sim(~{}{\bf 1^{\prime\prime}},~{}14),~{}~{}(r_{H_{d}},k_{H_{d}})\sim(~{}{\bf 1},~{}15)~{},$ (78)
and the choice of modular forms $Y_{\bf 1^{\prime\prime}}^{(8)}$ (or possibly $Y_{\bf 1}^{(12)}$, $Y_{\bf 1^{\prime}}^{(16)}$ etc with proper $k_{H_{u}}$ and $k_{H_{d}}$) with $k_{3}=-21$ to generate suppressed $\mu_{eff}$ parameter. In this case, the effective $\mu_{eff}$ parameter is
$\mu_{eff}$ $=$ $\frac{\alpha}{\bar{N}_{Y_{\bf 1^{\prime\prime}}^{(8)}}}\Lambda_{SUSY}\left[e^{\frac{\pi t_{0}}{3}}(2t_{0})^{-1}\right]^{-21/2}144e^{-\frac{4\pi t_{0}}{3}}(2t_{0})^{11/2}~{},$ (79)
$=$ $4.46{\alpha}\Lambda_{SUSY}e^{-29\pi t_{0}/6}(2t_{0})^{16}~{},$
with the global normalization factor ${\bar{N}_{Y_{\bf 1^{\prime}}^{(8)}}}=32.2794$.
It can be calculated that the value of $\mu_{eff}\lesssim 10^{-20}M_{SUSY}$ can be satisfied for $t_{0}\gtrsim 5.58$. If we choose the modular form from direct product of $(Y_{\bf 1^{\prime}}^{(4)})^{k+2}$ with larger $k\in N$, smaller $t_{0}$ is needed to get electroweak scale $\mu_{eff}$ parameter. Tiny value of $\mu_{eff}$ is an joint effect of suppression from powers of $q$ (or $(2\Im\tau)^{-1}$) and the asymptotic suppression behavior of typical modular forms away from the fixed point $i\infty$.
6 Conclusions
It is known that the solutions of SUSY $\mu$ problem involve the forbidden of the bare $\mu$-term via some symmetry and the re-generation of the effective $\mu$-term after the symmetry breaking. We propose that, the symmetry that responsible for the solution of the SUSY $\mu$-problem is a modular flavor symmetry. The bare $\mu$-term is forbidden by modular symmetry after assigning non-vanishing modular weights and/or non-trivial singlet representations for $H_{u},H_{d}$ superfields. Then, the effective $\mu$-term will be re-generated after the modulus field is stabilized, with the corresponding modulus VEV breaks the modular symmetry completely or partially (at the fixed points).
Ordinary single modulus stabilization mechanism can be affected by SUSY breaking contributions. The goldstino, which is necessarily present as a consequence of UV SUSY breaking and mixes with gravitino, can be promoted into a constrained nilpotent superfield $X_{nl}$ and provides a way for an arbitrary non-supersymmetric low-energy theory to realize SUSY non-linearly. We discuss the stabilization mechanism of a single modulus field with the presence of SUSY breaking contributions, described in the non-linear SUSY realization scheme with nilpotent goldstino $X_{nl}$. The presence of $X_{nl}$ contributions can slightly change the original stabilized modulus VEV to a value located in the vicinity of the fixed points. Natural value of $\mu_{eff}$, which is much suppressed relative to the SUSY scale, can be the consequence of series expansion of typical modular forms in terms of the tiny deviation parameter (away from the fixed point $\omega$), which is similar to the mechanism to naturally generate mass hierarchies in the vicinity of the fixed points. For modulus VEV with medium $\Im\tau$, the much suppressed $\mu_{eff}$ is the joint consequence of suppression from powers of $q$ [or $(2\Im\tau)^{-1}$], with the exponent of the power determined by the assignments of the modular weights for $H_{u}$ and $H_{d}$, and the asymptotic suppression behavior of typical modular forms away from the fixed point $i\infty$.
Appendix A Properties of Modular $A_{4}$
The finite modular group $A_{4}\cong\Gamma_{3}$ can generated by
$S^{2}=(ST)^{3}=T^{3}=1~{}.$ (80)
There are 4 inequivalent irreducible representations of $A_{4}$: three singlets $\mathbf{1}$, $\mathbf{1}^{\prime}$, $\mathbf{1}^{\prime\prime}$ and a triplet $\mathbf{3}$.
The triplet representation $\mathbf{3}$ (in the basis where $T$ is diagonal) is given by
$S=\frac{1}{3}\begin{pmatrix}-1~{}&2~{}&2\\ 2~{}&-1~{}&2\\ 2~{}&2~{}&-1\end{pmatrix},~{}\quad~{}T=\begin{pmatrix}1~{}&~{}0~{}&~{}0\\ 0~{}&~{}\omega~{}&~{}0\\ 0~{}&~{}0~{}&~{}\omega^{2}\end{pmatrix}\,.$ (81)
The decompositions of the direct product of $A_{4}$ representations are
$\mathbf{1}^{\prime}\otimes\mathbf{1}^{\prime}=\mathbf{1}^{\prime\prime},~{}~{}~{}\mathbf{1}^{\prime}\otimes\mathbf{1}^{\prime\prime}=\mathbf{1},~{}~{}~{}\mathbf{1}^{\prime\prime}\otimes\mathbf{1}^{\prime\prime}=\mathbf{1}^{\prime}\,,$
$\mathbf{3}\otimes\mathbf{3}=\mathbf{1}\oplus\mathbf{1^{\prime}}\oplus\mathbf{1^{\prime\prime}}\oplus\mathbf{3}_{S}\oplus\mathbf{3}_{A}\,,$ (82)
where $\mathbf{3}_{S(A)}$ denotes the symmetric (antisymmetric) combinations, respectively.
Given two $A_{4}$ triplet representations $\alpha=(\alpha_{1},\alpha_{2},\alpha_{3})$ and $\beta=(\beta_{1},\beta_{2},\beta_{3})$, the irreducible representations obtained from their direct products are:
$\mathbf{1}=\alpha_{1}\beta_{1}+\alpha_{2}\beta_{3}+\alpha_{3}\beta_{2}\,,$
$\mathbf{1}^{\prime}=\alpha_{3}\beta_{3}+\alpha_{1}\beta_{2}+\alpha_{2}\beta_{1}\,,$
$\mathbf{1}^{\prime\prime}=\alpha_{2}\beta_{2}+\alpha_{1}\beta_{3}+\alpha_{3}\beta_{1}\,,$
$\mathbf{3}_{S}=(2\alpha_{1}\beta_{1}-\alpha_{2}\beta_{3}-\alpha_{3}\beta_{2},2\alpha_{3}\beta_{3}-\alpha_{1}\beta_{2}-\alpha_{2}\beta_{1},2\alpha_{2}\beta_{2}-\alpha_{1}\beta_{3}-\alpha_{3}\beta_{1})\,,$
$\mathbf{3}_{A}=(\alpha_{2}\beta_{3}-\alpha_{3}\beta_{2},\alpha_{1}\beta_{2}-\alpha_{2}\beta_{1},\alpha_{3}\beta_{1}-\alpha_{1}\beta_{3})\,.$ (83)
Modular weight $k=2$ and the representation ${\bf r}={\bf 3}$ are given as $Y^{(2)}_{\mathbf{3}}=\left(Y_{1},Y_{2},Y_{3}\right)^{T}$ with
$Y_{1}(\tau)$ $=$ $\frac{i}{2\pi}\left[\frac{\eta^{\prime}(\tau/3)}{\eta(\tau/3)}+\frac{\eta^{\prime}((\tau+1)/3)}{\eta((\tau+1)/3)}+\frac{\eta^{\prime}((\tau+2)/3)}{\eta((\tau+2)/3)}-\frac{27\eta^{\prime}(3\tau)}{\eta(3\tau)}\right],$
$Y_{2}(\tau)$ $=$ $\frac{-i}{\pi}\left[\frac{\eta^{\prime}(\tau/3)}{\eta(\tau/3)}+\omega^{2}\frac{\eta^{\prime}((\tau+1)/3)}{\eta((\tau+1)/3)}+\omega\frac{\eta^{\prime}((\tau+2)/3)}{\eta((\tau+2)/3)}\right],$
$Y_{3}(\tau)$ $=$ $\frac{-i}{\pi}\left[\frac{\eta^{\prime}(\tau/3)}{\eta(\tau/3)}+\omega\frac{\eta^{\prime}((\tau+1)/3)}{\eta((\tau+1)/3)}+\omega^{2}\frac{\eta^{\prime}((\tau+2)/3)}{\eta((\tau+2)/3)}\right]\,,$ (84)
where $\eta(\tau)$ is the Dedekind eta-function,
$\eta(\tau)=q^{1/24}\prod_{n=1}^{\infty}(1-q^{n}),~{}~{}~{}q=e^{2\pi i\tau}\,.$ (85)
The $q-$expansions of $Y_{1,2,3}(\tau)$ are given as
$Y_{1}(\tau)=1+12q+36q^{2}+12q^{3}+84q^{4}+72q^{5}+\dots\,,$
$Y_{2}(\tau)=-6q^{1/3}(1+7q+8q^{2}+18q^{3}+14q^{4}+\dots)\,,$
$Y_{3}(\tau)=-18q^{2/3}(1+2q+5q^{2}+4q^{3}+8q^{4}+\dots)\,.$ (86)
Acknowledgements.
This work was supported by the National Natural Science Foundation of China under grant numbers Nos.12075213 and 12335005, by the Natural Science Foundation for Distinguished Young Scholars of Henan Province under grant number 242300421046.

### v3

@@
 Abstract
-We propose to solve the SUSY $\mu$-problem in the framework of modular flavor symmetry.
-The bare $\mu$-term is forbidden by modular symmetry and the effective $\mu$-term is re-generated after the modulus field is stabilized.
-We discuss the stabilization mechanism of a single modulus field with the presence of SUSY breaking contributions described in the non-linear realized SUSY scheme with nilpotent goldstino $X_{nl}$ superfield.
-The presence of $X_{nl}$ contributions can possibly slightly change the original stabilized modulus VEV to a value located in the vicinity of some fixed points.
-Natural value of $\mu_{eff}$, which is much smaller than the SUSY scale, can either be the consequence of expansion of typical modular forms in terms of the tiny deviation parameter (away from the fixed point $\omega$), or be the joint consequence of suppression from both powers of $q$ [or $(2\Im\tau)^{-1}$] and the asymptotic suppression behavior of typical modular forms away from the fixed point $i\infty$.
+We propose a solution to the SUSY $\mu$-problem within the framework of modular flavor symmetry.
+The explicit $\mu$-term is prohibited by modular symmetry, and an effective $\mu$-term is regenerated following the stabilization of the modulus field.
+We examine the stabilization mechanism of a single modulus field with the presence of SUSY breaking contributions described by the non-linear SUSY realization scheme involving a nilpotent Goldstino $\textbf{X}_{nl}$ superfield.
+The inclusion of $\textbf{X}_{nl}$ contributions can slightly alter the original stabilized modulus VEV to a value near some fixed point.
+A natural small $\mu_{eff}$, significantly smaller than the SUSY scale, can result from either the expansion of typical modular forms using a small deviation parameter near the fixed point $\omega$, or from the combined effects of suppression by powers of $q^{1/24}$ [or $(2\Im\tau)^{-1}$] along with the asymptotic suppression behavior of typical modular forms away from the fixed point $i\infty$, taking the form of appropriate power of the tiny deviation parameter.
 1 Introduction
-Low energy supersymmetry (SUSY) is well motivated theoretically and regarded as one of the most appealing extensions of the Standard Model.
-Not only the quadratic divergences related to the Higgs mass can be eliminated, but also successful gauge coupling unification can be realized in the framework of low energy SUSY.
-Especially, the observed 125 GeV scalar, discovered by both the ATLAS ATLAS and CMS CMS collaborations of LHC, falls miraculously within the narrow $115\sim 135$ GeV ’window’ required by minimal supersymmetric standard model (MSSM) and it can be seen as a triumph of low energy SUSY.
-Although the framework of low energy SUSY is attracting, it still has several unsatisfactory features.
-In addition to the problem of null observation of conclusive signals for superparticles at the LHC, the MSSM is also bothered by the SUSY $\mu$-problem: the $\mu$-parameter is a SUSY-respecting parameter appearing in the superpotential, while it is required to be roughly of order the soft SUSY breaking scale to allow a Higgs VEV of order 174 GeV without too much miraculous cancelation.
-There are many proposals to tackle the $\mu$-problem, for example, by singlet extension to next-to-minimal supersymmetric standard model (NMSSM) or by non-renormalizable Lagrangian terms in the superpotential via Kim-Nilles mechanism KN (or in the Kahler potential via Giudice-Masiero mechanism GM ).
-Most solutions consist of two steps Bae:2019dgg : First, the appearance of $\mu$-parameter is forbidden, usually by some discrete or continuous symmetry; Second, the $\mu$ term is re-generated via some mechanism such as symmetry breaking.
-One can see in Bae:2019dgg some detailed realizations of the two steps for the $\mu$-problem.
-We would like to tackle the $\mu$-problem with a similar but different approach.
+Low-energy supersymmetry (SUSY) is well motivated from a theoretical standpoint and is considered to be one of the most compelling extensions of the Standard Model.
+It not only address the quadratic divergence problem associated with the Higgs mass but also enables successful gauge coupling unification.
+Notably, the observed 125 GeV scalar, discovered by both the ATLAS ATLAS and CMS CMS collaborations of LHC, falls miraculously within the narrow $115\sim 135$ GeV ’window’ predicted by the minimal supersymmetric standard model (MSSM).
+This can be regarded as strong supports for low energy SUSY.
+Although the framework of low energy SUSY is attractive, it still has several unsatisfactory features.
+In addition to the lack of conclusive signals for superparticles at the LHC, the MSSM is also bothered by the SUSY $\mu$-problem.
+The $\mu$-parameter, a SUSY preserving parameter in the superpotential, however, is required (without too much miraculous cancelation) to be roughly of the order of the soft SUSY breaking scale, as determined by the minimization of the MSSM Higgs potential.
+Many proposals have been put forward to address the $\mu$-problem, for example, by singlet extension of MSSM to next-to-minimal supersymmetric standard model (NMSSM) or by introducing non-renormalizable terms in the superpotential via Kim-Nilles mechanism KN (or in the Kahler potential via Giudice-Masiero mechanism GM ).
+Most solutions involve two steps Bae:2019dgg : First, the appearance of $\mu$-parameter is prohibited, usually by some discrete or continuous symmetry.
+Second, the $\mu$ term is regenerated via some mechanism such as symmetry breaking.
+For detailed realizations of these two steps, one can refer to Bae:2019dgg .
+We propose to tackle the $\mu$-problem with a similar but different approach.
 The symmetry adopted to forbid the $\mu$-parameter is a modular flavor symmetry.
-Besides, the re-generation of $\mu$ term is realized after modular symmetry breaking with the modulus stabilization mechanism accompanied by typical SUSY breaking contributions.
-Naturally suppressed $\mu$-parameter can thus be realized.
+The regeneration of $\mu$ term is realized following the modular symmetry breaking, with the modulus stabilization mechanism accompanied by typical SUSY breaking contributions.
 Modular symmetry is a geometrical symmetry of torus $T^{2}$ and it corresponds to a change of the corresponding cycle basis.
-Modular symmetry can naturally arise in string theory, for example, (exact) moduli spaces of Calabi-Yau threefolds, like the quintic and fairly generically in large classes of elliptic Calabi-Yau threefolds.
+Modular symmetry can naturally arise in string theory, for example, in the (exact) moduli spaces of Calabi-Yau threefolds like the quintic, and more generically in large classes of elliptic Calabi-Yau threefolds.
 Recently, modular flavor symmetry was proposed in Feruglio:2017spp to explain the flavor puzzles of the standard model (plus the massive neutrinos), which no longer necessitates the introduction of the flavon fields that may cause difficulties in flavor model buildings.
 For reviews, see Feruglio:2019ybq ; Kobayashi:2023zzc ; Ding:2023htn ; Ding:2024ozt .
-The modular symmetry can constrain stringently the superpotential as the relevant Yukawa couplings are modular forms.
+In the modular flavor symmetry framework, the superpotential is tightly constrained by the modular symmetry, with the relevant Yukawa couplings being modular forms of a certain level $N$, which are functions of a modulus field.
 The matter superfields and the modular forms in the superpotential transform in some representations of certain finite modular group $\Gamma_{N}\equiv\bar{\Gamma}/\bar{\Gamma}(N),N\leq 5$ (or its double cover $\Gamma_{N}^{\prime}$).
-In modular flavor symmetry framework, the masses and mixings for the quarks and leptons can be predicted with very few input parameters.
-Modular flavor models based on $\Gamma_{2}\simeq S_{3}$ Kobayashi:2018vbk ; Okada:2019xqk ; Du:2020ylx , $\Gamma_{3}\simeq A_{4}$ Kobayashi:2018vbk ; Petcov:2018snn ; Criado:2018thu ; Kobayashi:2018scp ; Okada:2018yrn ; Novichkov:2018yse ; Okada:2020ukr ; Petcov:2022fjf ; Nomura:2024ctl ; Pathak:2024sei , $\Gamma_{4}\simeq S_{4}$ Penedo:2018nmg ; Petcov:2018snn ; Novichkov:2018ovf ; Kobayashi:2019xvz ; Wang:2020dbp ; Qu:2021jdy , $\Gamma_{5}\simeq A_{5}$ Petcov:2018snn ; Novichkov:2018nkm ; Ding:2019xna ; Yao:2020zml ; deMedeirosVarzielas:2022ihu ; Abbas:2024bbv and non-holomorphic modular flavor models Qu:2024rns ; Nomura:2024atp ; Ding:2024inn ; Li:2024svh had been adopted in various studies to explain the low energy flavor structures of the Standard Model.
-Besides, modular flavor symmetries can combine with the GUT framework to increase greatly the prediction ability of the models, such as the combination of $\Gamma_{2},\Gamma_{3}$ with $SU(5)$ deAnda:2018ecu ; Kobayashi:2019rzp ; Du:2020ylx ; Chen:2021zty ; Zhao:2021jxg ; King:2021fhl ; Ding:2021zbg , SO(10) Ding:2021eva ; Ding:2022bzs or flipped SU(5) Charalampous:2021gmf ; Du:2022lij ; King:2024gsd GUT models.
-The vacuum expectation value (VEV) of the modulus field can be the only source of flavour symmetry breaking, it can also be the only source of CP symmetry breaking.
-The VEV of the modulus field $\tau$ will in general break the modular symmetry completely except three inequivalent fixed points $\tau=i\infty,i,e^{i2\pi/3}$ in the fundamental domain.
-The modular forms, thus the Yukawa coupling, and the fermion mass matrices will be fixed after the modular flavour symmetry is fully or partially broken by VEV of $\tau$.
-From a top-down perspective, the value of modulus VEV can be determined dynamically with modular-invariant scalar potential in 4d $N=1$ supergravity, which is also crucial to give nonzero masses for modulus field to render it phenomenological acceptable.
-Experiences from heterotic string theory on moduli stabilization target:space ; Leedom:2022zdm ; Gonzalo:2018guu indicate that the scalar potential favors minima at the boundary of the fundamental domain or at points inside the fundamental domain that close to the fixed point $\omega=e^{i2\pi/3}$ with negative vacuum energy.
-In the presence of additional ”matter”-like fields, it is possible to obtain dS vacua with additional uplifting contributions, such as the introduction of an anti-D3 brane that provides a positive uplifting potential, which would induces a SUSY breaking vacuum shift to a tiny positive vacuum energy.
-The presence of the anti-brane like sources can be described in terms of a nilpotent goldstino superfield KS in a supersymmetric effective field.
-Although SUSY is non-linearly realized, the couplings of chiral, gauge and gravity superfields coupled to the constrained nilpotent goldstino superfield $X$ can be used to describe spontaneous SUSY breaking in supergravity Burgess:2021juk ; Aparicio:2015psl ; NL:SUGRA1 ; NL:SUGRA2 ; NL:SUGRA3 ; NL:SUGRA4 ; NL:SUGRA5 ; NL:SUGRA6 and the presence of the anti-brane type sources with a supersymmetric action.
+In this framework, the masses and mixings of quarks and leptons can be predicted with very few input parameters.
+Modular flavor models based on $\Gamma_{2}\simeq S_{3}$ Kobayashi:2018vbk ; Okada:2019xqk ; Du:2020ylx , $\Gamma_{3}\simeq A_{4}$ Kobayashi:2018vbk ; Petcov:2018snn ; Criado:2018thu ; Kobayashi:2018scp ; Okada:2018yrn ; Novichkov:2018yse ; Okada:2020ukr ; Petcov:2022fjf ; CentellesChulia:2023osj ; Kumar:2023moh ; Nomura:2024ctl ; Pathak:2024sei , $\Gamma_{4}\simeq S_{4}$ Penedo:2018nmg ; Petcov:2018snn ; Novichkov:2018ovf ; Kobayashi:2019xvz ; Wang:2020dbp ; Qu:2021jdy , $\Gamma_{5}\simeq A_{5}$ Petcov:2018snn ; Novichkov:2018nkm ; Ding:2019xna ; Yao:2020zml ; deMedeirosVarzielas:2022ihu ; Abbas:2024bbv and non-holomorphic modular flavor models Qu:2024rns ; Nomura:2024atp ; Ding:2024inn ; Li:2024svh had been used in various studies to explain the low energy flavor structures of the Standard Model.
+Modular symmetry can also be used to solve the strong CP problem Feruglio:2023uof ; Feruglio:2024ytl .
+Moreover, modular flavor symmetries can be combined with the GUT framework to further enhance the predictive power of the models, such as combining $\Gamma_{2},\Gamma_{3}$ with $SU(5)$ deAnda:2018ecu ; Kobayashi:2019rzp ; Du:2020ylx ; Chen:2021zty ; Zhao:2021jxg ; King:2021fhl ; Ding:2021zbg , SO(10) Ding:2021eva ; Ding:2022bzs or flipped SU(5) Charalampous:2021gmf ; Du:2022lij ; King:2024gsd GUT models.
+The vacuum expectation value (VEV) of the modulus field can be the only source of flavour symmetry breaking.
+It can also be the only source of CP symmetry breaking Novichkov:2019sqv ; Feruglio:2021dte ; Feruglio:2024ytl .
+The VEV of the modulus field $\tau$ generally break the modular symmetry completely, except at three inequivalent fixed points $\tau=i\infty,i,e^{i2\pi/3}$ in the fundamental domain.
+The modular forms (thus the Yukawa couplings) and the fermion mass matrices are fixed once the modular flavour symmetry is fully or partially broken by the VEV of $\tau$.
+From a top-down perspective, the value of modulus VEV can be determined dynamically with a modular-invariant scalar potential in 4d $N=1$ supergravity.
+This is also crucial for giving nonzero masses to the modulus field, rendering it phenomenological acceptable.
+Experiences from heterotic string theory on moduli stabilization target:space ; Leedom:2022zdm ; Gonzalo:2018guu indicate that the scalar potential favors minima at the boundary of the fundamental domain or at points inside the fundamental domain that are close to the fixed point $\omega=e^{i2\pi/3}$ with negative vacuum energy.
+In the presence of additional ”matter”-like fields, it is possible to obtain a dS vacua with additional uplifting contributions, such as the introduction of an anti-D3 brane that provides a positive uplifting potential, which would induce a SUSY breaking vacuum shift to a tiny positive vacuum energy.
+The presence of the anti-brane-like sources can be described in terms of a nilpotent Goldstino superfield KS in a supersymmetric effective field theory.
+In the framework of non-linearly realized SUSY, the couplings of chiral, gauge and gravity superfields to the constrained nilpotent Goldstino superfield $X$ can be used to describe spontaneous SUSY breaking in supergravity Burgess:2021juk ; Aparicio:2015psl ; NL:SUGRA1 ; NL:SUGRA2 ; NL:SUGRA3 ; NL:SUGRA4 ; NL:SUGRA5 ; NL:SUGRA6 and the presence of the anti-brane type sources with a supersymmetric action.
+Contributions from nilpotent chiral superfield can uplift the scalar potential responsible for the stabilization of the modulus field.
 This paper is organized as follows.
 In Sec 2, we briefly review the $\Gamma_{N}$ modular flavor symmetry.
-In Sec 3, we discuss the solution of SUSY $\mu$-problem with modular symmetry.
-In Sec 4, we discuss the stabilization of singlet modulus field, with the presence of SUSY breaking contributions described in the non-linear realized SUSY scheme with nilpotent goldstino $X_{nl}$ superfield.
-In Sec 5, we discuss the re-generation of naturally small $\mu_{eff}$ after modulus stabilization.
-Sec 6 contains our conclusions.
+In Sec 3, we discuss the solution of SUSY $\mu$-problem using modular flavor symmetry.
+In Sec 4, we discuss the stabilization of singlet modulus field in the presence of SUSY breaking contributions described by the non-linear realized SUSY scheme with the nilpotent Goldstino ${\textbf{X}}_{nl}$ superfield.
+In Sec 5, we discuss the regeneration of a naturally small $\mu_{eff}$ after modulus stabilization, when the modulus field is stabilized near some fixed points, such as $\omega\equiv e^{i2\pi/3}$ and $i\infty$ Sec 6 contains our conclusions.
 2 Modular $\Gamma_{N}$ symmetry
-The full modular group is the group of linear fraction transformations acting on the complex modulus $\tau$ in the upper half complex plane
+The full modular group is defined as the set of linear fractional transformations acting on the complex modulus $\tau$ in the upper half complex plane
 $\gamma=\left(\begin{array}[]{cc}a&b\\ c&d\end{array}\right):\tau\longmapsto\gamma\tau=\frac{a\tau+b}{c\tau+d},~{}~{}\text{with}~{}~{}a,b,c,d\in\mathbb{Z},~{}~{}ad-bc=1,~{}~{}~{}{\rm Im}\tau>0\,.$ (1)
-It can be generated by two generators $S$ and $T$
+This group can be generated by two generators, $S$ and $T$, which act as follows
 $S:\tau\mapsto-\frac{1}{\tau},~{}~{}~{}~{}\quad T:\tau\mapsto\tau+1\,,$ (2)
-which satisfy
+These generators satisfy the relations
 $S^{4}=1,~{}~{}(ST)^{3}=1,~{}~{}S^{2}T=TS^{2}.$ (3)
-The two generators can be represented by the following $SL(2,Z)$ matrices,
+The $S,T$ generators can be represented by the following $SL(2,Z)$ matrices,
 $S=\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right),~{}~{}~{}\quad T=\left(\begin{array}[]{cc}1&1\\ 0&1\end{array}\right)\,.$ (4)
-Finite modular groups $\Gamma_{N}$ Feruglio:2017spp are defined by $\Gamma_{N}=\overline{\Gamma}/\overline{\Gamma}(N)$, with $\overline{\Gamma}\equiv SL(2,Z)/\{\pm 1\}$ and $\overline{\Gamma}(N)$ defined by
-$\overline{\Gamma}(N)=\left\{\begin{array}[]{c}\Gamma(N),~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}N\geq 3\\ \Gamma(N)/\{\pm 1\},~{}~{}~{}N=2\end{array}\right.$ (7)
+Finite modular groups $\Gamma_{N}$ Feruglio:2017spp are defined as $\Gamma_{N}=\overline{\Gamma}/\overline{\Gamma}(N)$, where $\overline{\Gamma}\equiv SL(2,Z)/\{\pm 1\}$ and $\overline{\Gamma}(N)$ is given by
+$\overline{\Gamma}(N)=\left\{\begin{array}[]{c}\Gamma(N),~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}N\geq 3\\ \Gamma(N)/\{\pm 1\},~{}~{}~{}N=2\end{array}\right.~{}~{}~{},$ (7)
 with $\Gamma(N)$ the principal congruence subgroup of level $N$
@@
 Modular forms of weight $2k$ and level $N$ span a linear space of finite dimension.
-They can be arranged into some modular multiplets $Y_{\mathbf{r}}^{2k}$ that transform as irreducible representation $\mathbf{r}$ of the finite modular group $\Gamma_{N}$ up to automorphy factor Feruglio:2017spp , i.e.
+They can be arranged into some modular multiplets $Y_{\mathbf{r}}^{(2k)}$ that transform as irreducible representation $\mathbf{r}$ of the finite modular group $\Gamma_{N}$ up to automorphy factor Feruglio:2017spp , i.e.
 $Y_{\mathbf{r}}^{(2k)}(\gamma\tau)=(c\tau+d)^{2k}\rho_{\mathbf{r}}(\gamma)Y_{\mathbf{r}}^{(2k)},~{}~{}~{}~{}\forall~{}\gamma\in\overline{\Gamma}\,,$ (10)
@@
 3 Natural $\mu$-term in modular flavor model
-As noted in Bae:2019dgg , to solve the SUSY $\mu$ problem, the appearance of $\mu$-parameter in the superpotential should be forbidden by some symmetry.
-We propose that the superfields $H_{u}$, $H_{d}$ transform non-trivially under modular transformation so as that ordinary $\mu H_{u}H_{d}$ term is forbidden by modular symmetry.
+As noted in Bae:2019dgg , to solve the SUSY $\mu$ problem, the appearance of a $\mu$ term in the superpotential should be prohibited by some symmetry.
+We propose that the superfields $H_{u}$, $H_{d}$ transform non-trivially under modular transformations, thereby prohibiting the ordinary $\mu H_{u}H_{d}$ term by modular symmetry.
 The relevant Lagrangian can be given as
@@
 $+$ $\int d^{2}\theta\left[\alpha\Lambda_{SUSY}\frac{Y_{N}^{k_{r}}(\tau)}{\eta(\tau)^{2k_{3}}}H_{u}H_{d}+\cdots\right],$
-with the coefficient for bilinear $H_{u}H_{d}$ term given by
-$\mu=\alpha\Lambda_{SUSY}\frac{Y_{N}^{(k_{r})}(\tau)}{\eta(\tau)^{2k_{3}}}~{},$ (13)
-for $k_{H_{u}}+k_{H_{d}}=k_{r}-k_{3}+3$ with $62k_{3}\in 24\mathbb{Z}+$ 6 to guarantee the modular invariance of
+where the coefficient for bilinear $H_{u}H_{d}$ term is given by
+$\tilde{\mu}=\alpha\Lambda_{SUSY}\frac{Y_{N}^{(k_{r})}(\tau)}{\eta(\tau)^{2k_{3}}}~{},$ (13)
+for $k_{H_{u}}+k_{H_{d}}=k_{r}-k_{3}+3$ to ensure the modular invariance of
 $e^{G}\equiv e^{K}|W|^{2}~{}.$ (14)
 Here, $\alpha\sim{\cal O}(1)$ is a dimensionless parameter and $\Lambda_{SUSY}$ is typically of order the SUSY scale.
-The Dedekind $\eta$-function is a modular form with modular weight $1/2$ under modular transformation up to a phase, whose transformation under $\gamma$ can be given as
-$\eta(\gamma\tau)=\epsilon_{\eta}(\gamma)(c\tau+d)^{1/2}\eta(\tau)~{},~{}~{}\forall\gamma\in\overline{\Gamma}~{}.$ (15)
-with $\epsilon_{\eta}(\gamma)$ depending on $\gamma$ (but not on $\tau$) and being given in terms of the Rademacher $\Phi(\gamma)$ function
-$\epsilon_{\eta}(\gamma)=\exp\left[\frac{2\pi i}{24}\Phi(\gamma)-i\frac{\pi}{4}\right].$ (16)
+The Dedekind $\eta$-function is a modular form with modular weight $1/2$ under modular transformation up to a phase.
+Its transformation under $\gamma$ is given by
+$\eta(\gamma\tau)=\epsilon_{\eta}(\gamma)(c\tau+d)^{1/2}\eta(\tau)~{},~{}~{}\forall\gamma\in\overline{\Gamma}~{},$ (15)
+where $\epsilon_{\eta}(\gamma)$ depends on $\gamma$ (but not on $\tau$) and can be expressed in terms of the Rademacher $\Phi(\gamma)$ function
+$\epsilon_{\eta}(\gamma)=\exp\left[\frac{2\pi i}{24}\Phi(\gamma)-i\frac{\pi}{4}\right],$ (16)
 that satisfies $\epsilon_{\eta}(\gamma)^{24}=1$ DHoker:2022dxx .
-The expansion of $\eta(\tau)$ function is
+The expansion of the $\eta(\tau)$ function is
 $\eta(\tau)$ $=$ $q^{1/24}\sum\limits_{n=-\infty}^{\infty}(-1)^{n}q^{n(3n-1)/2},$ (17)
@@
 $=$ $q^{1/24}\left[1-q-q^{2}+q^{5}+q^{7}-q^{12}-\cdots\right],~{}~{}~{}~{}~{}q\equiv e^{i2\pi\tau}~{}.$
-It is obvious that the Kahler potential for superfields with non-trivial modular weights, for example, the $H_{u}$ and $H_{d}$ superfields, are non-minimal.
-Therefore, after the moduli fields $\tau$ acquire a VEV by modulus stabilization mechanism, the $H_{u},H_{d}$ fields need to be normalized properly to get their canonical forms, which amounts to the replacements
-$H_{u}\rightarrow H_{u}(2\Im\tau)^{k_{u}/2}~{},~{}~{}~{}H_{d}\rightarrow H_{d}(2\Im\tau)^{k_{d}/2}~{}.$ (18)
-So, when the modulus field is fixed by modulus stabilization, the effective $\mu$-parameter will be re-generated to be
-$\mu_{eff}=\alpha\Lambda_{SUSY}\frac{Y_{N}^{k_{r}}(\tau_{0})}{\eta(\tau_{0})^{2k_{3}}}(2\Im\tau_{0})^{(k_{u}+k_{d})/2}=\alpha\Lambda_{SUSY}\frac{Y_{N}^{k_{r}}(\tau_{0})}{\eta(\tau_{0})^{2k_{3}}}(2\Im\tau_{0})^{(k_{r}-k_{3}+3)/2}.$ (19)
-We need to explain why such $\mu_{eff}$ is much smaller than $\Lambda_{SUSY}$ and be of order the soft SUSY breaking scale $m_{soft}$.
-We indicate that the smallness of $\mu_{eff}$ is the consequence of large suppression factor, either origin from some power of deviation of modulus VEV from typical fixed point or from the non-trivial power (that determined by the choice of modular weights for $H_{u}$ and $H_{d}$) of some factor that related to the modulus VEV.
-All of the suppression factors are determined by the modular stabilization mechanism, with the stabilized modulus VEVs affected by SUSY breaking effects.
-4 Stabilization of modulus field with SUSY breaking effects
-To stabilize the modulus field, effective supergravity scalar potentials from weight $-3$ modular form superpotential can be used to determine the minimum.
-The modulus VEV can always be restricted to the fundamental domain $D_{H}$
-${D}_{F}\equiv\left\{\tau\in\mathcal{H}:-\frac{1}{2}\leq\Re{\tau}<\frac{1}{2},\,|\tau|>1\right\}\cup\left\{\tau\in\mathcal{H}:-\frac{1}{2}<\Re\tau\leq 0,\,|\tau|=1\right\}\,,$ (20)
-of the modular group, as any choice of the modulus VEV in the upper-half plane can be uniquely related to a $\tau\in D_{F}$ in the fundamental domain via a modular transformation, which is physically equivalent to such a choice.
-Discussions on the modulus stabilization with single modulus and multiple moduli can be found in Kobayashi:2019xvz ; NPP:2201.02020 ; PLNSR:2304.14437 ; King:2023snq ; King:2024ssx ; Ding:2024neh ; Higaki:2024jdk ; Kobayashi:2023spx ; Funakoshi:2024yxg .
-To guarantee the modular invariance and Kahler transformation invariance of
+It is evident that the Kahler potential for superfields with non-trivial modular weights, such as $H_{u}$ and $H_{d}$ superfields, is non-minimal.
+Therefore, after the modulus field $\tau$ acquires a VEV through the modulus stabilization mechanism, the $H_{u},H_{d}$ fields need to be normalized properly to get their canonical kinetic forms.
+This normalization corresponds to the following replacements
+$H_{u}\rightarrow H_{u}(2\Im\tau)^{k_{H_{u}}/2}~{},~{}~{}~{}H_{d}\rightarrow H_{d}(2\Im\tau)^{k_{H_{d}}/2}~{}.$ (18)
+Thus, when the modulus field is stabilized, the effective $\mu$-parameter is regenerated as
+$\mu_{eff}=\alpha\Lambda_{SUSY}\frac{Y_{N}^{k_{r}}(\tau_{0})}{\eta(\tau_{0})^{2k_{3}}}(2\Im\tau_{0})^{(k_{H_{u}}+k_{H_{d}})/2}=\alpha\Lambda_{SUSY}\frac{Y_{N}^{k_{r}}(\tau_{0})}{\eta(\tau_{0})^{2k_{3}}}(2\Im\tau_{0})^{(k_{r}-k_{3}+3)/2}.$ (19)
+We need to explain why $\mu_{eff}$ is much smaller than $\Lambda_{SUSY}$ and of order the electroweak scale or soft SUSY breaking scale $m_{soft}$.
+We will show that, the smallness of $\mu_{eff}$ is the consequence of large suppression factors, either originate from some power of deviation parameter that characterizes the deviation of modulus VEV from typical fixed point, or from non-trivial power (determined by the choice of modular weights for $H_{u}$ and $H_{d}$) of a small factor related to the modulus VEV.
+All these suppression factors are determined by the modulus stabilization mechanism, with the stabilized modulus VEVs being affected by SUSY breaking effects.
+4 Stabilization of modulus field with nilpotent Goldstino ${\textbf{X}}_{nl}$ superfield
+To stabilize the modulus field, we can utilize effective supergravity scalar potentials derived from a weight $-3$ superpotential to determine the minimum.
+The modulus VEV can always be restricted to the fundamental domain
+${D}_{H}\equiv\left\{\tau\in\mathcal{H}:-\frac{1}{2}\leq\Re{\tau}<\frac{1}{2},\,|\tau|>1\right\}\cup\left\{\tau\in\mathcal{H}:-\frac{1}{2}<\Re\tau\leq 0,\,|\tau|=1\right\}\,,$ (20)
+of the modular group, as any choice of the modulus VEV in the upper-half plane can be uniquely related to a $\tau\in D_{H}$ in the fundamental domain via a modular transformation, which is physically equivalent to such a choice.
+Discussions on modulus stabilization with single modulus and multiple moduli can be found in Kobayashi:2019xvz ; Kobayashi:2020hoc ; Kobayashi:2020uaj ; Ishiguro:2020tmo ; NPP:2201.02020 ; PLNSR:2304.14437 ; King:2023snq ; King:2024ssx ; Ding:2024neh ; Higaki:2024pql ; Higaki:2024jdk ; Kobayashi:2023spx ; Funakoshi:2024yxg .
+To ensure the modular invariance and Kahler transformation invariance of
 $e^{G}\equiv e^{K}|W|^{2}~{},$ (21)
-the superpotential for modulus should take the following form
-$W(\tau)=c_{1}^{3}\frac{H(\tau)}{\eta(\tau)^{6}}.$ (22)
-for the Kahler potential of the form
+the superpotential for the modulus should take the form
+$W(\tau)=c_{1}^{3}\frac{H(\tau)}{\eta(\tau)^{6}},$ (22)
+with the Kahler potential
 $K_{mod}=-3\log(2\Im\tau)~{},$ (23)
-in the unit $M_{Pl}=1$ and $c_{1}M_{Pl}$ denotes the typical scale for the superpotential for dimensionless number $c_{1}$.
+in the unit $M_{Pl}=1$. Here, $c_{1}M_{Pl}$ denotes the typical mass scale for the superpotential, where $c_{1}$ is a dimensionless number.
 The modular invariant holomorphic function $H(\tau)$ is given by
@@
 $\equiv$ $\left(\frac{G_{6}(\tau)}{\eta(\tau)^{12}}\right)^{m}\left(\frac{G_{4}(\tau)}{\eta(\tau)^{12}}\right)^{n}{\cal P}\left(j(\tau)\right)~{},$
-where ${\cal P}$ denotes a polynomial with respect to $j(\tau)$. The Klein j-function is defined as a modular form of zero weight
+where ${\cal P}\left(j(\tau)\right)$ denotes a polynomial with respect to $j(\tau)$. The Klein $j$-function is defined as a modular form of zero weight
 $j(\tau)$$\equiv$ $\frac{3^{6}5^{3}}{\pi^{12}}\frac{G_{4}(\tau)^{3}}{\eta(\tau)^{24}},$ (25)
-$\approx$ $744+q^{-1}+196884q+21493760q^{2}+864299970q^{3}+{\cal O}(q^{4}),$
+$\approx$ $744+q^{-1}+196884q+21493760q^{2}+864299970q^{3}+{\cal O}(q^{4})~{},$
 with $j(i\infty)$ diverges, $j(\omega)=0$ and $j(i)=1728$.
-We should note that, in such a stabilization mechanism, fixed points $z=i,\omega\equiv e^{i2\pi/3}$ are always the extrema of the scalar potential, although may not be the global minimum.
-With proper SUSY breaking contributions, possibly from the ”matter”-like sector, the stabilized modulus value, which located at/near the fixed points $z_{FP}=\omega,i$, can slightly deviate from the original global minimum and possibly uplift the vacuum to dS (de Sitter) one.
-In fact, the classical 10d supergravity Maldacena-Nunez no-go result MN suggests that any constructions that lead to dS vacuum must include non-classical ingredients as D-branes, O-planes, or quantum corrections.
-To describe the scenario in which SUSY is linearly realized in a gravitational sector and non-linearly realized in the uplifting SUSY breaking sector, (for example, from anti-D branes) the formalism of non-linearly realized SUSY as formulated by KS can be used for such a system.
-It was proved in Burgess:2021juk with the non-linear SUSY realization scheme that the supergravity form is stable even if the supermultiplets in SUSY breaking sector are split badly enough to allow some of its members to be integrated out while the others are not.
+We should note that, in such a stabilization mechanism, the fixed points $z=i,\omega\equiv e^{i2\pi/3}$ are always the extrema of the scalar potential, although they may not be the global minimum.
+With proper SUSY breaking contributions, possibly from the ”matter”-like sector, the stabilized modulus value, which located at/near the fixed points $z_{FP}=\omega,i$, can slightly deviate from the original global minimum.
+The minimum of the scalar potential can also possibly be uplifted to dS (de Sitter) type.
+In fact, the classical 10d supergravity Maldacena-Nunez no-go result MN suggests that any constructions leading to a dS vacuum must include non-classical ingredients, such as D-branes, O-planes or quantum corrections.
+To describe the scenario in which SUSY is linearly realized in a gravitational sector and non-linearly realized in the uplifting SUSY breaking sector (for example, from anti-D branes), the formalism of non-linearly realized SUSY as formulated by KS can be used for such a system.
+It was proved in Burgess:2021juk that, with the non-linear SUSY realization scheme, the supergravity form is stable even if the supermultiplets in SUSY breaking sector are split badly enough to allow some of its members to be integrated out while the others are not.
 When SUSY is spontaneously broken in supergravity, the Goldstino is eaten by the gravitino via the super-Higgs mechanism.
-When such a breaking happens at low energies compared with the Planck mass, the goldstino couplings can be described by introducing a (constrained) independent superfield in the supergravity effective action.
+When such a breaking happens at low energies compared with the Planck scale, the Goldstino couplings can be conveniently described by a (constrained) independent superfield in the supergravity effective action.
 The Goldstino provides a way for an arbitrary non-supersymmetric low-energy theory to realize SUSY non-linearly.
 The emergence of the constraints may also be viewed as an effective field theory limit in which certain mass parameters become very large and the corresponding modes decouple.
-Coupling of non-linearly realized SUSY with nilpotent goldstino $X_{nl}^{2}=0$ can be described by
-${\cal L}\supseteq\int d^{4}\theta X_{nl}^{*}X_{nl}+\left[\int d^{2}\theta f_{0}X_{nl}+h.c.\right]~{},$ (26)
-with the chiral superfield
-$X_{nl}=\frac{\psi_{X}\psi_{X}}{2F_{X}}+\sqrt{2}\theta\psi_{X}+\theta^{2}F_{X}~{},$ (27)
-is the IR limit of certain microscopic superfield.
-As noted in Aparicio:2015psl , for the anti-D3 brane in the KKLT scenario, the representation in terms of $X$ is very convenient and enough to capture all the anti-D3 brane degrees of freedoms since it allows to treat the effect of anti-D3 brane in terms of standard supergravity couplings of matter and moduli superfields to the nilpotent goldstino.
-From the perspective of low energy effective theory, regardless of the UV origin of SUSY breaking, such a non-linear realization scheme is also convenient for modulus stablization and vacuum uplifting discussions.
-When coupled to supergravity, the most general form is given by
+The coupling of non-linearly realized SUSY with a nilpotent chiral Goldstino superfield, subjected to constraints ${\textbf{X}}_{nl}^{2}=0$, can be described by
+${\cal L}\supseteq\int d^{4}\theta\textbf{X}_{nl}^{\dagger}\textbf{X}_{nl}+\left[\int d^{2}\theta f_{0}\textbf{X}_{nl}+h.c.\right]~{}.$ (26)
+The nilpotent constraints for the chiral Goldstino superfield can be solved to be
+$\textbf{X}_{nl}=\frac{\psi_{X}\psi_{X}}{2F_{X}}+\sqrt{2}\theta\psi_{X}+\theta^{2}F_{X}~{},$ (27)
+where all the fields are the functions of $y^{\mu}\equiv x^{\mu}+i\theta\sigma^{\mu}\bar{\theta}$.
+The lowest component of $\textbf{X}_{nl}$ is eliminated by the nilpotent constraints $\textbf{X}_{nl}^{2}=0$ and is no longer an independent complex scalar field.
+This leaves only the Goldstino field $\psi_{X}$ and an independent auxiliary field $F_{X}$ in $\textbf{X}_{nl}$ superfield.
+The Goldstino field $\psi_{X}$, which is the fermionic component of $\textbf{X}_{nl}$, eventually mixes with the gravitino to give it a mass through the super-Higgs mechanism.
+The lowest component of $\textbf{X}_{nl}$, which is denoted as ${X}_{nl}$ for the Goldstino bilinear ${\psi_{X}\psi_{X}}$, obviously satisfies ${X}_{nl}^{2}=0$ due to its anti-commuting Grassman variable nature.
+The chiral superfield $\textbf{X}_{nl}$ can be seen as the long-distance expression of the microscopic superfield X that satisfies $\bar{D}^{\dot{\alpha}}{\cal J}_{\alpha\dot{\alpha}}=D_{\alpha}\textbf{X}$ for Ferrara-Zumino multiplet ${\cal J}_{\alpha\dot{\alpha}}$ KS , which can be used even when SUSY is spontaneously broken and can be followed consistently along the renormalization group flow down to the deep IR.
+As noted in Aparicio:2015psl , for the anti-D3 brane in the KKLT scenario, the representation in terms of $\textbf{X}_{nl}$ is very convenient and enough to capture all the anti-D3 brane degrees of freedoms since it allows to treat the effect of anti-D3 brane in terms of standard supergravity couplings of matter and moduli superfields to the nilpotent Goldstino superfield.
+From the perspective of low energy effective field theory, regardless of the UV origin of SUSY breaking, such a non-linear realization scheme is also convenient for discussions on modulus stabilization and vacuum uplifting.
+When coupled to supergravity, the most general form (at the two-derivative level) for the action in superspace is given by
 […]
 with
-$K={X}^{*}_{nl}X_{nl}~{},~{}~{}~{}W=f_{0}X+W_{0}~{}.$ (29)
-and $f_{0}=-F_{X}$.
+$K=\textbf{X}_{nl}^{\dagger}\textbf{X}_{nl}~{},~{}~{}~{}W=f_{0}\textbf{X}_{nl}+c_{1}^{3}~{},$ (29)
+where without loss of generality we take $f_{0}$ and $c_{1}$ to be real.
+Integrating out the auxillary field $F_{X}$ gives $F_{X}=-(W_{\textbf{X}_{nl}})^{*}=-f_{0}$.
 To accommodate the modular flavor symmetry, the Kahler potential and superpotential needs to be adapted to
-$K$ $\supseteq$ $-3\log(2\Im\tau)+\frac{{X}^{*}_{nl}X_{nl}}{(2\Im\tau)^{k_{X}}}~{},$
-$W$ $\supseteq$ $\left(f_{0}X_{nl}\left[\eta(\tau)\right]^{2k_{X}}+c_{1}^{3}\right)\frac{H(\tau)}{\eta^{6}(\tau)}.$ (30)
-The scalar potential for $X$ and $\tau$ can be obtained from
+$K$ $\supseteq$ $-3\log(2\Im\tau)+\frac{\textbf{X}^{\dagger}_{nl}\textbf{X}_{nl}}{(2\Im\tau)^{k_{X}}}~{},$
+$W$ $\supseteq$ $\left(f_{0}\textbf{X}_{nl}\left[\eta(\tau)\right]^{2k_{X}}+c_{1}^{3}\right)\frac{H(\tau)}{\eta^{6}(\tau)}.$ (30)
+The supergravity scalar potential can be derived from
 $V=e^{K}\left[K^{\alpha\bar{\beta}}\nabla_{\alpha}W\nabla_{\bar{\beta}}W-3W\overline{W}\right],$ (31)
-with
+where the Kahler-covariant derivative of the superpotential is
 $\nabla_{\alpha}W=W_{\alpha}+K_{\alpha}W~{}.$ (32)
 From the Kahler potential, we can calculate the inverse Kahler metric
-$K^{\bar{\alpha}\beta}$ $=$ $\left(\begin{array}[]{cc}\frac{(2\Im\tau)^{2}}{3}\left[1-\frac{k_{X}}{3}\frac{|X|^{2}}{(2\Im\tau)^{k_{X}}}\right]&\frac{i(2\Im\tau)k_{X}X}{3}\\ -\frac{i(2\Im\tau)k_{X}{X}^{*}}{3}&(2\Im\tau)^{k_{X}}\left[1+\frac{k_{X}^{2}|X|^{2}}{3(2\Im\tau)^{k_{X}}}\right]\end{array}\right)~{}.$ (35)
-Here the indices $\alpha\sim(\tau,X_{nl})$ and the nilpotency of $X_{nl}$ is used in the calculation.
-The scalar potential can be calculated to be
+$K^{{\alpha}\bar{\beta}}$ $=$ $\left(\begin{array}[]{cc}\frac{(2\Im\tau)^{2}}{3}\left[1-\frac{k_{X}}{3}\frac{\textbf{X}^{\dagger}_{nl}\textbf{X}_{nl}}{(2\Im\tau)^{k_{X}}}\right]&\frac{i(2\Im\tau)k_{X}\textbf{X}_{nl}}{3}\\ -\frac{i(2\Im\tau)k_{X}\textbf{X}^{\dagger}_{nl}}{3}&(2\Im\tau)^{k_{X}}\left[1+\frac{k_{X}^{2}(\textbf{X}^{\dagger}_{nl}\textbf{X}_{nl})}{3(2\Im\tau)^{k_{X}}}\right]\end{array}\right)~{}.$ (35)
+Here the indices $\alpha,\beta\sim(\tau,\textbf{X}_{nl})$ and the nilpotency of $\textbf{X}_{nl}$ is used in the calculation.
+The scalar potential for the modulus field $\tau$ and $X_{nl}$ can be calculated as
 $V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})$ (36)
@@
 -- $\left.3\left|\tilde{c}_{W}\right|^{2}\left|{H(\tau)}\right|^{2}\right\}~{},$
-with
+where
 $\tilde{W}_{M}$$\equiv$ $H^{\prime}(\tau)-\frac{3iH(\tau)}{2\pi}\hat{G}_{2}(\tau,\tau^{*})~{},$
@@
 $\tilde{W}_{X}$ $=$ $f_{0}\eta(\tau)^{2k_{X}}\left(1+\frac{|X_{nl}|^{2}}{2(\Im\tau)^{k_{X}}}\right)+\frac{{X}^{*}_{nl}c_{1}^{3}}{(2\Im\tau)^{k_{X}}}.$ (37)
-and the constraint $X_{nl}^{2}=0$ is used to eliminate some of the high order $X_{nl}$ terms.
-We also use the expression for the derivative of Dedekind $\eta$ function, which is given by
+The constraint $X_{nl}^{2}=0$ for the lowest component of $\textbf{X}_{nl}$ is used to eliminate some of the high order $X_{nl}$ terms.
+We also use the following expression for the derivative of Dedekind $\eta$ function
 $\frac{\eta^{\prime}(\tau)}{\eta(\tau)}=\frac{i}{4\pi}G_{2}~{},$ (38)
@@
 $\hat{G}_{2}(\tau,{\tau}^{*})=G_{2}(\tau)-\frac{\pi}{\Im\tau}~{}.$ (40)
-The generic complicate form of the scalar potential may obscure the discussions of the stabilization mechanism. So, we would like to concentrate on its typical simplification forms
+The generic complicate form of the scalar potential may obscure the discussions on the stabilization mechanism. Therefore, we would like to concentrate on its typical simplification forms
   • Scenario with $k_{X}=0$:
-    The scalar potential can be simplified to be
+    When $k_{X}=0$, the scalar potential simplifies significantly.
+    The potential can then be written as
     $V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})$ (41)
@@
     $+$ $\left|H(\tau)\right|^{2}\left(f^{2}_{0}+2f_{0}^{2}{|X_{nl}|^{2}}+2\Re\left[f_{0}c_{1}^{3}{X}_{nl}^{*}\right]+|X_{nl}|^{2}c_{1}^{6}\frac{}{}\right)~{}$
-    -- $\left.3\left(f_{0}^{2}|X_{nl}|^{2}+c_{1}^{6}+2\Re\left[f_{0}X_{nl}c_{1}^{3}\right]\right)\left|{H(\tau)}\right|^{2}\right\}~{}$
-    $+$ $\frac{|X_{nl}|^{2}}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}c_{1}^{3}\right|^{2}+\left|H(\tau)\right|^{2}f^{2}_{0}-3c_{1}^{6}\left|{H(\tau)}\right|^{2}\right\}~{},$
-    with
+    -- $\left.3\left|{H(\tau)}\right|^{2}\left(f_{0}^{2}|X_{nl}|^{2}+c_{1}^{6}+2\Re\left[f_{0}X_{nl}c_{1}^{3}\right]\right)\right\}~{}$
+    $+$ $\frac{|X_{nl}|^{2}}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}c_{1}^{6}+\left|H(\tau)\right|^{2}f^{2}_{0}-3c_{1}^{6}\left|{H(\tau)}\right|^{2}\right\}~{},$
+    where
     $\tilde{W}_{M}$$\equiv$ $H^{\prime}(\tau)-\frac{3iH(\tau)}{2\pi}\hat{G}_{2}(\tau,\tau^{*})~{},$
-    $\tilde{c}_{M}$ $=$ $f_{0}X_{nl}+c_{1}^{3}~{},$ (42)
-    Note the constraints $X_{nl}^{2}=\bar{X}_{nl}^{2}=0$ are imposed for the scalar potential.
-    The modulus VEV can be determined by the extrema conditions of the scalar potential
+    $\tilde{c}_{M}$ $=$ $f_{0}X_{nl}+c_{1}^{3}~{}.$ (42)
+    Note the constraints $X_{nl}^{2}=({X}^{*}_{nl})^{2}=0$ are used in the scalar potential.
+    In this scenario, the modulus VEV can be determined by the extreme conditions of the scalar potential
     $\frac{\partial}{\partial\tau}V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})=\frac{\partial}{\partial\tau^{*}}V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})=0~{},$
-    and the minimum condition can be determined by positive definiteness of Hessian matrix in terms of
+    and the minimum condition can be determined by the positive definiteness of Hessian matrix in terms of
     $\frac{\partial^{2}}{\partial x^{2}}V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})$ $=$ $2\left(\frac{\partial^{2}}{\partial\tau\partial\tau^{*}}+\Re\frac{\partial^{2}}{\partial\tau\partial\tau}\right)V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*}),$
@@
     $\frac{\partial^{2}}{\partial x\partial y}V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})$ $=$ $-2\Im\frac{\partial^{2}}{\partial\tau\partial\tau}V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*}),$ (43)
-    with $\tau\equiv x+iy\equiv\rho_{\tau}e^{i\theta_{\tau}}$, as well as those conditions for $X_{nl}$.
-It was argued in target:space ; Ding:2024neh that, modular invariance and $\tau\leftrightarrow-\tau^{*}$ symmetry of the scalar potential ensures that the first derivative of scalar potential along certain directions at the boundary of the fundamental domain vanishes
+    where $\tau\equiv x+iy\equiv\rho_{\tau}e^{i\theta_{\tau}}$.
+It was argued in target:space ; Ding:2024neh ; Higaki:2024pql that the modular invariance and $\tau\leftrightarrow-\tau^{*}$ symmetry of the scalar potential ensure that the first derivative of scalar potential along certain directions at the boundary of the fundamental domain vanishes
     $\left.\frac{\partial}{\partial x}V\right|_{x=\pm 1/2,0}=0~{},~{}~{}~{}\left.\frac{\partial}{\partial\rho_{\tau}}V\right|_{\rho_{\tau}=1}=0~{}.~{}~{}~{}$ (44)
-    Then we will determine the extrema condition
+    We then determine the extreme conditions
     $\frac{\partial}{\partial\rho}V=\frac{\partial}{\partial\theta}V=0~{},$ (45)
-    for $X_{nl}\equiv\rho e^{i\theta}$. It can be calculated that the extreme conditions for $X_{nl}$ reduce to the constraints
-    $\frac{\partial}{\partial\theta}V=0~{},\Rightarrow\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}-2\left|H(\tau)\right|^{2}=0~{},$ (46)
-    and
-    $\frac{\partial}{\partial\rho}V=0~{},$$\Rightarrow$ $\left|H(\tau)\right|^{2}f_{0}^{2}=0~{},$ (49)
-    $\Rightarrow$ $\left\{\begin{array}[]{l}\left|H(\tau)\right|^{2}=0,~{}~{}~{}~{}~{}~{}{\rm for}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}f_{0}\neq 0~{}.\\ ~{}~{}~{}~{}~{}~{}~{}f_{0}=0~{},~{}~{}~{}~{}{\rm and/or}~{}~{}~{}\left|H(\tau)\right|^{2}\neq 0.\end{array}\right.$
-    So, we have the minimum conditions for $X_{nl}$
-    $\left|\tilde{W}_{M}\right|^{2}=\left|H(\tau)\right|^{2}=0~{},$ (50)
-    in the case $f_{0}\neq 0$.
+    for $X_{nl}\equiv\rho e^{i\theta}$. We have the following discussions for some non-generic simplified scenarios:
+    • Case $\rho=0$:
+      This corresponds to $X_{nl}=0$, reducing to the case with only a single $\tau$ modulus field, which has already been extensively discussed target:space ; NPP:2201.02020 .
+    • Case $f_{0}=0$:
+      The extrema of the scalar potential are located at the parameter points with $X_{nl}=0$.
+Again, the modulus VEV of the corresponding extrema is the same as the case with only a single $\tau$ modulus field.
+    • Case $c_{1}=0$:
+      The extrema of the scalar potential are also located at the parameter point with $X_{nl}=0$, and the corresponding potential energy is
+      $\left.V\right|_{X_{nl}=0}=\frac{1}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left|{H(\tau)}\right|^{2}f_{0}^{2}\geq 0~{},$ (46)
+      for values of $\tau$ within the fundamental region.
+To obtain a vanishing (very tiny positive) vacuum energy $\Lambda^{4}\sim 10^{-120}M_{Pl}^{4}\approx 0$, we can find the parameter points that are solutions of ${H(\tau)}=0$ and also satisfy the minimum condition (with respect to $\tau$ and $\tau^{*}$) with $\tilde{W}_{M}=0$.
+It can be checked that solutions of ${H(\tau)}=0$ can also satisfy the minimum condition $\tilde{W}_{M}=0$, which will be clear shortly.
+Therefore, the stabilized modulus VEV is located at the solutions of ${H(\tau)}=0$ with (almost) vanishing vacuum energy in this case.
+    We now focus on the general case with $f_{0},c_{1},X_{nl}\neq 0$. It can be calculated that the extreme condition for $\theta$ is
+    $\frac{\partial}{\partial\theta}V=-\frac{2f_{0}c_{1}^{3}}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\rho\sin\theta\left[\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}-2\left|H(\tau)\right|^{2}\right]=0.$ (47)
+    First, we should note that, as $[(2\Im\tau)|\eta(\tau)|^{4}]^{-1}$ diverges when $\Im\tau=\infty$, the $\tau=i\infty$ point is not a solution of the extreme condition for $\theta$.
+    For generic $\theta$ (with $\theta\neq 0,\pi$ to render $X_{nl}$ not real) in the case $f_{0},c_{1},\rho\neq 0$, the extreme condition for $\theta$ reduces to the constraint
+    $\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}-2\left|H(\tau)\right|^{2}=0~{}.$ (48)
+    After substituting this constraint in the extreme condition for $\rho$, we have
+    $\frac{\partial}{\partial\rho}V=4\rho f_{0}^{2}\frac{\left|H(\tau)\right|^{2}}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}=0~{}.$ (49)
+    Such an extreme condition for $\rho$ can be satisfied with
+    $\left|H(\tau)\right|^{2}=0~{}.$ (50)
+    Thus, we have the extreme conditions for the case with $f_{0},c_{1}\neq 0$ and a non-vanishing non-real $X_{nl}$
+    $\left|\tilde{W}_{M}\right|^{2}=\left|H(\tau)\right|^{2}=0~{}.$ (51)
     Substituting back into the scalar potential, we can see that the potential energy at the minimum vanishes.
-So, the presence non-trivial $X_{nl}$ sector, whose contributions can uplift the scalar potential, can naturally lead to a flat Minkowski universe.
-    We will concentrate on the constraint $H(\tau)=0$ when $f_{0}\neq 0$.
-When $f_{0}=0$, the scalar potential for modulus fields will not be affected by $X_{nl}$ as the two sector decouples in the $k_{X}=0$ limit.
-    From the expression of $H(\tau)$, the constraint $H(\tau)=0$ requires either the condition $j(\tau)^{n/3}[j(\tau)-1728]^{m/2}=0$ or ${\cal P}(j(\tau))=0$.
+Therefore, the presence of a non-trivial $X_{nl}$ sector, whose contributions can uplift the scalar potential, can naturally lead to a flat Minkowski universe in this case.
+    From the expression of $H(\tau)$, the constraint $H(\tau)=0$ requires
+    $j(\tau)^{n/3}[j(\tau)-1728]^{m/2}=0,~{}~{}{\rm or}~{}~{}~{}~{}{\cal P}(j(\tau))=0.$ (52)
     Obviously, $\tau=i,\omega$ are its solutions for $m\geq 1$ and $n\geq 1$, respectively.
-As ${\cal P}(j(\tau))$ is an arbitrary polynomial function of $j(\tau)$, from the fundamental theorem of algebra, arbitrary complex value of $j(\tau)$ can be the root of typical chosen polynomial equation ${\cal P}(j(\tau))=0$, leading to no interesting constraint at all.
-So, we would like to focus on the constraint
+As ${\cal P}(j(\tau))$ is an arbitrary polynomial function of $j(\tau)$, from the fundamental theorem of algebra, any complex value of $j(\tau)$ can be the root of typical chosen polynomial equation ${\cal P}(j(\tau))=0$, leading to no interesting constraint at all unless we specify the concrete form of ${\cal P}(j(\tau))$.
+    Then, we would like to focus on the constraint
     $\tilde{W}_{M}\equiv H^{\prime}(\tau)-\frac{3iH(\tau)}{2\pi}\hat{G}_{2}(\tau,\tau^{*})=0~{},$ (53)
-    that satisfies $H(\tau)=0$. Such a constraint reduces to
+    which satisfies $H(\tau)=0$. Such a constraint reduces to
     $H^{\prime}(\tau)=\left\{j(\tau)^{n/3}[j(\tau)-1728]^{m/2}{\cal P}(j(\tau))\right\}^{\prime}=0~{},$ (54)
@@
     $\left.+j(\tau)^{n/3}[j(\tau)-1728]^{m/2}\frac{\partial{\cal P}(j(\tau))}{\partial j(\tau)}\right]=0~{},$
-    $\Rightarrow$ $H(\tau)j^{\prime}(\tau)\left[\frac{n}{3}\frac{1}{j(\tau)}+\frac{m}{2}\frac{1}{j(\tau)-1728}+\frac{1}{{\cal P}(j(\tau))}\frac{\partial{\cal P}(j(\tau))}{\partial j(\tau)}\right]=0~{},$
-    So $H^{\prime}(\tau)=0$ can trivially hold for $H(\tau)=0$.
+    $\Rightarrow$ $H(\tau)j^{\prime}(\tau)\left[\frac{n}{3}\frac{1}{j(\tau)}+\frac{m}{2}\frac{1}{j(\tau)-1728}+\frac{1}{{\cal P}(j(\tau))}\frac{\partial{\cal P}(j(\tau))}{\partial j(\tau)}\right]=0~{}.$
+    So $H^{\prime}(\tau)=0$ can trivially hold when $H(\tau)=0$.
     On the other hand, $H^{\prime}(\tau)=0$ can also be satisfied by the other two conditions
     $\left[\frac{n}{3}\frac{1}{j(\tau)}+\frac{m}{2}\frac{1}{j(\tau)-1728}+\frac{1}{{\cal P}(j(\tau))}\frac{\partial{\cal P}(j(\tau))}{\partial j(\tau)}\right]=0~{},$ (55)
-    and $j^{\prime}(\tau)=0$.
+    or $j^{\prime}(\tau)=0$.
     We have the following discussions for the condition (55)
-    • $n\neq 0$ and $m=0$ ($m\neq 0$ and $n=0$) or $mn\neq 0$:
-      The condition (53) reduces to
+    • $m+n\neq 0$:
+      When $n\neq 0$ and $m=0$, condition (55) reduces to
       $\frac{d{\cal P}(j(\tau))}{{\cal P}(j(\tau))}=-\frac{n}{3}\frac{dj(\tau)}{j(\tau)}~{},$ (56)
@@
       ${\cal P}(j(\tau))=C_{0}\left[j(\tau)\right]^{-n/3}~{}.$ (57)
-      So, unless $n\in 3\mathbb{Z}$ is negative, ${\cal P}(j(\tau))$ is not a polynomial of $j(\tau)$.
-Similarly, we can deduce that $m\in 2\mathbb{Z}$ should be negative.
-Such a deduction can also hold for $m\neq 0$ with $n\neq 0$ case, which also requires $m=n=0$.
+      Unless $n\in 3\mathbb{Z}$ is negative, ${\cal P}(j(\tau))$ is not a polynomial of $j(\tau)$.
+Similarly, when $m\neq 0$ and $n=0$, we can deduce that $m\in 2\mathbb{Z}$ should be negative.
+For $m\neq 0$ and $n\neq 0$, similar deduction can give
+      ${\cal P}(j(\tau))=C_{0}^{\prime}\left[j(\tau)\right]^{-n/3}\left[{j(\tau)-1728}\right]^{-m/2}~{},$ (58)
+      which also requires both $m\in 2\mathbb{Z}$ and $n\in 3\mathbb{Z}$ being negative to guarantee ${\cal P}(j(\tau))$ being a polynomial of $j(\tau)$.
 As $m,n$ are defined to be positive integers, this condition requires $m=n=0$.
     • $m=0$ and $n=0$:
-      With the constraints that $m=n=0$, the condition (53) reduces to
+      With the constraints $m=n=0$, the condition (55) reduces to
       $\frac{\partial{\cal P}(j(\tau))}{\partial j(\tau)}=0~{},$ (59)
@@
       ${\cal P}(j(\tau))={\rm Constant}.$ (60)
-      as a constant. As $H(\tau)=0$, the constant should be zero.
-    So, the condition (53) requires $H(\tau)$ to be zero. As the constraint (53) is trivial, we would like to discuss the condition $j^{\prime}(\tau)=0$. From the derivative of the Klein j-function
+      As $H(\tau)=0$, this constant should be zero.
+    Therefore, the condition (55) requires a vanishing $H(\tau)$. For the condition $j^{\prime}(\tau)=0$, from the derivative of the Klein $j$-function
     $\frac{\partial}{\partial\tau}j(\tau)=-2\pi i\frac{{E_{6}(\tau)}{E_{4}^{2}(\tau)}}{[\eta(\tau)]^{24}}=0~{},$ (61)
-    we can see that its solutions are $\omega$ and $i$.
-    So, we can conclude that, in the scenario with $k_{X}=0$ for $f_{0}\neq 0$, the extrema of the scalar potential can only be determined by $H(\tau)=0$ (and the extrema conditions from $X_{nl}$ are also satisfied by $z_{0}=\omega,i$).
-The corresponding vacuum is Minkowski type.
-As noted previously, the constraint $H(\tau)=0$ predicts that the fixed points $z_{0}=\omega,i$ are the local extrema.
+    we can see that its solutions are $\omega$ and $i$, leading to no new additional constraints. Therefore, the extreme conditions (51) can reduce to the requirement of vanishing $H(\tau)$.
+    Before we conclude this scenario, it should be noted that $\theta=0,\pi$, which corresponds to real $X_{nl}=\pm\rho$ with positive $\rho$, are also solutions of the the extreme condition (47) for $\theta$.
+In this case, the scalar potential reduces to
+    $V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})$ (62)
+    $=$ $\frac{1}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}\left(f_{0}^{2}\rho^{2}+c_{1}^{6}\pm 2f_{0}\rho c_{1}^{3}\right)\right.$
+    $+$ $\left|H(\tau)\right|^{2}\left(f^{2}_{0}+2f_{0}^{2}\rho^{2}\pm 2f_{0}c_{1}^{3}\rho+\rho^{2}c_{1}^{6}\frac{}{}\right)~{}$
+    -- $\left.3\left|{H(\tau)}\right|^{2}\left(f_{0}^{2}\rho^{2}+c_{1}^{6}\pm 2f_{0}\rho c_{1}^{3}\right)\right\}~{}$
+    $+$ $\frac{\rho^{2}}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}c_{1}^{6}+\left|H(\tau)\right|^{2}f^{2}_{0}-3c_{1}^{6}\left|{H(\tau)}\right|^{2}\right\}~{}.$
+    The scalar potential, which is a quadratic polynomial of $\rho$, can be written as
+    $V=\frac{1}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left[C_{1}\rho^{2}\pm C_{2}\rho+C_{3}\right]~{},$ (63)
+    where
+    $C_{1}$ $=$ $\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}(f_{0}^{2}+c_{1}^{6})-2\left|H(\tau)\right|^{2}c_{1}^{6}~{},$
+    $C_{2}$ $=$ $\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}2f_{0}c_{1}^{3}-4\left|H(\tau)\right|^{2}f_{0}c_{1}^{3}~{},$
+    $C_{3}$ $=$ $\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}c_{1}^{6}+\left|H(\tau)\right|^{2}\left(f_{0}^{2}-3c_{1}^{6}\right)~{}.$ (64)
+    The scalar potential has a minimum with the corresponding $\rho$ parameter
+    $\rho_{0}=\mp\frac{C_{2}}{2C_{1}}~{},$ (65)
+    and the potential energy
+    $\left.V\right|_{\rho=\rho_{0}}=C_{3}-\frac{C_{2}^{2}}{2C_{1}}~{}.$ (66)
+    The modulus VEV is determined by the minimum conditions for the potential (66) with respect to $\tau,\tau^{*}$, and the condition $C_{2}^{2}=2C_{1}C_{3}$ should be established to obtain a vanishing (almost vanishing tiny positive) vacuum energy.
+It can be checked from previous discussions that solutions of ${H(\tau)}=0$ can also satisfy such minimum conditions.
+    Thus, we can conclude that, in the scenario with $k_{X}=0$ for $f_{0}\neq 0$, the extrema of the scalar potential are located at the points determined by $H(\tau)=0$, and the corresponding vacuum is of Minkowski type.
+As noted previously, the constraint $H(\tau)=0$ predicts that the fixed points $z_{0}=\omega,i$ are always the local extrema unless $m=n=0$.
 Other solutions of $\tau$ from ${\cal P}(j(\tau))=0$ are essentially possible because of the arbitrariness of the polynomial form of ${\cal P}(j(\tau))$.
-So, other minimum, which is determined by (44) to locate at the boundary of the fundamental domain or at the imaginary axis, is also possible.
+Therefore, other minima, which are determined by (44) to lie at the boundary of the fundamental domain or at the imaginary axis, can also be possible.
     In fig.1, we show the energies of the scalar potential at modulus values taken within the fundamental domain with various parameter choices.
-It can be seen that the energies are non-negative when $f_{0}\neq 0$.
+It can be seen that the potential energies are non-negative when $f_{0}\neq 0$.
     Figure 1: The values of $V(X_{nl},\bar{X}_{nl},\tau,{\tau}^{*})$ within the fundamental domain in the scenario with $k_{X}=0$ for various parameter choices when $f_{0}\neq 0$.
+We adopt $\mathcal{P}(j)=1$ for the left panel and $\mathcal{P}(j)=j-4.5764069879884376\times 10^{-11}$ for the right panel.
   • Scenario with small $X_{nl}$ (when $|X_{nl}|\rightarrow 0$):
-    In this case, at small $|X_{nl}|$ region, the $|X_{nl}|^{2}$ terms can be neglected in addition to the . constraint $X_{nl}^{2}=0$. The scalar potential can be simplified to
+    In this case, at small $|X_{nl}|$ region, the $|X_{nl}|^{2}$ terms can be neglected in addition to the constraint $X_{nl}^{2}=0$.
+Note that the $|X_{nl}|^{2}$ terms will dominate the scalar potential at very large $|X_{nl}|$ values, ensuring that the potential can be bounded from below.
+    The scalar potential can be simplified to
     $V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})$ (67)
@@
     It can be seen easily from the linear terms of $X_{nl}$ that the extrema conditions for $X_{nl}$ will not lead to trivial constraints for $\tau$ with arbitrary choices of $k_{X}$.
-So, other modulus values lying on the boundary of fundamental domain or imaginary axis other than the fixed point can possibly be the minimum.
-The (otherwise stabilized) modulus VEV (without contributions from $X_{nl}$ sector) can be destabilized by non-vanishing $X_{nl}$.
-Note that the $|X_{nl}|^{2}$ terms will dominate the scalar potential at very large $|X_{nl}|$ values, ensuring that the potential can be bounded from below.
+Therefore, other modulus values lying on the boundary of fundamental domain or imaginary axis other than the fixed point, can possibly be the minimum.
+The initially stabilized modulus VEV without contributions from $X_{nl}$ sector can then be destabilized by new contributions from non-vanishing $X_{nl}$.
 We will discuss the extrema conditions for the most general case in our subsequent studies.
-5 Re-generation of naturally small $\mu_{eff}$ after modulus stabilization
+It should be noted that the modulus stabilization mechanism with nilpotent $\textbf{X}_{nl}$ superfield can be fairly generic and ubiquitous, which is independent on the following discussions on the regeneration of small $\mu_{eff}$ term following modulus stabilization.
+5 Regeneration of naturally small $\mu_{eff}$ following modulus stabilization
+[…]
 It is known that the fixed points $\tau_{FP}=i,\omega,i\infty$ for modular theories based on the $\Gamma_{N}$ invariance can preserve $Z_{2}^{S},Z_{3}^{ST},Z_{N}^{T}$ residual symmetries, respectively Novichkov:2021evw .
-One can easily seen from the modular invariance properties of scalar potential that the self-dual fixed points should be the local extrema of the scalar potential, since they act non-trivially (with residual symmetries) on the derivatives of the potential and must vanish.
-Although they may not be the global minima, experiences from modulus stabilization in string theory indicate that the minima should always lie at/near $\tau_{FP}=i,\omega\equiv e^{i2\pi/3}$ for the tree-level scalar potential, as the scalar potential in general diverges for $\tau\rightarrow i\infty$.
-It was found in Novichkov:2021evw that the hierarchies among the charged leptons and quarks mass can follow solely from the properties of the modular forms in the vicinity of symmetric points, avoiding the fine-tuning of the constant parameters (or without the need to introducing extra weighton-like scalar fields).
-Similar ideas, augmented with additional suppression powers of $q$ (or $(2\Im\tau)^{-1}$), can be used to generate the hierarchy between $\Lambda_{SUSY}$ and $\mu_{eff}$.
+We can infer from the modular invariance of scalar potential that the fixed points should be the local extrema of the scalar potential, as they act non-trivially (with residual symmetries) on the derivatives of the potential and must vanish.
+Although they may not be the global minima, experiences from modulus stabilization in string theory indicate that the minima should always lie at/near $\tau_{FP}=i,\omega$ for the tree-level scalar potential, as the scalar potential generally diverges for $\tau\rightarrow i\infty$.
+It was found in Novichkov:2021evw ; Feruglio:2021dte that the hierarchies among the charged leptons and quarks mass can follow solely from the properties of the modular forms (strictly speaking, the representation of the mass term bilinear with respect to the residual group) in the vicinity of symmetric points, avoiding the fine-tuning of the constant parameters (or without the need to introducing extra weighton-like scalar fields).
+Similar ideas, augmented with additional suppression powers of $q^{1/24}$ (or $(2\Im\tau)^{-1}$), can be used to generate the hierarchy between $\Lambda_{SUSY}$ and $\mu_{eff}$.
 Studies on target-space modular symmetry in string theory indicates target:space that all extrema lie on the boundary of the fundamental region or on the imaginary axis.
-We anticipate that, due to the SUSY breaking contributions from $X_{nl}$ sector, the extrema can change with respect to the change of $f_{0}$ so as that the VEV $\tau_{0}$ can be pushed slightly away from the fixed points $\tau_{FP}=i,\omega\equiv e^{i2\pi/3}$.
-Such deviation can lead to the small parameters within the $\mu_{eff}$.
-Besides, although the scalar potential for modulus field $\tau$ in general diverges when $\tau\rightarrow i\infty$, the scalar potential can still possibly lead to the stabilization of $\tau$ to a value with $\Im\tau_{0}$ being medium because of the contributions from $X_{nl}$.
-Then, for proper choice of $k_{3}$, the effective $\mu$ parameter can naturally be much smaller than the SUSY scale $\Lambda_{SUSY}$ by a large suppression factor, which origins from both powers of $q$ [or $(2\Im\tau)^{-1}$] and typical suppression factor from modular forms at $\tau_{0}$ (with medium $\Im\tau_{0}$).
-5.1 Natural $\mu_{eff}$ with modulus Stabilization near fixed points $\omega\equiv e^{i2\pi/3}$
-The Higgs superfields $H_{u},H_{d}$ in MSSM should always be singlets under modular flavor symmetry.
-They can be assigned with some singlet representations and non-trivial modular weights for their superpotential.
-For modular $A_{4}$ symmetry, at the point $\omega$, the following modular forms
-$Y_{\bf 1}^{(4)}(\omega)=Y_{\bf 1}^{(8)}(\omega)=Y_{\bf 1^{\prime}}^{(8)}(\omega)=Y_{\bf 1}^{(12)}(\omega)=Y_{\bf 1^{\prime}}^{(12)}(\omega)=Y_{\bf 1}^{(16)}(\omega)=0~{},$ (61)
-vanish as well as $Y_{\bf 1^{\prime\prime}}^{(16)}(\omega)$ etc.
-The small deviation of modulus VEV away from $\omega$ by modulus stabilization mechanism can act as the desired suppression parameter for $\mu_{eff}$.
-Expressions of modular forms with larger modular weights can be constructed from $Y_{\bf 3}^{(2)}\equiv(Y_{1},Y_{2},Y_{3})$ by
-$Y_{\bf 1^{\prime}}^{(8)}$ $=$ $(Y_{1}^{2}+2Y_{2}Y_{3})(Y_{3}^{2}+2Y_{1}Y_{2})~{},~{}~{}~{}Y_{\bf 1}^{(8)}=(Y_{1}^{2}+2Y_{2}Y_{3})^{2}~{},$
-$Y_{\bf 1^{\prime}}^{(8)}$ $=$ $(Y_{1}^{2}+2Y_{2}Y_{3})^{3}(Y_{3}^{2}+2Y_{1}Y_{2})~{},~{}~{}~{}Y_{\bf 1}^{(16)}=(Y_{1}^{2}+2Y_{2}Y_{3})^{4}~{}.$ (62)
-From the expansion Petcov:2022fjf at the point $z_{0}=\omega+\epsilon$ (with $\epsilon$ depending on the choices of $f_{0}$ satisfying $|\epsilon|\ll 0$) near $\omega$ point
+We anticipate that, due to SUSY breaking contributions (for example, from $X_{nl}$ sector), the extrema can change with respect to the change of $f_{0}$ so as that the VEV $\tau_{0}$ can be altered slightly away from the fixed points $\tau_{FP}=i,\omega$.
+Such deviations from fixed points by SUSY breaking effects can lead to large suppression factors that regenerate $\mu_{eff}$ of order the electroweak scale.
+Besides, although the scalar potential for modulus field $\tau$ in general diverges when $\tau\rightarrow i\infty$, the presence of $X_{nl}$ contributions can alter the scalar potential, making it possible to stabilize $\tau$ at an intermediate value.
+Consequently, for an appropriate choice of $k_{3}$, the effective $\mu_{eff}$ parameter can naturally be much smaller than the SUSY scale $\Lambda_{SUSY}$, due to a large suppression factor originating from both powers of $q$ [or $(2\Im\tau)^{-1}$] and typical suppression behavior of modular forms away from the fixed point $i\infty$ (at $\tau_{0}$ with medium $\Im\tau_{0}$).
+To regenerate the effective $\mu$-term with a stabilized modulus field, we need to construct a modular weight $-3$ bilinear for $H_{u}$ and $H_{d}$ using proper modular forms.
+For a modular invariant bilinear of the form $\psi_{i}^{c}M_{ij}(\tau)\psi_{j}$, where $M(\tau)_{ij}$ is a modular form of level $N$ and weight $K$, the zero entries of $M(\tau)_{ij}$ at a symmetric point in the mass matrices will generically become non-zero when $\tau$ deviates slightly from the symmetric point.
+The magnitudes of these residual symmetry breaking entries are controlled by the size of the departure $\epsilon$ from $\tau_{sym}$ and the field transformation properties under the residual symmetry group, which may depend on the modular weights Petcov:2023vws ; Novichkov:2021evw ; Petcov:2022fjf .
+That is, the entries of the mass matrices in the vicinity of the symmetric point are subjected to corrections of ${\cal O}(\epsilon^{l})$, where the powers $l$ corresponding to the representations of the residual symmetry group and depends only on how the representations of the fermion fields in the mass term bilinear decompose under the considered residual symmetry group.
+The degree of suppression, given by the integer $l$, can take values $l=0,1,\cdots,N-1$ in the case of $\tau_{sym}=i\infty$, $l=0,1,2$ in the case of $\tau_{sym}=\omega$ and $l=0,1$ in the case of $\tau_{sym}=i$.
+In Novichkov:2021evw , the authors systematically classify the residual symmetry representations for different choices of $\Gamma_{N}$ representations of matter fields.
+Such a formalism can also be applied to the regeneration of small $\mu_{eff}$ for the $H_{u}H_{d}$ bilinear to achieve a desired suppression factor ${\cal O}(\epsilon^{l})$ of the deviation parameter $\epsilon$ (from a given symmetric point).
+In our case, we need to specify the representations of the Higgs superfields $H_{u},H_{d}$ in MSSM under the residual symmetry group $\mathbb{Z}^{T}_{N}$ (with $N=3$) for $\tau_{sym}=i\infty$ and the residual group $\mathbb{Z}^{ST}_{3}$ for $\tau_{sym}=\omega$.
+Obviously, $H_{u},H_{d}$ should always be singlets under the modular flavor symmetry $\Gamma_{N}$.
+They can only be assigned some singlet representations of $\Gamma_{N}$ with non-trivial modular weights.
+The decompositions of the representation $({\bf r};k)$ multiplet of $\Gamma_{N}^{\prime}$ under the residual symmetry can be found in the appendix of Novichkov:2021evw .
+So, the desired suppression factor for the regenerated $\mu_{eff}$ can be directly obtained with proper choice of singlet representations and modular weights.
+5.1 Natural $\mu_{eff}$ with modulus stabilization near the fixed point $\omega\equiv e^{i2\pi/3}$
+In the vicinity of $\omega$, the degree of suppression for a bilinear is proven Novichkov:2021evw to be given by ${\cal O}(\epsilon^{l})$ with $l$ taking values $l=0,1,2$.
+In this study, we choose to directly calculate the suppression factor and its coefficient for the bilinear $H_{u}H_{d}$ using explicit model construction involving appropriate modular forms.
+It should be emphasized again that the exponentiation power of deviation parameter in the suppression factor depends only on the $\mathbb{Z}^{ST}_{3}$ (or $\mathbb{Z}^{T}_{3}$ for modular VEVs near $i\infty$) charge of $H_{u},H_{d}$ instead of the explicit form of modular forms.
+For modular $A_{4}$ symmetry, at the fixed point $\omega$, the following modular forms satisfy
+$Y_{\bf 1}^{(4)}(\omega)=Y_{\bf 1}^{(8)}(\omega)=Y_{\bf 1^{\prime}}^{(8)}(\omega)=0~{}.$ (70)
+The small deviation of modulus VEV away from $\omega$, induced by the modulus stabilization mechanism, can act as the desired suppression parameter for $\mu_{eff}$.
+Expressions of modular forms with larger modular weights can be constructed from $Y_{\bf 3}^{(2)}\equiv(Y_{1},Y_{2},Y_{3})$, for example,
+$Y_{\bf 1^{\prime}}^{(8)}$ $=$ $(Y_{1}^{2}+2Y_{2}Y_{3})(Y_{3}^{2}+2Y_{1}Y_{2})~{},~{}~{}~{}Y_{\bf 1}^{(8)}=(Y_{1}^{2}+2Y_{2}Y_{3})^{2}~{}.$ (71)
+From the expansion Petcov:2022fjf of modular functions at the point $z_{0}=\omega+\epsilon$ (with $\epsilon$ depending on the choices of $f_{0}$ such that $|\epsilon|\ll 0$)
 $\frac{Y_{2}(z_{0})}{Y_{1}(z_{0})}$ $\simeq$ $\omega\left(1+2.235i\epsilon-2.95717\epsilon^{2}+\cdots\right)~{},$
 $\frac{Y_{3}(z_{0})}{Y_{1}(z_{0})}$ $\simeq$ $-\frac{1}{2}\omega^{2}\left(1+4.47i\epsilon-10.9096\epsilon^{2}+\cdots\right),$ (72)
-one can arrive at the expansion for $Y_{\bf 1}^{(16)}(z_{0})$ etc as
-$Y_{\bf 1^{\prime}}^{(8)}(\omega+\epsilon)$ $\simeq$ $(10.5823+6.10969i)\epsilon-(39.9459-69.1884i)\epsilon^{2}+\cdots~{},$
-$Y_{\bf 1}^{(16)}(\omega+\epsilon)$ $\simeq$ $1325.96\epsilon^{4}+18871.7i\epsilon^{5}+\cdots~{},$ (64)
+we can obtain the expansion for $Y_{\bf 1}^{(8)}(z_{0})$ etc as
+$Y_{\bf 1}^{(8)}(\omega+\epsilon)$ $\simeq$ $-36.4138\epsilon^{2}-259.129i\epsilon^{3}+\cdots~{},$ (73)
 with $Y_{1}(\omega)\approx 0.948674$.
@@
 $W\supseteq\alpha\Lambda_{SUSY}\frac{Y_{\bf r}^{(k_{r})}(\tau)}{\eta(\tau)^{2k_{3}}}H_{u}H_{d}~{},$ (74)
-with $k_{H_{u}}+k_{H_{d}}=k_{r}-k_{3}+3$ for $62k_{3}\in 24\mathbb{Z}+$ 6 and ${\bf r}(H_{u})\otimes{\bf r}(H_{d})\otimes{\bf r}\supset{\bf 1}$.
+where $k_{H_{u}}+k_{H_{d}}=k_{r}-k_{3}+3$ and ${\bf r}(H_{u})\otimes{\bf r}(H_{d})\otimes{\bf r}\supset{\bf 1}$.
 When the modulus field is stabilized close to the point $\omega$, the effective $\mu$-term is given by
-$\mu_{eff}$ $=$ $\alpha\Lambda_{SUSY}\eta^{-2k_{3}}(\omega+\epsilon)Y_{\bf r}^{(k_{r})}(\omega+\epsilon)(2\Im\tau)^{(k_{u}+k_{d})/2}~{},$ (66)
+$\mu_{eff}$ $=$ $\alpha\Lambda_{SUSY}\eta^{-2k_{3}}(\omega+\epsilon)Y_{\bf r}^{(k_{r})}(\omega+\epsilon)(2\Im\tau)^{(k_{H_{u}}+k_{H_{d}})/2}~{},$ (75)
 $=$ $\alpha\Lambda_{SUSY}\eta^{-2k_{3}}(\omega)Y_{\bf r}^{(k_{r})}(\omega+\epsilon)(\sqrt{3})^{(k_{r}-k_{3}+3)/2}$
-$\approx$ $\alpha\Lambda_{SUSY}\left[0.711506\cdot e^{-0.523599}\right]^{-k_{3}/2}Y_{\bf r}^{(k_{r})}(\omega+\epsilon)(\sqrt{3})^{k_{r}+3/2}~{}.$
-So, to obtain much suppressed $\mu_{eff}$, we either need very large negative value $k_{3}$ or large $k_{r}$ to get large powers of $\epsilon$.
-However, it can be calculated that the suppression from $k_{3}$ power is not very efficient to get tiny number.
-So we choose to adopt large powers of $\epsilon$ for suppressed $\mu_{eff}$ with proper modular weights and representations for $H_{u},H_{d}$.
-For example, we can adopt the choice
-$(r_{H_{u}},k_{H_{u}})\sim(~{}{\bf 1^{\prime}},~{}6),~{}~{}(r_{H_{d}},k_{H_{d}})\sim(~{}{\bf 1^{\prime\prime}},~{}7)~{},$ (67)
-and the choice of modular forms $Y_{\bf 1}^{(16)}$ (or possibly $Y_{\bf 1}^{(4)},Y_{\bf 1^{\prime}}^{(8)},Y_{\bf 1}^{(12)}$ with proper $k_{H_{u}}$ and $k_{H_{d}}$). So, we have
-$\mu_{eff}$ $=$ $\frac{\alpha}{\bar{N}_{Y_{\bf 1}^{(16)}}}\eta(\omega)^{-6}Y_{\bf 1}^{(16)}(\omega+\epsilon)(\sqrt{3})^{(k_{u}+k_{d})/2}~{},$ (68)
-$\approx$ $\frac{\alpha}{360.275}(0.8e^{-0.131i})^{-6}\Lambda_{SUSY}1325.96\epsilon^{4}(\sqrt{3})^{13/2}~{},$
-$\approx$ $4.98\times 10^{2}e^{0.785i}\alpha\Lambda_{SUSY}\epsilon^{4}~{}.$
-Here $\bar{N}_{Y_{\bf 1}^{(16)}}=360.275$ denotes the global normalization factor Petcov:2023fwh , which does not depend on the VEV of the modulus $\tau$ and defined by
+$\approx$ $\alpha\Lambda_{SUSY}\left[0.711506\cdot e^{-0.523599i}\right]^{-k_{3}/2}Y_{\bf r}^{(k_{r})}(\omega+\epsilon)(\sqrt{3})^{k_{r}+3/2}~{}.$
+To achieve a significantly suppressed $\mu_{eff}$, we either need a very large negative value $k_{3}$ or a tiny $\epsilon$.
+However, calculations show the suppression from $k_{3}$ power is not very efficient for obtaining a tiny number.
+Therefore, we choose to use a small $\epsilon$ to suppress $\mu_{eff}$ with appropriate modular weights and representations for $H_{u},H_{d}$.
+For instance, we can choose
+$(r_{H_{u}},k_{H_{u}})\sim(~{}{\bf 1^{\prime}},~{}4),~{}~{}(r_{H_{d}},k_{H_{d}})\sim(~{}{\bf 1^{\prime\prime}},~{}4)~{},$ (76)
+and the modular forms $Y_{\bf 1}^{(8)}$ (or, for example, $Y_{\bf 1}^{(4)}$ with suitable $k_{H_{u}}$ and $k_{H_{d}}$). Thus, we have
+$\mu_{eff}$ $=$ $\frac{\alpha\Lambda_{SUSY}}{|\bar{N}_{Y_{\bf 1}^{(8)}}|}\eta(\omega)^{-6}Y_{\bf 1}^{(8)}(\omega+\epsilon)(\sqrt{3})^{(k_{H_{u}}+k_{H_{d}})/2}~{},$ (77)
+$\approx$ $-0.927\times 10^{3}e^{i\frac{\pi}{4}}\alpha\Lambda_{SUSY}\epsilon^{2}~{}.$
+Here $|\bar{N}_{Y_{\bf 1}^{(8)}}|=1.3423$ denotes the global normalization factor Petcov:2023fwh for $Y_{\bf 1}^{(8)}$, which is independent of the modulus VEV and is defined by
 $(\overline{{\rm N}}^{(K)}_{\rm Y})^{2}=\int_{D}\int(N^{(K)}_{\rm Y}(\tau,\tau^{*}))^{2}\,\frac{dxdy}{y^{2}}\equiv\int_{D}\int\sum_{i}|Y^{(K)}_{{\bf r}i}(\tau)|^{2}(2\,y)^{K}\frac{dx\,dy}{y^{2}}\,,$ (78)
@@
 […]
-for non-cusp modular forms, within which $T$ is a real constant, the region $D_{T}$ is the lower part of the fundamental domain bounded from above by the line $y=T$, $T>1$, and $|a_{0}|^{2}$ is the constant $q$-independent term in the $q$-expansion of $\sum_{i}|Y^{(K)}_{{\bf r}i}(\tau)|^{2}$.
-For $\epsilon\sim 10^{-5}$ and $\Lambda_{SUSY}$ typically of order the reduced Planck scale, we can obtain the desired $\mu_{eff}$ at electroweak scale.
-Note that the smallness of deviation parameter $\epsilon$ (away from $\omega$) can be the consequence of SUSY breaking that depends on the choice of $f_{0}$, for example, if the parameter choice with $f_{0}=0$ leads to the global minimum at $\omega$.
-In general, adopting the modular form $Y_{\bf 1}^{(4(k+2))}$ with $(k\in N)$, we can obtain suppression factor $\epsilon^{k+2}$ for tiny $\epsilon(f_{0})$.
-We show some of the benchmark points that can lead to the suppression factor in Table 5.1.
-$m$ $n$ $f_{0}/M_{\mathrm{Pl}}^{2}$ $c_{1}/M_{\mathrm{Pl}}$ $X_{nl}/M_{\mathrm{Pl}}$ $\epsilon$ 00 $3$ $0.2$ $10.0$ $5.194\times 10^{-4}$ 00 $3$ $0.2$ $0.1$ $0.2$ $8.767\times 10^{-4}$ 00 $2$ $0.2$ $0.1$ $0.02$ $4.227\times 10^{-4}$ 00 $3$ $0.2$ $0.1$ $0.02$ $6.916\times 10^{-4}$ $1$ $3$ $0.2$ $0.1$ $0.02$ $9.396\times 10^{-4}$ $1$ $3$ $0.2$ $0.1$ 00 $4.711\times 10^{-4}$ Table 1: Parameter choices for scenario with $k_{X}=0$, which can stabilize the modulus at the point $\omega+\epsilon$ (for $|\epsilon\sim 10^{-4}|$).
+for non-cusp modular forms, where $T$ is a real constant, $D_{T}$ is the region below the line $y=T$ in the fundamental domain with $T>1$, and $|a_{0}|^{2}$ is the constant $q$-independent term in the $q$-expansion of $\sum_{i}|Y^{(K)}_{{\bf r}i}(\tau)|^{2}$.
+From the decompositions of $\Gamma_{N}$ representations under the residual symmetry group in appendix of Novichkov:2021evw , it can be verified that the suppression factor of $\epsilon$ for the $H_{u}(~{}{\bf 1^{\prime}},~{}4)$ and $H_{d}(~{}{\bf 1^{\prime\prime}},~{}4)$ bilinear is given by $({k_{H_{u}}+k_{H_{d}}+3})~{}{\rm modulo}~{}3=2$, which is consistent with the result from our explicit construction.
+For $\epsilon\sim 10^{-5}$, $\alpha\sim{\cal O}(1)$ and $\Lambda_{SUSY}$ typically of the order $\sqrt{F}\sim\sqrt{m_{3/2}M_{Pl}}\sim 10^{10}$ GeV for $m_{3/2}\sim{\cal O}({\rm 10^{2}~{}GeV})$ in gravity mediation SUSY breaking mechanism, we can obtain a desired $\mu_{eff}$ at electroweak scale.
+If $\Lambda_{SUSY}$ is naively taken to be the Planck scale, smaller $\epsilon\sim 10^{-10}$ or a tiny $\alpha$ is needed to achieve an electroweak-scale $\mu_{eff}$.
+Note that the smallness of deviation parameter $\epsilon$ (away from $\omega$) can be a consequence of SUSY breaking, which depends on the choice of $f_{0}$.
+For example, if the parameter choice with $f_{0}=0$ leads to the global minimum at $\omega$, SUSY breaking contributions may slightly shift the minimum to some point near $\omega$.
+We present some benchmark points that can lead to small suppression factors in Table 5.1.
+$m$ $n$ $f_{0}/M_{\mathrm{Pl}}^{2}$ $c_{1}/M_{\mathrm{Pl}}$ $\epsilon$ $\mathcal{P}(j(\tau))$ 00 $3$ $0.2$ $0.01$ $4.711\times 10^{-5}$ $1$ 00 $3$ $0.2$ $0.1$ $2.529\times 10^{-5}$ $1$ 00 $2$ $0.2$ $0.1$ $6.758\times 10^{-6}$ $1$ 00 $3$ $0.2$ $0.1$ $4.711\times 10^{-5}$ $1$ $1$ $3$ $0.2$ $0.1$ $4.711\times 10^{-5}$ $1$ 00 $3$ $0.2$ $0.1$ $2.920\times 10^{-5}$ $j-4146.6830+184.0955i$ $1$ $3$ $0.2$ $0.1$ $6.587\times 10^{-5}$ $j-4144.8675+185.3946i$ Table 1: Parameter choices for the scenario with $k_{X}=0$, which can stabilize the modulus at the point $\omega+\epsilon$ (for $|\epsilon\sim 10^{-5}|$).
 5.2 Natural $\mu_{eff}$ with modulus stabilization at medium $\Im\tau$
-Although the scalar potential for modulus in general diverges for $\tau\rightarrow i\infty$, the introduction of nilpotent Goldstino $X$ superfield or multiple moduli/matter fields may alter the shape of the scalar potential at intermediate values of $\tau$, making it possible the stabilization of modulus at $\tau_{0}$ with medium $\Im\tau_{0}$ (for example, $\Im\tau_{0}\sim 5-10$).
-Detailed discussions on such a possibility will be given in our subsequent studies.
-In this work, we just concentrate on the implications of modulus stabilization at medium $\Im\tau$ on the SUSY $\mu$-problem.
-Note that the stabilized VEV $\tau_{0}$ (with medium $\Im\tau_{0}$) can be regarded to lie close to the $i\infty$ fixed point.
-To obtain the suppressed effective $\mu$-parameter, in addition to the suppression factor from powers of $q^{1/24}$ or $(2\Im\tau)^{-1}$, we would also like to choose the modular function that vanishes at $\Im\tau=i\infty$.
+When the modulus field is stabilized near $i\infty$, similar to the case where modulus VEV lies in the vicinity of $\omega$, the exponentiation power $l$ of deviation parameter $\epsilon$ in the suppression factor for the $H_{u}H_{d}$ bilinear can also be proven Novichkov:2021evw to depend only on the $\mathbb{Z}^{T}_{3}$ charge of $H_{u},H_{d}$, rather than the explicit form of modular forms.
+Since the coefficient of $\epsilon^{l}$ cannot be simply determined and may involve additional suppression factors, we opt to directly calculate the suppression factor and the coefficient for the bilinear $H_{u}H_{d}$ using an explicit model construction.
+[…]
+To obtain the suppressed effective $\mu$-parameter, in addition to possible suppression factor from powers of $q^{1/24}$ (from $\eta(\tau)$) or $(2\Im\tau)^{-1}$ (from normalization factor), we would also like to choose the modular function that vanishes at $\Im\tau=i\infty$.
 For example, the modular functions satisfy
 $Y_{\bf 1^{\prime}}^{(4)}(i\infty)=Y_{\bf 1^{\prime}}^{(8)}(i\infty)=Y_{\bf 1^{\prime\prime}}^{(8)}(i\infty)=0~{},$ (81)
-as well as the vanishing singlet components from their direct products, such as the modular functions $Y_{\bf 1}^{(12)}(i\infty)$,$Y_{\bf 1^{\prime}}^{(16)}(i\infty)$.
-The asymptotic behavior of the previous modular forms near $i\infty$ within the fundamental domain can be calculated from Petcov:2023vws (for $z_{0}$ with $\Im z_{0}$ being sufficient large)
-$Y_{\bf 1^{\prime}}^{(4)}(z_{0})$ $=$ $-12p\epsilon+96p^{4}\epsilon^{4}~{},$
-$Y_{\bf 1^{\prime}}^{(8)}(z_{0})$ $=$ $-6p\epsilon-1230p^{4}\epsilon^{4}~{},$
-$Y_{\bf 1^{\prime\prime}}^{(8)}(z_{0})$ $=$ $144p^{2}\epsilon^{2}-2304p^{5}\epsilon^{5}~{},$ (73)
+as well as the vanishing singlet components from their direct products, such as the modular function $Y_{\bf 1}^{(12)}(i\infty)$.
+The asymptotic behavior of the previous modular forms near $i\infty$ within the fundamental domain can be calculated from Petcov:2023vws (for $z_{0}$ with $\Im z_{0}$ sufficiently large)
+$Y_{\bf 1^{\prime}}^{(4)}(z_{0})$ $=$ $-12p\epsilon+96p^{4}\epsilon^{4}+{\cal O}(\epsilon^{7})~{},$
+$Y_{\bf 1^{\prime}}^{(8)}(z_{0})$ $=$ $-6p\epsilon-1230p^{4}\epsilon^{4}+{\cal O}(\epsilon^{7})~{},$
+$Y_{\bf 1^{\prime\prime}}^{(8)}(z_{0})$ $=$ $144p^{2}\epsilon^{2}-2304p^{5}\epsilon^{5}+{\cal O}(\epsilon^{8})~{},$ (82)
 with
@@
 where $\epsilon\ll 1$.
-Let us suppose that the modulus could be stabilized at the imaginary axis (within the fundamental domain) with $z_{0}=it_{0}$ for medium large $t_{0}$, when proper SUSY breaking contributions with non-vanishing $f_{0}$ are taken into account.
-The effective $\mu$-term from the bilinear term for $H_{u}H_{d}$ in (65), which satisfies $k_{H_{u}}+k_{H_{d}}=k_{r}-k_{3}+3$ for $62k_{3}\in 24\mathbb{Z}+$ 6, is given by
-$\mu_{eff}$ $=$ $\alpha\Lambda_{SUSY}\eta(it_{0})^{-2k_{3}}Y_{\bf r}^{(k_{r})}(it_{0})(2t_{0})^{(k_{u}+k_{d})/2}~{},$ (75)
-$\approx$ $\alpha\Lambda_{SUSY}\eta(it_{0})^{-2k_{3}}Y_{\bf r}^{(k_{r})}(it_{0})(2t_{0})^{(k_{r}-k_{3}+3)/2}~{},$
-$\approx$ $\alpha\Lambda_{SUSY}e^{\frac{k_{3}\pi t_{0}}{6}}(2t_{0})^{-\frac{k_{3}}{2}}Y_{\bf r}^{(k_{r})}(it_{0})(2t_{0})^{\frac{k_{r}+3}{2}}~{}.$
-with proper normalization factors for $H_{u},H_{d}$ from the Kahler potential.
+Let us suppose that the modulus field could be stabilized at the imaginary axis (within the fundamental domain) with $z_{0}=it_{0}$ for medium large $t_{0}$, when proper SUSY breaking contributions with non-vanishing $f_{0}$ are taken into account.
+Naively, if the desired suppression factor $10^{-20}$ to regenerate $\mu_{eff}$ comes solely from $\epsilon^{2}$ when $\Lambda_{SUSY}$ is naively taken to be the Planck scale, very large $t_{0}$ (here $t_{0}\gtrsim 10$) is needed.
+Such a large $t_{0}$ possibly causes difficulties for modulus stabilization realizations, as the scalar potential for modulus field generally diverges for $\tau\rightarrow i\infty$.
+Therefore, it is desirable to introduce additional suppression factors other than $\epsilon$ so that medium large $t_{0}$ can lead to desired suppression factors.
+The effective $\mu$-term from the bilinear term for $H_{u}H_{d}$ in (74), which satisfies $k_{H_{u}}+k_{H_{d}}=k_{r}-k_{3}+3$, is given by
+$\mu_{eff}$ $=$ $\alpha\Lambda_{SUSY}\eta(it_{0})^{-2k_{3}}Y_{\bf r}^{(k_{r})}(it_{0})(2t_{0})^{(k_{H_{u}}+k_{H_{d}})/2}~{},$ (84)
+$\approx$ $\alpha\Lambda_{SUSY}e^{\frac{k_{3}\pi t_{0}}{6}}(2t_{0})^{-\frac{k_{3}}{2}}Y_{\bf r}^{(k_{r})}(it_{0})(2t_{0})^{\frac{k_{r}+3}{2}}~{},$
+where proper normalization factors for $H_{u}$,$H_{d}$ from the Kahler potential are taken into account.
 For fixed $k_{r}$, we require that the factor
-$\left[e^{\frac{\pi t_{0}}{3}}(2t_{0})^{-1}\right]^{\frac{k_{3}}{2}}\ll 1.$ (76)
-As the factor
+$\left[e^{\frac{\pi t_{0}}{3}}(2t_{0})^{-1}\right]^{\frac{k_{3}}{2}}\ll 1,$ (85)
+is tiny to act as an additional suppression factor other than $\epsilon$. As the factor
 $e^{\frac{\pi t_{0}}{3}}(2t_{0})^{-1}>1~{},$ (86)
@@
 We can choose the following modular weights and representations
-$(r_{H_{u}},k_{H_{u}})\sim(~{}{\bf 1^{\prime\prime}},~{}14),~{}~{}(r_{H_{d}},k_{H_{d}})\sim(~{}{\bf 1},~{}15)~{},$ (78)
-and the choice of modular forms $Y_{\bf 1^{\prime\prime}}^{(8)}$ (or possibly $Y_{\bf 1}^{(12)}$, $Y_{\bf 1^{\prime}}^{(16)}$ etc with proper $k_{H_{u}}$ and $k_{H_{d}}$) with $k_{3}=-21$ to generate suppressed $\mu_{eff}$ parameter.
+$(r_{H_{u}},k_{H_{u}})\sim(~{}{\bf 1},~{}16),~{}~{}(r_{H_{d}},k_{H_{d}})\sim(~{}{\bf 1^{\prime}},~{}16)~{},$ (87)
+and the choice of modular form $Y_{\bf 1^{\prime\prime}}^{(8)}$ (or, for example, $Y_{\bf 1}^{(12)}$, $Y_{\bf 1^{\prime}}^{(16)}$ etc with proper $k_{H_{u}}$ and $k_{H_{d}}$) with $k_{3}=-21$ to generate a suppressed $\mu_{eff}$ parameter.
 In this case, the effective $\mu_{eff}$ parameter is
 $\mu_{eff}$ $=$ $\frac{\alpha}{\bar{N}_{Y_{\bf 1^{\prime\prime}}^{(8)}}}\Lambda_{SUSY}\left[e^{\frac{\pi t_{0}}{3}}(2t_{0})^{-1}\right]^{-21/2}144e^{-\frac{4\pi t_{0}}{3}}(2t_{0})^{11/2}~{},$ (88)
-$=$ $4.46{\alpha}\Lambda_{SUSY}e^{-29\pi t_{0}/6}(2t_{0})^{16}~{},$
-with the global normalization factor ${\bar{N}_{Y_{\bf 1^{\prime}}^{(8)}}}=32.2794$.
-It can be calculated that the value of $\mu_{eff}\lesssim 10^{-20}M_{SUSY}$ can be satisfied for $t_{0}\gtrsim 5.58$.
-If we choose the modular form from direct product of $(Y_{\bf 1^{\prime}}^{(4)})^{k+2}$ with larger $k\in N$, smaller $t_{0}$ is needed to get electroweak scale $\mu_{eff}$ parameter.
-Tiny value of $\mu_{eff}$ is an joint effect of suppression from powers of $q$ (or $(2\Im\tau)^{-1}$) and the asymptotic suppression behavior of typical modular forms away from the fixed point $i\infty$.
+$=$ $7.05{\alpha}\Lambda_{SUSY}e^{-29\pi t_{0}/6}(2t_{0})^{16}~{},$
+where ${\bar{N}_{Y_{\bf 1^{\prime\prime}}^{(8)}}}=20.4152$ is the global normalization factor for the modular form $Y_{\bf 1^{\prime\prime}}^{(8)}$.
+Calculations show that the value $\mu_{eff}\lesssim 10^{-20}M_{SUSY}$ can be achieved for $t_{0}\gtrsim 5.58$, if $\Lambda_{SUSY}$ is naively taken to be the Planck scale.
+For $\Lambda_{SUSY}$ typically of the order $\sqrt{F}\sim\sqrt{m_{3/2}M_{Pl}}\sim 10^{10}$ GeV, it is sufficient to choose $t_{0}\gtrsim 3.2$ to satisfy $\mu_{eff}\lesssim 10^{-8}M_{SUSY}$.
+The tiny value of $\mu_{eff}$ results from the combined effect of suppression from powers of $q$ (or $(2\Im\tau)^{-1}$) and the asymptotic suppression behavior of typical modular forms away from the fixed point $i\infty$.
+It can also be verified that the suppression power of $\epsilon=\exp[-\frac{2\pi}{3}\Im z_{0}]$ for $H_{u}(~{}{\bf 1},~{}16)$ and $H_{d}(~{}{\bf 1^{\prime}},~{}16)$ bilinear is given by the power factor $l$ of $(\rho_{i}^{c}\rho_{j})^{*}=\zeta^{l}$ for $\zeta=\exp(2\pi i/3)$, which can be read out from Novichkov:2021evw to be $(1\cdot\zeta)^{*}=\zeta^{2}$ and predict $l=2$.
+This power factor for $\epsilon$ is consistent with the result from our explicit construction.
 6 Conclusions
-It is known that the solutions of SUSY $\mu$ problem involve the forbidden of the bare $\mu$-term via some symmetry and the re-generation of the effective $\mu$-term after the symmetry breaking.
-We propose that, the symmetry that responsible for the solution of the SUSY $\mu$-problem is a modular flavor symmetry.
-The bare $\mu$-term is forbidden by modular symmetry after assigning non-vanishing modular weights and/or non-trivial singlet representations for $H_{u},H_{d}$ superfields.
-Then, the effective $\mu$-term will be re-generated after the modulus field is stabilized, with the corresponding modulus VEV breaks the modular symmetry completely or partially (at the fixed points).
+In this study, we propose a novel solution to the SUSY $\mu$-problem within the framework of modular flavor symmetry.
+The explicit $\mu$-term is prohibited by modular symmetry after assigning non-vanishing modular weights and/or non-trivial singlet representations to the $H_{u},H_{d}$ superfields.
+The effective $\mu$-term is then regenerated following the stabilization of the modulus field, with the corresponding modulus VEV breaks the modular symmetry partially (at the fixed points) or completely (in the vicinity of the fixed points).
+Since naturalness requires $\mu_{eff}$ to lie at the electroweak scale or soft SUSY breaking scale, a large suppression factor leading to $\mu_{eff}\ll\Lambda_{SUSY}$ is required when the modulus field is stabilized.
+Such a large suppression factor can naturally arise from a modulus VEV in the vicinity of certain fixed points.
 Ordinary single modulus stabilization mechanism can be affected by SUSY breaking contributions.
-The goldstino, which is necessarily present as a consequence of UV SUSY breaking and mixes with gravitino, can be promoted into a constrained nilpotent superfield $X_{nl}$ and provides a way for an arbitrary non-supersymmetric low-energy theory to realize SUSY non-linearly.
-We discuss the stabilization mechanism of a single modulus field with the presence of SUSY breaking contributions, described in the non-linear SUSY realization scheme with nilpotent goldstino $X_{nl}$.
-The presence of $X_{nl}$ contributions can slightly change the original stabilized modulus VEV to a value located in the vicinity of the fixed points.
-Natural value of $\mu_{eff}$, which is much suppressed relative to the SUSY scale, can be the consequence of series expansion of typical modular forms in terms of the tiny deviation parameter (away from the fixed point $\omega$), which is similar to the mechanism to naturally generate mass hierarchies in the vicinity of the fixed points.
-For modulus VEV with medium $\Im\tau$, the much suppressed $\mu_{eff}$ is the joint consequence of suppression from powers of $q$ [or $(2\Im\tau)^{-1}$], with the exponent of the power determined by the assignments of the modular weights for $H_{u}$ and $H_{d}$, and the asymptotic suppression behavior of typical modular forms away from the fixed point $i\infty$.
+The Goldstino, which is necessarily present as a consequence of UV SUSY breaking and mixes with gravitino, can be promoted into the constrained nilpotent superfield $\textbf{X}_{nl}$ and provides a way for an arbitrary non-supersymmetric low-energy theory to realize SUSY non-linearly.
+We discuss the stabilization mechanism of a single modulus field in the presence of SUSY breaking contributions, described by the non-linear SUSY realization scheme with a nilpotent Goldstino $\textbf{X}_{nl}$.
+The presence of $\textbf{X}_{nl}$ contributions can slightly alter the original stabilized modulus VEV to a value near the fixed points.
+Natural value of $\mu_{eff}$, which is significantly suppressed relative to the SUSY scale, can be achieved from the expansion of typical modular forms using a tiny deviation parameter near the fixed points.
+This is similar to the mechanism that can naturally generate mass hierarchies for quarks and leptons in the vicinity of certain fixed points.
+The exponentiation power of the deviation parameter in the suppression factor, in fact, depends solely on the $\mathbb{Z}_{3}^{ST}$ charge (for fixed point $\omega$), or $\mathbb{Z}^{T}_{3}$ charge (for fixed point $i\infty$) of $H_{u}$, $H_{d}$, rather than the explicit form of modular forms.
+For a modulus VEV with an intermediate value of $\Im\tau$, the significantly suppressed $\mu_{eff}$ results from the combined effects of suppression by powers of $q^{1/24}$ [or $(2\Im\tau)^{-1}$], where the exponentiation power determined by the modular weights of $H_{u}$ and $H_{d}$, along with the asymptotic suppression behavior of typical modular forms away from the fixed point $i\infty$, taking the form of appropriate power of the tiny deviation parameter.
+The modulus stabilization mechanism with nilpotent $\textbf{X}_{nl}$ superfield can be fairly generic and ubiquitous.
+The regeneration of small $\mu_{eff}$ term following modulus stabilization, which just needs a modulus VEV near some fixed points, is independent on the concrete modulus stabilization mechanism.
 Appendix A Properties of Modular $A_{4}$

### v4

@@
 We examine the stabilization mechanism of a single modulus field with the presence of SUSY breaking contributions described by the non-linear SUSY realization scheme involving a nilpotent Goldstino $\textbf{X}_{nl}$ superfield.
-The inclusion of $\textbf{X}_{nl}$ contributions can slightly alter the original stabilized modulus VEV to a value near some fixed point.
 A natural small $\mu_{eff}$, significantly smaller than the SUSY scale, can result from either the expansion of typical modular forms using a small deviation parameter near the fixed point $\omega$, or from the combined effects of suppression by powers of $q^{1/24}$ [or $(2\Im\tau)^{-1}$] along with the asymptotic suppression behavior of typical modular forms away from the fixed point $i\infty$, taking the form of appropriate power of the tiny deviation parameter.
+A natural small $\mu_{eff}$ can also be achieved by a weighton-like mechanism for $H_{u}H_{d}$ bilinear.
 1 Introduction
@@
 Modular symmetry is a geometrical symmetry of torus $T^{2}$ and it corresponds to a change of the corresponding cycle basis.
-Modular symmetry can naturally arise in string theory, for example, in the (exact) moduli spaces of Calabi-Yau threefolds like the quintic, and more generically in large classes of elliptic Calabi-Yau threefolds.
+Modular symmetry can naturally arise in string theory, for example, in the (exact) moduli spaces of Calabi-Yau threefolds like the quintic, and more generically in large classes of elliptic Calabi-Yau threefolds CY:modular1 ; CY:modular2 .
 Recently, modular flavor symmetry was proposed in Feruglio:2017spp to explain the flavor puzzles of the Standard Model (plus the massive neutrinos), which no longer necessitates the introduction of the flavon fields that may cause difficulties in flavor model buildings.
@@
 In this framework, the masses and mixings of quarks and leptons can be predicted with very few input parameters.
-Modular flavor models based on $\Gamma_{2}\simeq S_{3}$ Kobayashi:2018vbk ; Okada:2019xqk ; Du:2020ylx , $\Gamma_{3}\simeq A_{4}$ Kobayashi:2018vbk ; Petcov:2018snn ; Criado:2018thu ; Kobayashi:2018scp ; Okada:2018yrn ; Novichkov:2018yse ; Okada:2020ukr ; Petcov:2022fjf ; CentellesChulia:2023osj ; Kumar:2023moh ; Nomura:2024ctl ; Pathak:2024sei , $\Gamma_{4}\simeq S_{4}$ Penedo:2018nmg ; Petcov:2018snn ; Novichkov:2018ovf ; Kobayashi:2019xvz ; Wang:2020dbp ; Qu:2021jdy , $\Gamma_{5}\simeq A_{5}$ Petcov:2018snn ; Novichkov:2018nkm ; Ding:2019xna ; Yao:2020zml ; deMedeirosVarzielas:2022ihu ; Abbas:2024bbv and non-holomorphic modular flavor models Qu:2024rns ; Nomura:2024atp ; Ding:2024inn ; Li:2024svh had been used in various studies to explain the low energy flavor structures of the Standard Model.
+Modular flavor models based on $\Gamma_{2}\simeq S_{3}$ Kobayashi:2018vbk ; Okada:2019xqk ; Du:2020ylx , $\Gamma_{3}\simeq A_{4}$ Kobayashi:2018vbk ; Petcov:2018snn ; Criado:2018thu ; Kobayashi:2018scp ; Okada:2018yrn ; Novichkov:2018yse ; Okada:2020ukr ; Petcov:2022fjf ; CentellesChulia:2023osj ; Kumar:2023moh ; Nomura:2024ctl ; Pathak:2024sei , $\Gamma_{4}\simeq S_{4}$ Penedo:2018nmg ; Petcov:2018snn ; Novichkov:2018ovf ; Kobayashi:2019xvz ; Wang:2020dbp ; Qu:2021jdy , $\Gamma_{5}\simeq A_{5}$ Petcov:2018snn ; Novichkov:2018nkm ; Ding:2019xna ; Yao:2020zml ; deMedeirosVarzielas:2022ihu ; Abbas:2024bbv and non-holomorphic modular flavor models Qu:2024rns ; Nomura:2024atp ; Ding:2024inn ; Li:2024svh had been studied in various works to explain the low energy flavor structures of the Standard Model.
 Modular symmetry can also be used to solve the strong CP problem Feruglio:2023uof ; Feruglio:2024ytl .
@@
 The vacuum expectation value (VEV) of the modulus field can be the only source of flavour symmetry breaking.
-It can also be the only source of CP symmetry breaking Novichkov:2019sqv ; Feruglio:2021dte ; Feruglio:2024ytl .
+It can also be the only source of CP symmetry breaking Baur:2019kwi ; Novichkov:2019sqv ; Feruglio:2021dte ; Ishiguro:2020nuf ; Feruglio:2024ytl .
 The VEV of the modulus field $\tau$ generally break the modular symmetry completely, except at three inequivalent fixed points $\tau=i\infty,i,e^{i2\pi/3}$ in the fundamental domain.
@@
 The presence of the anti-brane-like sources can be described in terms of a nilpotent Goldstino superfield KS in a supersymmetric effective field theory.
-In the framework of non-linearly realized SUSY, the couplings of chiral, gauge and gravity superfields to the constrained nilpotent Goldstino superfield $X$ can be used to describe spontaneous SUSY breaking in supergravity Burgess:2021juk ; Aparicio:2015psl ; NL:SUGRA1 ; NL:SUGRA2 ; NL:SUGRA3 ; NL:SUGRA4 ; NL:SUGRA5 ; NL:SUGRA6 and the presence of the anti-brane type sources with a supersymmetric action.
+In the framework of non-linearly realized SUSY, the couplings of chiral, gauge and gravity superfields to the constrained nilpotent Goldstino superfield $\textbf{X}_{nl}$ can be used to describe spontaneous SUSY breaking in supergravity Burgess:2021juk ; Aparicio:2015psl ; NL:SUGRA1 ; NL:SUGRA2 ; NL:SUGRA3 ; NL:SUGRA4 ; NL:SUGRA5 ; NL:SUGRA6 and the presence of the anti-brane type sources with a supersymmetric action.
 Contributions from nilpotent chiral superfield can uplift the scalar potential responsible for the stabilization of the modulus field.
@@
 In Sec 4, we discuss the stabilization of singlet modulus field in the presence of SUSY breaking contributions described by the non-linear realized SUSY scheme with the nilpotent Goldstino ${\textbf{X}}_{nl}$ superfield.
-In Sec 5, we discuss the regeneration of a naturally small $\mu_{eff}$ after modulus stabilization, when the modulus field is stabilized near some fixed points, such as $\omega\equiv e^{i2\pi/3}$ and $i\infty$ Sec 6 contains our conclusions.
+In Sec 5, we discuss the regeneration of a naturally small $\mu_{eff}$ after modulus stabilization, when the modulus field is stabilized near some fixed points, such as $\omega\equiv e^{i2\pi/3}$ and $i\infty$ .
+Sec 6 contains our conclusions.
 2 Modular $\Gamma_{N}$ symmetry
@@
 This group can be generated by two generators, $S$ and $T$, which act as follows
-$S:\tau\mapsto-\frac{1}{\tau},~{}~{}~{}~{}\quad T:\tau\mapsto\tau+1\,,$ (2)
+$S:\tau\mapsto-\frac{1}{\tau},~~~~\quad T:\tau\mapsto\tau+1\,.$ (2)
 These generators satisfy the relations
@@
 The $S,T$ generators can be represented by the following $SL(2,Z)$ matrices,
-$S=\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right),~{}~{}~{}\quad T=\left(\begin{array}[]{cc}1&1\\ 0&1\end{array}\right)\,.$ (4)
+$S=\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right),~~~\quad T=\left(\begin{array}[]{cc}1&1\\ 0&1\end{array}\right)~.$ (4)
 Finite modular groups $\Gamma_{N}$ Feruglio:2017spp are defined as $\Gamma_{N}=\overline{\Gamma}/\overline{\Gamma}(N)$, where $\overline{\Gamma}\equiv SL(2,Z)/\{\pm 1\}$ and $\overline{\Gamma}(N)$ is given by
 $\overline{\Gamma}(N)=\left\{\begin{array}[]{c}\Gamma(N),~~~~~~~~~~~N\geq 3\\ \Gamma(N)/\{\pm 1\},~~~N=2\end{array}\right$.~~~, (7)
-with $\Gamma(N)$ the principal congruence subgroup of level $N$
+with $\Gamma(N)$ the principal congruence subgroup of level NN
 $\Gamma(N)=\left\{\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in SL(2,\mathbb{Z}),~b=c=0\,(\mathrm{mod~N}),a=d=1\,(\mathrm{mod~N})\right\}\,.$ (8)
@@
 ${\cal L}$ $\supseteq$ $\int d^{4}\theta\left[-3\log(2\Im\tau)+\frac{H_{u}^{\dagger}H_{u}}{(2\Im\tau)^{k_{u}}}+\frac{H_{d}^{\dagger}H_{d}}{(2\Im\tau)^{k_{d}}}+\cdots\right]$ (12)
-$+$ $\int d^{2}\theta\left[\alpha\Lambda_{SUSY}\frac{Y_{N}^{k_{r}}(\tau)}{\eta(\tau)^{2k_{3}}}H_{u}H_{d}+\cdots\right],$
+$+$ $\int d^{2}\theta\left[\alpha\Lambda_{SUSY}\frac{Y_{N}^{(k_{r})}(\tau)}{\left[\eta(\tau)\right]^{2k_{3}}}H_{u}H_{d}+\cdots\right],$
 where the coefficient for bilinear $H_{u}H_{d}$ term is given by
-$\tilde{\mu}=\alpha\Lambda_{SUSY}\frac{Y_{N}^{(k_{r})}(\tau)}{\eta(\tau)^{2k_{3}}}~{},$ (13)
+$\tilde{\mu}=\alpha\Lambda_{SUSY}\frac{Y_{N}^{(k_{r})}(\tau)}{\left[\eta(\tau)\right]^{2k_{3}}}~,$ (13)
 for $k_{H_{u}}+k_{H_{d}}=k_{r}-k_{3}+3$ to ensure the modular invariance of
@@
 $\eta(\gamma\tau)=\epsilon_{\eta}(\gamma)(c\tau+d)^{1/2}\eta(\tau)~,~~\forall\gamma\in\overline{\Gamma}~,$ (15)
-where $\epsilon_{\eta}(\gamma)$ depends on $\gamma$ (but not on $\tau$) and can be expressed in terms of the Rademacher $\Phi(\gamma)$ function
-$\epsilon_{\eta}(\gamma)=\exp\left[\frac{2\pi i}{24}\Phi(\gamma)-i\frac{\pi}{4}\right],$ (16)
-that satisfies $\epsilon_{\eta}(\gamma)^{24}=1$ DHoker:2022dxx .
+where $\epsilon_{\eta}(\gamma)$ depends on $\gamma$ (but not on $\tau$) and can be expressed as
+$\epsilon_{\eta}(\gamma)$ $=$ $\exp\left[\frac{\pi i}{12}\omega(a,b,c,d)-i\frac{\pi}{4}\right],$~ (16)
+that satisfies $\epsilon_{\eta}(\gamma)^{24}=1$ DHoker:2022dxx . Here
+$\omega(a,b,c,d)=\left(\frac{a+d}{c}+12s(-d,c)\right),$ (17)
+is an integer and s(d,c)s(-d,c) is a Dedekind sum.
 The expansion of the $\eta(\tau)$ function is
@@
 Thus, when the modulus field is stabilized, the effective $\mu$-parameter is regenerated as
-$\mu_{eff}=\alpha\Lambda_{SUSY}\frac{Y_{N}^{k_{r}}(\tau_{0})}{\eta(\tau_{0})^{2k_{3}}}(2\Im\tau_{0})^{(k_{H_{u}}+k_{H_{d}})/2}=\alpha\Lambda_{SUSY}\frac{Y_{N}^{k_{r}}(\tau_{0})}{\eta(\tau_{0})^{2k_{3}}}(2\Im\tau_{0})^{(k_{r}-k_{3}+3)/2}.$ (19)
+$\mu_{eff}=\alpha\Lambda_{SUSY}\frac{Y_{N}^{(k_{r})}(\tau_{0})}{\left[\eta(\tau_{0})\right]^{2k_{3}}}(2\Im\tau_{0})^{(k_{H_{u}}+k_{H_{d}})/2}=\alpha\Lambda_{SUSY}\frac{Y_{N}^{(k_{r})}(\tau_{0})}{\left[\eta(\tau_{0})\right]^{2k_{3}}}(2\Im\tau_{0})^{(k_{r}-k_{3}+3)/2}.$ (20)
 We need to explain why $\mu_{eff}$ is much smaller than $\Lambda_{SUSY}$ and of order the electroweak scale or soft SUSY breaking scale $m_{soft}$.
@@
 All these suppression factors are determined by the modulus stabilization mechanism, with the stabilized modulus VEVs being affected by SUSY breaking effects.
+The smallness of $\mu_{eff}$ can also be achieved by a weighton-like mechanism King:2020qaj for $H_{u}H_{d}$ bilinear.
 4 Stabilization of modulus field with nilpotent Goldstino ${\textbf{X}}_{nl}$ superfield
-To stabilize the modulus field, we can utilize effective supergravity scalar potentials derived from a weight $-3$ superpotential to determine the minimum.
+To stabilize the modulus field, we can utilize effective supergravity scalar potentials derived from a weight 3-3 superpotential to determine the minimum.
 The modulus VEV can always be restricted to the fundamental domain
@@
 the superpotential for the modulus should take the form
-$W(\tau)=c_{1}^{3}\frac{H(\tau)}{\eta(\tau)^{6}},$ (22)
+$W(\tau)=c_{1}^{3}\frac{H(\tau)}{\left[\eta(\tau)\right]^{6}},$ (23)
 with the Kahler potential
@@
 $H(\tau)_{m,n}$ $=$ $\left(j(\tau)-1728\right)^{m/2}j(\tau)^{n/3}{\cal P}\left(j(\tau)\right),$ (25)
-$\equiv$ $\left(\frac{G_{6}(\tau)}{\eta(\tau)^{12}}\right)^{m}\left(\frac{G_{4}(\tau)}{\eta(\tau)^{12}}\right)^{n}{\cal P}\left(j(\tau)\right)~{},$
-where ${\cal P}\left(j(\tau)\right)$ denotes a polynomial with respect to $j(\tau)$. The Klein $j$-function is defined as a modular form of zero weight
-$j(\tau)$$\equiv$ $\frac{3^{6}5^{3}}{\pi^{12}}\frac{G_{4}(\tau)^{3}}{\eta(\tau)^{24}},$ (25)
+$\equiv$ $\left(\frac{G_{6}(\tau)}{\left[\eta(\tau)\right]^{12}}\right)^{m}\left(\frac{G_{4}(\tau)}{\left[\eta(\tau)\right]^{8}}\right)^{n}{\cal P}\left(j(\tau)\right)~,$
+where ${\cal P}\left(j(\tau)\right)$ denotes a polynomial with respect to $j(\tau)$ and $m,n$ are non-negative integers target:space .
+The definitions of $G_{4}(\tau)$ and $G_{6}(\tau)$ are given in the Appendix B.
+The Klein $j$-function is defined as a modular form of zero weight
+$j(\tau)$ $\equiv$ $\frac{3^{6}5^{3}}{\pi^{12}}\frac{G_{4}(\tau)^{3}}{\left[\eta(\tau)\right]^{24}},$ (26)
 $\approx$ $744+q^{-1}+196884q+21493760q^{2}+864299970q^{3}+{\cal O}(q^{4})~,$
 with $j(i\infty)$ diverges, $j(\omega)=0$ and $j(i)=1728$.
+Note that $j(\tau)$ has a triple zero at $\tau=\omega$, and $j(\tau)-1728$ has a double zero at $\tau=i$.
+The derivative $j^{\prime}(\tau)$ vanishes only at these values: $j^{\prime}(\omega)=j^{\prime}(i)=0$ Klein:jfunction .
 We should note that, in such a stabilization mechanism, the fixed points $z=i,\omega\equiv e^{i2\pi/3}$ are always the extrema of the scalar potential, although they may not be the global minimum.
@@
 The chiral superfield $\textbf{X}_{nl}$ can be seen as the long-distance expression of the microscopic superfield X that satisfies $\bar{D}^{\dot{\alpha}}{\cal J}_{\alpha\dot{\alpha}}=D_{\alpha}\textbf{X}$ for Ferrara-Zumino multiplet ${\cal J}_{\alpha\dot{\alpha}}$ KS , which can be used even when SUSY is spontaneously broken and can be followed consistently along the renormalization group flow down to the deep IR.
-As noted in Aparicio:2015psl , for the anti-D3 brane in the KKLT scenario, the representation in terms of $\textbf{X}_{nl}$ is very convenient and enough to capture all the anti-D3 brane degrees of freedoms since it allows to treat the effect of anti-D3 brane in terms of standard supergravity couplings of matter and moduli superfields to the nilpotent Goldstino superfield.
+As noted in Aparicio:2015psl , for the anti-D3 brane in the KKLT Kachru:2003aw scenario, the representation in terms of $\textbf{X}_{nl}$ is very convenient and enough to capture all the anti-D3 brane degrees of freedoms since it allows to treat the effect of anti-D3 brane in terms of standard supergravity couplings of matter and moduli superfields to the nilpotent Goldstino superfield.
 From the perspective of low energy effective field theory, regardless of the UV origin of SUSY breaking, such a non-linear realization scheme is also convenient for discussions on modulus stabilization and vacuum uplifting.
 When coupled to supergravity, the most general form (at the two-derivative level) for the action in superspace is given by
-[…]
+${\cal L}\supseteq\int d^{2}\Theta 2{\cal{E}}\left[\frac{3}{8}\left(\overline{\cal{D}}^{2}-8{\cal R}\right)\,e^{-K/3}+W\right]+\hbox{h.c.}\,$, (29)
 with
 $K=\textbf{X}_{nl}^{\dagger}\textbf{X}_{nl}~,~~~W=f_{0}\textbf{X}_{nl}+c_{1}^{3}~,$ (30)
-where without loss of generality we take $f_{0}$ and $c_{1}$ to be real.
-Integrating out the auxillary field $F_{X}$ gives $F_{X}=-(W_{\textbf{X}_{nl}})^{*}=-f_{0}$.
-To accommodate the modular flavor symmetry, the Kahler potential and superpotential needs to be adapted to
+[…]
+Integrating out the auxillary field $F_{X}$ gives $F_{X}=-(W_{\textbf{X}_{nl}})^{*}=-f_{0}$. To accommodate the modular flavor symmetry, the Kahler potential and superpotential needs to be adapted to
 $K$ $\supseteq$ $-3\log(2\Im\tau)+\frac{\textbf{X}^{\dagger}_{nl}\textbf{X}_{nl}}{(2\Im\tau)^{k_{X}}}~,$
-$W$ $\supseteq$ $\left(f_{0}\textbf{X}_{nl}\left[\eta(\tau)\right]^{2k_{X}}+c_{1}^{3}\right)\frac{H(\tau)}{\eta^{6}(\tau)}.$ (30)
-The supergravity scalar potential can be derived from
+$W$ $\supseteq$ $\left(f_{0}\textbf{X}_{nl}\left[\eta(\tau)\right]^{2k_{X}}+c_{1}^{3}\right)\frac{H(\tau)}{[\eta(\tau)]^{6}}.$ (31)
+To ensure the modular invariance of the superpotential, the modular weight $k_{X}$ should satisfy $k_{X}\in 12\mathbb{Z}$. The supergravity scalar potential can be derived from
 $V=e^{K}\left[K^{\alpha\bar{\beta}}\nabla_{\alpha}W\nabla_{\bar{\beta}}W-3W\overline{W}\right],$ (32)
@@
 From the Kahler potential, we can calculate the inverse Kahler metric
-$K^{{\alpha}\bar{\beta}}$ $=$ $\left(\begin{array}[]{cc}\frac{(2\Im\tau)^{2}}{3}\left[1-\frac{k_{X}}{3}\frac{\textbf{X}^{\dagger}_{nl}\textbf{X}_{nl}}{(2\Im\tau)^{k_{X}}}\right]&\frac{i(2\Im\tau)k_{X}\textbf{X}_{nl}}{3}\\ -\frac{i(2\Im\tau)k_{X}\textbf{X}^{\dagger}_{nl}}{3}&(2\Im\tau)^{k_{X}}\left[1+\frac{k_{X}^{2}(\textbf{X}^{\dagger}_{nl}\textbf{X}_{nl})}{3(2\Im\tau)^{k_{X}}}\right]\end{array}\right)~{}.$ (35)
-Here the indices $\alpha,\beta\sim(\tau,\textbf{X}_{nl})$ and the nilpotency of $\textbf{X}_{nl}$ is used in the calculation.
-The scalar potential for the modulus field $\tau$ and $X_{nl}$ can be calculated as
+$K^{{\alpha}\bar{\beta}}$ $=$ $\left(\begin{array}[]{cc}\frac{(2\Im\tau)^{2}}{3}\left[1-\frac{k_{X}}{3}\frac{{X}^{*}_{nl}{X}_{nl}}{(2\Im\tau)^{k_{X}}}\right]&\frac{i(2\Im\tau)k_{X}{X}_{nl}}{3}\\ -\frac{i(2\Im\tau)k_{X}{X}^{*}_{nl}}{3}&(2\Im\tau)^{k_{X}}\left[1+\frac{k_{X}^{2}({X}^{*}_{nl}{X}_{nl})}{3(2\Im\tau)^{k_{X}}}\right]\end{array}\right)~.$ (36)
+Here the indices $\alpha,\beta\sim(\tau,{X}_{nl})$ and the nilpotency of ${X}_{nl}$ is used in the calculation.
+From the general expression of the SUGRA scalar potential in (32), the scalar potential for the modulus field $\tau$ and $X_{nl}$ can be calculated as
 $V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})$ (37)
@@
 $+$ $2\Re\left[\frac{i(2\Im\tau)k_{X}X_{nl}}{3}(\tilde{W}_{X}H(\tau))^{*}\left(\tilde{W}_{M}\tilde{c}_{M}+H(\tau)\tilde{W}_{T}\right)\right]$
--- $\left.3\left|\tilde{c}_{W}\right|^{2}\left|{H(\tau)}\right|^{2}\right\}~{},$
+- $\left.3\left|\tilde{c}_{M}\right|^{2}\left|{H(\tau)}\right|^{2}\right\}~,$
 where
 $\tilde{W}_{M}$ $\equiv$ $H^{\prime}(\tau)-\frac{3iH(\tau)}{2\pi}\hat{G}_{2}(\tau,\tau^{*})~,$
-$\tilde{W}_{T}$$\equiv$ $k_{X}f_{0}X_{nl}\left[\eta(\tau)\right]^{2k_{X}}\left[\frac{i}{2\pi}\hat{G}_{2}(\tau,\tau^{*})+\frac{i}{2\Im\tau}+\frac{ic_{1}^{3}X_{nl}^{*}}{f_{0}(2\Im\tau)^{k_{X}+1}\eta(\tau)^{2k_{X}}}\right]~{},$
-$\tilde{c}_{M}$ $=$ $f_{0}X_{nl}\left[\eta(\tau)\right]^{2k_{X}}+c_{1}^{3}~{},$
-$\tilde{W}_{X}$ $=$ $f_{0}\eta(\tau)^{2k_{X}}\left(1+\frac{|X_{nl}|^{2}}{2(\Im\tau)^{k_{X}}}\right)+\frac{{X}^{*}_{nl}c_{1}^{3}}{(2\Im\tau)^{k_{X}}}.$ (37)
+$\tilde{W}_{T}$ $\equiv$ $k_{X}f_{0}X_{nl}\left[\eta(\tau)\right]^{2k_{X}}\left[\frac{i}{2\pi}\hat{G}_{2}(\tau,\tau^{*})+\frac{i}{2\Im\tau}+\frac{ic_{1}^{3}X_{nl}^{*}}{f_{0}(2\Im\tau)^{k_{X}+1}\left[\eta(\tau)\right]^{2k_{X}}}\right]~,$
+$\tilde{c}_{M}$ $\equiv$ $f_{0}X_{nl}\left[\eta(\tau)\right]^{2k_{X}}+c_{1}^{3}~,$
+$\tilde{W}_{X}$ $\equiv$ $f_{0}\left[\eta(\tau)\right]^{2k_{X}}\left(1+\frac{|X_{nl}|^{2}}{(2\Im\tau)^{k_{X}}}\right)+\frac{{X}^{*}_{nl}c_{1}^{3}}{(2\Im\tau)^{k_{X}}}.$ (38)
 The constraint $X_{nl}^{2}=0$ for the lowest component of $\textbf{X}_{nl}$ is used to eliminate some of the high order $X_{nl}$ terms.
 We also use the following expression for the derivative of Dedekind $\eta$ function
-$\frac{\eta^{\prime}(\tau)}{\eta(\tau)}=\frac{i}{4\pi}G_{2}~{},$ (38)
+$\frac{\eta^{\prime}(\tau)}{\eta(\tau)}=\frac{i}{4\pi}G_{2}(\tau)~,$ (39)
 where the holomorphic Eisenstein series $G_{2}$, which is a quasi-modular form
@@
 $\hat{G}_{2}(\tau,{\tau}^{*})=G_{2}(\tau)-\frac{\pi}{\Im\tau}~.$ (41)
-The generic complicate form of the scalar potential may obscure the discussions on the stabilization mechanism. Therefore, we would like to concentrate on its typical simplification forms
-  • Scenario with $k_{X}=0$:
-    When $k_{X}=0$, the scalar potential simplifies significantly.
+From the definition of $H(\tau)$, its derivative with respect to $\tau$ can be calculated to be
+$H^{\prime}(\tau)$ $=$ $\left\{j(\tau)^{n/3}[j(\tau)-1728]^{m/2}\mathcal{P}(j(\tau))\right\}^{\prime}~,$ (42)
+$=$ $\left\{\frac{n}{3}j(\tau)^{n/3-1}[j(\tau)-1728]^{m/2}\mathcal{P}(j(\tau))+\frac{m}{2}j(\tau)^{n/3}[j(\tau)-1728]^{m/2-1}\mathcal{P}(j(\tau))\right$.
+$+$ $\left.j(\tau)^{n/3}[j(\tau)-1728]^{m/2}\mathcal{P}^{\prime}(j(\tau))\right\}j^{\prime}(\tau)~.$
+For non-vanishing values of $H(\tau)$, we have
+$H^{\prime}(\tau)$ $=$ $H(\tau)j^{\prime}(\tau)\left[\frac{n}{3}\frac{1}{j(\tau)}+\frac{m}{2}\frac{1}{j(\tau)-1728}+\frac{1}{\mathcal{P}(j(\tau))}\frac{\partial\mathcal{P}(j(\tau))}{\partial j(\tau)}\right],$ (43)
+making obvious that $H(\tau)$ can be factored out from $\tilde{W}_{M}$ as
+$\tilde{W}_{M}$ $=$ $H(\tau)\left(j^{\prime}(\tau)\left[\frac{n}{3}\frac{1}{j(\tau)}+\frac{m}{2}\frac{1}{j(\tau)-1728}+\frac{1}{\mathcal{P}(j(\tau))}\frac{\partial\mathcal{P}(j(\tau))}{\partial j(\tau)}\right]-\frac{3i}{2\pi}\hat{G}_{2}(\tau,\tau^{*})\right).$
+Substituting the factorized form of $\tilde{W}_{M}$ back into the scalar potential, it becomes apparent that the $|H(\tau)|^{2}$ factor can be extracted from the expression when $H(\tau)\neq 0$.
+This observation plays an important role in locating the extrema of the scalar potential.
+For $m=n=0$, the trivial case
+$H^{\prime}(\tau)$ $=$ $\mathcal{P}^{\prime}(j(\tau))j^{\prime}(\tau)~,$ (44)
+vanishes at $z=\omega,i$ because $j^{\prime}(\omega)=j^{\prime}(i)=0$.
+For $m=1$ or $n=1,2$, from the results
+$\frac{j^{\prime}(i)}{[j(i)-1728]^{1/2}}$ $=$ $-i2\pi\frac{[E_{4}(i)]^{2}}{[\eta(i)]^{12}},~~\frac{j^{\prime}(\omega)}{[j(\omega)]^{1/3}}=-i2\pi\frac{E_{4}(\omega)E_{6}(\omega)}{[\eta(\omega)]^{16}}=0~,$
+$~\frac{j^{\prime}(\omega)}{[j(\omega)]^{2/3}}$ $=$ $-i2\pi\frac{E_{6}(\omega)}{[\eta(\omega)]^{8}}~,$ (45)
+we can see that $H^{\prime}(i)$ and $H^{\prime}(\tau)$ are finite with such choices of $m,n$, within which we use
+$\frac{\partial}{\partial\tau}j(\tau)=-i2\pi\frac{E_{6}(\tau)}{E_{4}(\tau)}j(\tau)=-i2\pi\frac{E_{6}(\tau)[E_{4}(\tau)]^{2}}{[\eta(\tau)]^{24}},$ (46)
+and the relations target:space
+$j(\tau)-1728=\left(\frac{E_{6}(\tau)}{[\eta(\tau)]^{12}}\right)^{2}~,~~~j(\tau)=\left(\frac{E_{4}(\tau)}{[\eta(\tau)]^{8}}\right)^{3}~,$ (47)
+as well as the values of $\eta$-function at $z=i,\omega$
+$\eta(i)=\frac{\Gamma(\frac{1}{4})}{2\pi^{3/4}}~,~~~~\eta(\omega)=e^{-i\frac{\pi}{24}}\frac{3^{\frac{1}{8}}[\Gamma(\frac{1}{3})]^{\frac{3}{2}}}{2\pi}~,$ (48)
+and some special value for $E_{4}(\tau)$ and $E_{6}(\tau)$ DHoker:2022dxx
+$E_{4}(i)=\frac{48[\Gamma(\frac{5}{4})]^{4}}{\pi^{2}[\Gamma(\frac{3}{4})]^{4}},~~E_{4}(\omega)=0,~~E_{6}(\omega)=\frac{729}{2\pi^{3}}\frac{[\Gamma(\frac{4}{3})]^{6}}{[\Gamma(\frac{5}{6})]^{6}}.$ (49)
+For $m>1$ and $n>2$, it is obvious that $H^{\prime}(\tau)$ is not singular at zero points of $H(\tau)$, therefore vanishes at $i$ and $\omega$, as $j^{\prime}(i)=j^{\prime}(\omega)=0$.
+So, we have
+$H^{\prime}(i)$ $=$ $\left\{\begin{array}[]{c}-i\pi(12)^{n}\frac{[E_{4}(i)]^{2}}{[\eta(i)]^{12}}P(1728),~~~m=1\\ 0,~~~~~~~~~~~~~~~~\quad~\quad\quad\quad\quad~~m\neq 1\end{array}\right$. (52)
+$H^{\prime}(\omega)$ $=$ $\left\{\begin{array}[]{c}-\frac{i2\pi}{3}(-12)^{3m/2}\frac{E_{6}(\omega)}{[\eta(\omega)]^{8}}P(0),~~~n=1\\ 0,~~~~~~~~~~~~~~~~~\quad\quad\quad\quad\quad~~n\neq 1\end{array}\right.$~.
+(55)
+The condition $H(\tau)=0$ can be reduced to
+$j(\tau)^{n/3}=0,\quad\text{or}\quad[j(\tau)-1728]^{m/2}=0,\quad\text{or}\quad\mathcal{P}(j(\tau))=0.$ (56)
+The corresponding solutions are $\tau=i,\omega$ (for $m,n\neq 0$) and $\tau=\tilde{\tau}_{i}$ that satisfies $P(j(\tilde{\tau}_{i}))=0$.
+Whereas $H^{\prime}(\tau)$ typically vanishes at $\tau=i$ and $\tau=\omega$ (barring special choices of $m$ and $n$), it generally does not vanish at $\tau=\tilde{\tau}_{i}$, the point where $P(j(\tilde{\tau}_{i}))=0$.
+However, for some special choices of $)P(j(\tilde{\tau}$)), $H^{\prime}(\tau)$ can still vanish at $\tau=\tilde{\tau}_{0}$.
+For example, when
+$P(j(\tau))=\prod\limits_{i}(j(\tau)-j(\tilde{\tau}_{i}))^{2}~,$ (57)
+then $\tau=\tilde{\tau}_{0}$ acts as a multiple root of $P(j(\tilde{\tau}))=0$ and $H^{\prime}(\tau)$ vanishes at such roots. We will discuss such interesting possibilities in the next subsection.
+The generic, complicated form of the scalar potential may obscure the discussions regarding the stabilization mechanism. Therefore, we would like to concentrate on its typical simplification forms.
+4.1 The case with $k_{X}=0$
+When $k_{X}=0$, the scalar potential can be simplified significantly.
+It is evident that, due to the presence of the $K_{\alpha}W$ terms within the Kahler-covariant derivative of the superpotential $\nabla_{\alpha}W$, the scalar potential still includes non-trivial interference terms between the $X_{nl}$ sector and the modulus field $\tau$ sector, despite the fact that the Kahler metric $K^{\alpha\bar{\beta}}$ is diagonal in this case.
 The potential can then be written as
@@
 $+$ $\left|H(\tau)\right|^{2}\left(f^{2}_{0}+2f_{0}^{2}{|X_{nl}|^{2}}+2\Re\left[f_{0}c_{1}^{3}{X}_{nl}^{*}\right]+|X_{nl}|^{2}c_{1}^{6}\frac{}{}\right)$~
-    -- $\left.3\left|{H(\tau)}\right|^{2}\left(f_{0}^{2}|X_{nl}|^{2}+c_{1}^{6}+2\Re\left[f_{0}X_{nl}c_{1}^{3}\right]\right)\right\}~{}$
+- $\left.3\left|{H(\tau)}\right|^{2}\left(f_{0}^{2}|X_{nl}|^{2}+c_{1}^{6}+2\Re\left[f_{0}X_{nl}c_{1}^{3}\right]\right)\right\}$~
 $+$ $\frac{|X_{nl}|^{2}}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}c_{1}^{6}+\left|H(\tau)\right|^{2}f^{2}_{0}-3c_{1}^{6}\left|{H(\tau)}\right|^{2}\right\}~,$
-    where
-    $\tilde{W}_{M}$$\equiv$ $H^{\prime}(\tau)-\frac{3iH(\tau)}{2\pi}\hat{G}_{2}(\tau,\tau^{*})~{},$
-    $\tilde{c}_{M}$ $=$ $f_{0}X_{nl}+c_{1}^{3}~{}.$ (42)
-    Note the constraints $X_{nl}^{2}=({X}^{*}_{nl})^{2}=0$ are used in the scalar potential.
+where we adopt
+$\tilde{W}_{M}$ $=$ $H^{\prime}(\tau)-\frac{3iH(\tau)}{2\pi}\hat{G}_{2}(\tau,\tau^{*})~,$
+$\tilde{c}_{M}$ $=$ $f_{0}X_{nl}+c_{1}^{3}~,$ (59)
+from the definitions in (38).
 In this scenario, the modulus VEV can be determined by the extreme conditions of the scalar potential
-    $\frac{\partial}{\partial\tau}V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})=\frac{\partial}{\partial\tau^{*}}V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})=0~{},$
-    and the minimum condition can be determined by the positive definiteness of Hessian matrix in terms of
-    $\frac{\partial^{2}}{\partial x^{2}}V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})$ $=$ $2\left(\frac{\partial^{2}}{\partial\tau\partial\tau^{*}}+\Re\frac{\partial^{2}}{\partial\tau\partial\tau}\right)V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*}),$
-    $\frac{\partial^{2}}{\partial y^{2}}V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})$ $=$ $2\left(\frac{\partial^{2}}{\partial\tau\partial\tau^{*}}-\Re\frac{\partial^{2}}{\partial\tau\partial\tau}\right)V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*}),$
-    $\frac{\partial^{2}}{\partial x\partial y}V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})$ $=$ $-2\Im\frac{\partial^{2}}{\partial\tau\partial\tau}V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*}),$ (43)
-    where $\tau\equiv x+iy\equiv\rho_{\tau}e^{i\theta_{\tau}}$.
-It was argued in target:space ; Ding:2024neh ; Higaki:2024pql that the modular invariance and $\tau\leftrightarrow-\tau^{*}$ symmetry of the scalar potential ensure that the first derivative of scalar potential along certain directions at the boundary of the fundamental domain vanishes
-    $\left.\frac{\partial}{\partial x}V\right|_{x=\pm 1/2,0}=0~{},~{}~{}~{}\left.\frac{\partial}{\partial\rho_{\tau}}V\right|_{\rho_{\tau}=1}=0~{}.~{}~{}~{}$ (44)
-    We then determine the extreme conditions
+$\frac{\partial}{\partial\tau}V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})=\frac{\partial}{\partial\tau^{*}}V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})=0~,$ (60)
+$\frac{\partial}{\partial X_{nl}}V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})=\frac{\partial}{\partial X_{nl}^{*}}V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})=0~.$ (61)
+Given that the scalar potential is real, the two equations on the right follow directly from those on the left via complex conjugation.
+To determine if a local extremum is a minimum, we need to assess the positive definiteness of the $4\times 4$ Hessian matrix
+$(H_{f})_{ij}\equiv\frac{\partial^{2}V}{\partial x_{i}\partial x_{j}}~,~~~{\rm for}~~x_{i}={\Re}\tau,{\Im}\tau,{\Re}X_{nl},{\Im}X_{nl}~,$ (62)
+at that extremum. The elements of the Hessian matrix can be rewritten by
+$\frac{\partial^{2}}{\partial x_{1}^{2}}V$ $=$ $2\left(\frac{\partial^{2}}{\partial\tau\partial\tau^{*}}+\Re\frac{\partial^{2}}{\partial\tau\partial\tau}\right)V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*}),$
+$\frac{\partial^{2}}{\partial x_{2}^{2}}V$ $=$ $2\left(\frac{\partial^{2}}{\partial\tau\partial\tau^{*}}-\Re\frac{\partial^{2}}{\partial\tau\partial\tau}\right)V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*}),$
+$\frac{\partial^{2}}{\partial x_{1}\partial x_{2}}V$ $=$ $-2\Im\frac{\partial^{2}}{\partial\tau\partial\tau}V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*}),$ (63)
+with similar terms for $X_{nl}$ and $X_{nl}^{*}$ in terms of $\rho$ and $\theta$
+$\frac{\partial^{2}}{\partial x_{3}^{2}}V$ $=$ $2\left(\frac{\partial^{2}}{\partial X_{nl}\partial X_{nl}^{*}}+\Re\frac{\partial^{2}}{\partial X_{nl}\partial X_{nl}}\right)V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*}),$
+$=$ $\left(\cos^{2}\theta\frac{\partial^{2}}{\partial\rho^{2}}-2\frac{\cos\theta\sin\theta}{\rho}\frac{\partial^{2}}{\partial\rho\partial\theta}+\frac{\sin^{2}\theta}{\rho^{2}}\frac{\partial^{2}}{\partial\theta^{2}}+\frac{\sin^{2}\theta}{\rho}\frac{\partial}{\partial\rho}+2\frac{\cos\theta\sin\theta}{\rho^{2}}\frac{\partial}{\partial\theta}\right)V~,$
+$\frac{\partial^{2}}{\partial x_{4}^{2}}V$ $=$ $2\left(\frac{\partial^{2}}{\partial X_{nl}\partial X_{nl}^{*}}-\Re\frac{\partial^{2}}{\partial X_{nl}\partial X_{nl}}\right)V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*}),$
+$=$ $\left(\sin^{2}\theta\frac{\partial^{2}}{\partial\rho^{2}}+2\frac{\sin\theta\cos\theta}{\rho}\frac{\partial^{2}}{\partial\rho\partial\theta}+\frac{\cos^{2}\theta}{\rho^{2}}\frac{\partial^{2}}{\partial\theta^{2}}+\frac{\cos^{2}\theta}{\rho}\frac{\partial}{\partial\rho}-2\frac{\sin\theta\cos\theta}{\rho^{2}}\frac{\partial}{\partial\theta}\right)V~,$
+$\frac{\partial^{2}}{\partial x_{3}\partial x_{4}}V$ $=$ $-2\Im\frac{\partial^{2}}{\partial X_{nl}\partial X_{nl}}V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*}),$
+$=$ $\left(\sin\theta\cos\theta\frac{\partial^{2}}{\partial\rho^{2}}+\frac{\cos 2\theta}{\rho}\frac{\partial^{2}}{\partial\rho\partial\theta}-\frac{\sin\theta\cos\theta}{\rho^{2}}\frac{\partial^{2}}{\partial\theta^{2}}-\frac{\sin\theta\cos\theta}{\rho}\frac{\partial}{\partial\rho}-\frac{\cos 2\theta}{\rho^{2}}\frac{\partial}{\partial\theta}\right)V~.$
+Other second derivatives of $V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})$ with respect to $\tau,\tau^{*}$ and $\rho,\theta$ are given by
+$\frac{\partial^{2}}{\partial x_{1}\partial x_{3}}V$ $=$ $\left(\frac{\partial^{2}}{\partial\tau\partial X_{nl}}+\frac{\partial^{2}}{\partial\tau^{*}\partial X_{nl}^{*}}+\frac{\partial^{2}}{\partial\tau\partial X_{nl}^{*}}+\frac{\partial^{2}}{\partial\tau^{*}\partial X_{nl}}\right)V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})~,$
+$=$ $\left[\cos\theta\left(\frac{\partial^{2}}{\partial\tau\partial\rho}+\frac{\partial^{2}}{\partial\tau^{*}\partial\rho}\right)-\frac{\sin\theta}{\rho}\left(\frac{\partial^{2}}{\partial\tau\partial\theta}+\frac{\partial^{2}}{\partial\tau^{*}\partial\theta}\right)\right]V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})~,$
+$\frac{\partial^{2}}{\partial x_{2}\partial x_{4}}V$ $=$ $-\left(\frac{\partial^{2}}{\partial\tau\partial X_{nl}}+\frac{\partial^{2}}{\partial\tau^{*}\partial X_{nl}^{*}}-\frac{\partial^{2}}{\partial\tau\partial X_{nl}^{*}}-\frac{\partial^{2}}{\partial\tau^{*}\partial X_{nl}}\right)V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})~,$
+$=$ $-\left[i\sin\theta\left(\frac{\partial^{2}}{\partial\tau\partial\rho}-\frac{\partial^{2}}{\partial\tau^{*}\partial\rho}\right)+i\frac{\cos\theta}{\rho}\left(\frac{\partial^{2}}{\partial\tau\partial\theta}-\frac{\partial^{2}}{\partial\tau^{*}\partial\theta}\right)\right]V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})~,$
+$\frac{\partial^{2}}{\partial x_{1}\partial x_{4}}V$ $=$ $-i\left(\frac{\partial^{2}}{\partial\tau\partial X_{nl}}-\frac{\partial^{2}}{\partial\tau^{*}\partial X_{nl}^{*}}+\frac{\partial^{2}}{\partial\tau\partial X_{nl}^{*}}-\frac{\partial^{2}}{\partial\tau^{*}\partial X_{nl}}\right)V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})~,$
+$=$ $\left[\sin\theta\left(\frac{\partial^{2}}{\partial\tau\partial\rho}+\frac{\partial^{2}}{\partial\tau^{*}\partial\rho}\right)+\frac{\cos\theta}{\rho}\left(\frac{\partial^{2}}{\partial\tau\partial\theta}+\frac{\partial^{2}}{\partial\tau^{*}\partial\theta}\right)\right]V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})~,$
+$\frac{\partial^{2}}{\partial x_{2}\partial x_{3}}V$ $=$ $i\left(\frac{\partial^{2}}{\partial\tau\partial X_{nl}}-\frac{\partial^{2}}{\partial\tau^{*}\partial X_{nl}^{*}}-\frac{\partial^{2}}{\partial\tau\partial X_{nl}^{*}}+\frac{\partial^{2}}{\partial\tau^{*}\partial X_{nl}}\right)V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})~,$
+$=$ $\left[i\cos\theta\left(\frac{\partial^{2}}{\partial\tau\partial\rho}-\frac{\partial^{2}}{\partial\tau^{*}\partial\rho}\right)-i\frac{\sin\theta}{\rho}\left(\frac{\partial^{2}}{\partial\tau\partial\theta}-\frac{\partial^{2}}{\partial\tau^{*}\partial\theta}\right)\right]V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})~.$
+The analytical expression for each element of the Hessian matrix is highly complex and only takes a simplified form at specific points, such as the fixed points $\omega$ and $i$.
+Since this study does not focus on the fixed points, we evaluate the positivity of the eigenvalues for the Hessian matrix at the extrema numerically, rather than through analysis of its analytical form, in the subsequent discussions.
+The extreme condition (60) $V_{\tau}=0$ can be calculated as
+$\frac{\partial}{\partial\tau}V$ $=$ $\frac{\frac{3i}{2\pi}\hat{G}_{2}(\tau,\tau^{*})}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\tilde{c}_{M}\right|^{2}\right$.
+(66)
+$+$ $\left.\left|H(\tau)\right|^{2}\left|f_{0}\left(1+{|X_{nl}|^{2}}\right)+{{X}^{*}_{nl}c_{1}^{3}}\right|^{2}-3\left|\tilde{c}_{M}\right|^{2}\left|{H(\tau)}\right|^{2}\right\}$
+$+$ $\frac{1}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{\frac{-4i\Im\tau}{3}\left|\tilde{W}_{M}\tilde{c}_{M}\right|^{2}+\frac{(2\Im\tau)^{2}}{3}\left|\tilde{c}_{M}\right|^{2}\tilde{W}_{M}^{*}\tilde{W}_{M}^{\prime}\right$.
+$+$ $\left.H(\tau)^{*}H^{\prime}(\tau)\left(\left|f_{0}\left(1+{|X_{nl}|^{2}}\right)+{{X}^{*}_{nl}c_{1}^{3}}\right|^{2}-3\left|\tilde{c}_{M}\right|^{2}\right)\right\}$
+$+$ $\frac{\frac{3i}{2\pi}\hat{G}_{2}(\tau,\tau^{*})|X_{nl}|^{2}}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\tilde{c}_{M}\right|^{2}+\left|H(\tau)\right|^{2}\left(f_{0}^{2}-3\left|\tilde{c}_{M}\right|^{2}\right)\right\}$
+$+$ $\frac{{|X_{nl}|^{2}}}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{\frac{-4i\Im\tau}{3}\left|\tilde{W}_{M}\tilde{c}_{M}\right|^{2}+\frac{(2\Im\tau)^{2}}{3}\left|\tilde{c}_{M}\right|^{2}\tilde{W}_{M}^{*}\tilde{W}_{M}^{\prime}\right$.
+$+$ $\left.H(\tau)^{*}H^{\prime}(\tau)\left(f_{0}^{2}-3\left|\tilde{c}_{M}\right|^{2}\right)\right\},$
+and its complex conjugate for $V_{\tau^{*}}$.
+Within this expression, $\tilde{W}_{M}^{\prime}$ is given by
+$\tilde{W}_{M}^{\prime}$ $=$ $H^{\prime\prime}(\tau)-\frac{3iH^{\prime}(\tau)}{2\pi}\hat{G}_{2}(\tau,\tau^{*})-\frac{3iH(\tau)}{2\pi}\hat{G}^{\prime}_{2}(\tau,\tau^{*})~.$ (67)
+When the value of $\tau$ satisfies $H(\tau)\neq 0$, $\tilde{W}_{M}^{\prime}$ can be rewritten as
+$\tilde{W}_{M}^{\prime}$ $=$ $H(\tau)[j^{\prime}(\tau)]^{2}\left[\frac{n}{3}\frac{1}{j(\tau)}+\frac{m}{2}\frac{1}{j(\tau)-1728}+\frac{1}{\mathcal{P}(j(\tau))}\frac{\partial\mathcal{P}(j(\tau))}{\partial j(\tau)}\right]^{2}$ (68)
+$+$ $H(\tau)\left\{j^{\prime}(\tau)\left[\frac{n}{3}\frac{1}{j(\tau)}+\frac{m}{2}\frac{1}{j(\tau)-1728}+\frac{1}{\mathcal{P}(j(\tau))}\frac{\partial\mathcal{P}(j(\tau))}{\partial j(\tau)}\right]\right\}^{\prime}$~
+- $H(\tau)\frac{3i}{2\pi}\hat{G}_{2}(\tau,\tau^{*})[j^{\prime}(\tau)]^{2}\left[\frac{n}{3}\frac{1}{j(\tau)}+\frac{m}{2}\frac{1}{j(\tau)-1728}+\frac{1}{\mathcal{P}(j(\tau))}\frac{\partial\mathcal{P}(j(\tau))}{\partial j(\tau)}\right]$
+- $\frac{3iH(\tau)}{2\pi}\hat{G}^{\prime}_{2}(\tau,\tau^{*})~,$
+from (42). The derivative of $\hat{G}^{\prime}_{2}(\tau,\tau^{*})$ with respect to $\tau$ is given as follows
+$\hat{G}^{\prime}_{2}(\tau,\tau^{*})=\frac{i\pi^{3}}{18}\left(E_{2}(\tau)^{2}-E_{4}(\tau)\right)-i\frac{2\pi}{(2\Im\tau)^{2}}~.$ (69)
+From equations (4) and (68), one can check that $|H(\tau)|^{2}$ acts as a positive multiplicative factor in $V_{\tau}$ when $H(\tau)\neq 0$.
+In addition to the solutions of $V_{\tau}/|H(\tau)|^{2}=0$ (for $H(\tau)\neq 0$), the constraints $\tilde{W}_{M}^{\prime}=0$ can be satisfied also by some solutions of $H(\tau)=0$, which we need to check point by point.
+From equation (42), we have the following discussions on the condition $H^{\prime}(\tau)=0$ when $H(\tau)=0$:
+  • The case $\tau=i$ is a solution of $H(\tau)=0$:
+    When $\tau=i$ is a solution of $H(\tau)=0$ for $m>0$, the condition $H^{\prime}(\tau)=0$ can be satisfied automatically for $m\neq 1$, see equation (55).
+It can also be satisfied when $\mathcal{P}(1728)=0$ for $m=1$, requiring $P(j(\tau))$ to take the form
+    $P(j(\tau))\propto(j(\tau)-1728)^{k}~,~~~~k\in\mathbb{N}~.$ (70)
+    Such factor of $P(j(\tau))$ can be absorbed into $(j(\tau)-1728)^{m/2}$ term in $H(\tau)$ after the redefinition of ’$m$’, rendering the redefined $\mathcal{P}(1728)\neq 0$.
+So, we conclude that $H^{\prime}(i)=0$ can be satisfied for $m>1$.
+Note that $H^{\prime}(i)=0$ for $m=0$, although $\tau=i$ is in general not a solution of $H(\tau)=0$ when $m=0$.
+  • The case $\tau=\omega$ is a solution of $H(\tau)=0$:
+    Similarly, when $\tau=\omega$ is a solution of $H(\tau)=0$ for $n>0$, the condition $H^{\prime}(\tau)=0$ can be satisfied automatically for $n\neq 1$.
+It can also be satisfied when $\mathcal{P}(0)=0$ for $n=1$, which requires $P(j(\tau))$ to take the form
+    $P(j(\tau))\propto[j(\tau)]^{k}~,~~~~k\in\mathbb{N}~.$ (71)
+    After the redefinition of ’$n$’, we conclude that the condition $H^{\prime}(\omega)=0$ can be satisfied for $n>1$.
+Note also $H^{\prime}(\omega)=0$ when $n=0$, although $\tau=\omega$ is in general not a solution of $H(\tau)=0$ when $n=0$..
+  • The case $\tau=\tilde{\tau}_{i}$ with $P(j(\tilde{\tau}_{i}))=0$:
+    The condition $H^{\prime}(\tau)=0$ can be satisfied for $P^{\prime}(j(\tilde{\tau}_{i}))=0$, which requires $\tau=\tilde{\tau}_{i}$ being a root of $P(j(\tau))$ with multiplicity two or greater.
+Therefore,
+    $P(j(\tau))\propto\pm g(j(\tau))\prod\limits_{i}(j(\tau)-j(\tilde{\tau}_{i}))^{k_{i}}~,~~~~k_{i}\in\mathbb{N}~.$ (72)
+    with $k_{i}\geq 2$ and $g(j(\tau))$ an arbitrary positive polynomial function of $j(\tau)$.
+Before we discuss the condition $H^{\prime\prime}(\tau)=0$ when $H(\tau)=0$, we need to know the values of $H^{\prime\prime}(\tau)$ at the fixed points. It can be calculated that
+$H^{\prime\prime}(i)$ $=$ $\left\{\begin{array}[]{c}-\frac{2\pi^{2}}{[\eta(i)]^{24}}[E_{4}(i)]^{4}\left[\frac{n}{3}P(1728)(12)^{n-3}+(12)^{n}P^{\prime}(1728)\right]~,~~~~m=0\\ ~0,~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~m=1\\ -\frac{2\pi^{2}}{[\eta(i)]^{24}}[E_{4}(i)]^{4}(12)^{n}P(1728)~,~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~m=2\\ ~0,~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~m>2\end{array}\right$.
+(77)
+$H^{\prime\prime}(\omega)$ $=$ $\left\{\begin{array}[]{c}0~,~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~n=0,1,3,4,5\\ -\frac{8\pi^{2}}{9}\frac{[E_{6}(\omega)]^{2}}{[\eta(\omega)]^{16}}(-12)^{\frac{m}{2}}P(0)~,~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~n=2\\ 0~.~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~n>5\end{array}\right$.
+(81)
+Here we use the following relations
+$\frac{j^{\prime\prime}(\omega)}{[j(\omega)]^{\frac{1}{3}}}$ $=$ $-\frac{8\pi^{2}}{3}\frac{[E_{6}(\omega)]^{2}}{[\eta(\omega)]^{16}}~,~~~\frac{j^{\prime\prime}(\omega)}{[j(\omega)]^{\frac{2}{3}}}=-\frac{8\pi^{2}}{3[\eta(\omega)]^{8}}\frac{[E_{6}(\omega)]^{2}}{E_{4}(\omega)}\rightarrow\infty~,$~
+$\frac{j^{\prime\prime}(i)}{[j(i)-1728]^{\frac{1}{2}}}$ $=$ $-\frac{2\pi^{2}}{[\eta(\omega)]^{12}}\frac{[E_{4}(i)]^{4}}{E_{6}(i)}\rightarrow\infty~,$ (82)
+and
+$\frac{[j^{\prime}(i)]^{2}}{[j(i)-1728]^{2-\frac{m}{2}}}$ $=$ $-\frac{4\pi^{2}}{[\eta(\omega)]^{12m}}\frac{[E_{6}(i)]^{2}[E_{4}(i)]^{4}}{[E_{6}(i)]^{4-m}},$~~~
+$\frac{[j^{\prime}(\omega)]^{2}}{[j(\omega)]^{2-\frac{n}{3}}}$ $=$ $-\frac{4\pi^{2}}{[\eta(\omega)]^{8n}}\frac{[E_{6}(\omega)]^{2}[E_{4}(\omega)]^{4}}{[E_{4}(\omega)]^{6-n}}~.$ (83)
+to cancel each divergent terms among the $j^{\prime\prime}(\tau)$ and $(j^{\prime}(\tau))^{2}$ related terms within $H^{\prime\prime}(\tau)$ at the fixed points when $m=1$ or $n=1$.
+When $H(\tau)=0$, we have the following discussions on the condition $H^{\prime\prime}(\tau)=0$:
+  • The case $\tau=i$ is a solution of $H(\tau)=0$:
+    When $\tau=i$ is a solution of $H(\tau)=0$ for $m>0$, the condition $H^{\prime\prime}(\tau)=0$ can be satisfied automatically for $m\neq 2$.
+For other choices of $m$, it can also be satisfied when $\mathcal{P}(1728)=\mathcal{P}^{\prime}(1728)=0$, requiring $P(j(\tau))$ to take the form
+    $P(j(\tau))\propto(j(\tau)-1728)^{k+1}~,~~~~k\in\mathbb{N}~.$ (84)
+    After the redefinition of $m$, we conclude that $H^{\prime\prime}(i)=0$ can be satisfied for $m\neq 0,2$.
+  • The case $\tau=\omega$ is a solution of $H(\tau)=0$:
+    Similarly, when $\tau=\omega$ is a solution of $H(\tau)=0$ for $n>0$, the condition $H^{\prime\prime}(\tau)=0$ can be satisfied automatically for $n\neq 2$.
+For other choices of $m$, it can also be satisfied when $\mathcal{P}(0)=0$, requiring $P(j(\tau))$ to take the form
+    $P(j(\tau))\propto[j(\tau)]^{k}~,~~~~k\in\mathbb{N}~.$ (85)
+    After the redefinition of $n$, we conclude that $H^{\prime\prime}(\omega)=0$ can be satisfied for $n\neq 0,2$.
+  • The case $\tau=\tilde{\tau}_{i}$ with $P(j(\tilde{\tau}_{i}))=0$:
+    The condition $H^{\prime\prime}(\tau)=0$ can be satisfied for $P^{\prime}(j(\tilde{\tau}_{i}))=P^{\prime\prime}(j(\tilde{\tau}_{i}))=0$, which requires $\tau=\tilde{\tau}_{i}$ being a root of $P(j(\tau))$ with multiplicity three or greater.
+Therefore, the most general form of $P(j(\tau))$ should be
+    $P(j(\tau))=\pm g(j(\tau))\prod\limits_{i}(j(\tau)-j(\tilde{\tau}_{i}))^{k_{i}}~,~~~~k_{i}\in\mathbb{N}~.$ (86)
+    with $k_{i}\geq 3$ and $g(j(\tau))$ an arbitrary positive polynomial function of $j(\tau)$.
+It is obvious from (66) that the stationary condition for $\tau$ is determined by the solutions of $W_{M}^{*}W_{M}^{\prime}=0$ and $|W_{M}|^{2}=0$ when $H(\tau)=0$.
+From (55) and (81), it is obvious that $W_{M}^{*}W_{M}^{\prime}=0$ always vanish at $\tau=i$ except when $m=1$ (or $\tau=\omega$ except when $n=1$).
+However, they will not vanish at $\tau=\tilde{\tau}_{i}$ in general, where $P(j(\tilde{\tau}_{i}))=0$.
+Therefore, solutions of equation $V_{\tau}/|H(\tau)|^{2}=0$ (for $H(\tau)\neq 0$) and fixed points $\tau=i,\omega$ are always stationary points with respect to $\tau$ and $\tau^{*}$.
+As the solutions of combined equations $V_{\tau}/|H(\tau)|^{2}=0$ (for $H(\tau)\neq 0$) and (61) in general can not be expressed in analytical forms and need numerical calculations, it is hard to get the desired location of the modulus VEV for our model buildings.
+So, we would like to concentrate on other solutions that can be determined analytically, with the corresponding modulus VEVs deviating slightly from the fixed points.
+Such non-fixed point solutions of modulus VEVs can come from $H(\tau)=0$.
+The value $\tau=\tilde{\tau}_{i}$ with $P(j(\tilde{\tau}_{i}))=0$ generally fails to be a solution of $V_{\tau}=0$ because the derivatives $H^{\prime}(\tau)$ (or the expression $H^{\prime\prime}(\tau)-\frac{3iH^{\prime}(\tau)}{2\pi}\hat{G}_{2}(\tau,\tau^{*})]$) are typically non-zero at this point.
+In order for $H^{\prime}(\tilde{\tau}_{i})=0$, $P(j(\tau))$ needs to take the form (72).
+In such a case, non-fixed point values of $\tau$ with $P(j(\tau))=0$ can also satisfy $P^{\prime}(j(\tau))=0$ to guarantee $\tilde{W}_{M}^{\prime}\tilde{W}^{*}_{M}=0$.
+We will concentrate on such interesting choices of $P(j(\tau))$ in our subsequent studies unless otherwise specified.
+As the $H(\tau)$ (with $m,n\neq 1$) factor can be effectively factored out from $\tilde{W}_{M}^{\prime}$ and $\tilde{W}_{M}$, any non-fixed point value of $\tau$ satisfying $|H(\tau)|^{2}=0$ is an extremum of the potential, as long as the condition $V_{X_{nl}}=0$ is met.
+It was argued in target:space ; Ding:2024neh ; Higaki:2024pql that modular invariance and $\tau\leftrightarrow-\tau^{*}$ symmetry of the scalar potential ensure that the first derivative of scalar potential on $\tau$ along certain directions at the boundary of the fundamental domain vanishes
+$\left.\frac{\partial}{\partial x_{1}}V\right|_{x_{1}=\pm 1/2,0}=0~,~~~\left.\frac{\partial}{\partial\rho_{\tau}}V\right|_{\rho_{\tau}=1}=0~,$~~~ (87)
+where $\tau\equiv x_{1}+ix_{2}\equiv\rho_{\tau}e^{i\theta_{\tau}}$.
+As the two independent directional derivatives vanishes at the finite fixed points $\tau=\omega$ and $\tau=i$, the scalar potential is always stationary with respect to $\tau$ at such finite fixed points King:2023snq ; Higaki:2024pql , agreeing with our previous discussions.
+The extrema conditions (61) can be rewritten as
 $\frac{\partial}{\partial\rho}V=\frac{\partial}{\partial\theta}V=0~,$ (88)
-    for $X_{nl}\equiv\rho e^{i\theta}$. We have the following discussions for some non-generic simplified scenarios:
-    • Case $\rho=0$:
-      This corresponds to $X_{nl}=0$, reducing to the case with only a single $\tau$ modulus field, which has already been extensively discussed target:space ; NPP:2201.02020 .
+with $X_{nl}\equiv\rho e^{i\theta}$. We have the following discussions for some special parameter choices:
   • Case $f_{0}=0$:
-      The extrema of the scalar potential are located at the parameter points with $X_{nl}=0$.
-Again, the modulus VEV of the corresponding extrema is the same as the case with only a single $\tau$ modulus field.
+    The extrema of the scalar potential are located at parameter points where $X_{nl}=0$.
+The modulus VEV associated with these extrema is identical to the case involving only a single modulus field $\tau$, which had been extensively studied in previous works target:space ; NPP:2201.02020 ; Leedom:2022zdm .
+It has been demonstrated that the minima of the scalar potential can slightly deviate from the left (right) cusp symmetric point and the boundary when $m\neq 0$ and $n=0$ for $H(\tau)$, resulting in negative vacuum energy.
   • Case $c_{1}=0$:
-      The extrema of the scalar potential are also located at the parameter point with $X_{nl}=0$, and the corresponding potential energy is
+    The extrema of the scalar potential are also located at the parameter points with $\rho=0$ (that is, $X_{nl}=0$). The corresponding potential energy is given by
     $\left.V\right|_{X_{nl}=0}=\frac{1}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left|{H(\tau)}\right|^{2}f_{0}^{2}\geq 0~,$ (89)
-      for values of $\tau$ within the fundamental region.
-To obtain a vanishing (very tiny positive) vacuum energy $\Lambda^{4}\sim 10^{-120}M_{Pl}^{4}\approx 0$, we can find the parameter points that are solutions of ${H(\tau)}=0$ and also satisfy the minimum condition (with respect to $\tau$ and $\tau^{*}$) with $\tilde{W}_{M}=0$.
-It can be checked that solutions of ${H(\tau)}=0$ can also satisfy the minimum condition $\tilde{W}_{M}=0$, which will be clear shortly.
-Therefore, the stabilized modulus VEV is located at the solutions of ${H(\tau)}=0$ with (almost) vanishing vacuum energy in this case.
-    We now focus on the general case with $f_{0},c_{1},X_{nl}\neq 0$. It can be calculated that the extreme condition for $\theta$ is
-    $\frac{\partial}{\partial\theta}V=-\frac{2f_{0}c_{1}^{3}}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\rho\sin\theta\left[\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}-2\left|H(\tau)\right|^{2}\right]=0.$ (47)
-    First, we should note that, as $[(2\Im\tau)|\eta(\tau)|^{4}]^{-1}$ diverges when $\Im\tau=\infty$, the $\tau=i\infty$ point is not a solution of the extreme condition for $\theta$.
-    For generic $\theta$ (with $\theta\neq 0,\pi$ to render $X_{nl}$ not real) in the case $f_{0},c_{1},\rho\neq 0$, the extreme condition for $\theta$ reduces to the constraint
-    $\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}-2\left|H(\tau)\right|^{2}=0~{}.$ (48)
-    After substituting this constraint in the extreme condition for $\rho$, we have
-    $\frac{\partial}{\partial\rho}V=4\rho f_{0}^{2}\frac{\left|H(\tau)\right|^{2}}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}=0~{}.$ (49)
-    Such an extreme condition for $\rho$ can be satisfied with
-    $\left|H(\tau)\right|^{2}=0~{}.$ (50)
-    Thus, we have the extreme conditions for the case with $f_{0},c_{1}\neq 0$ and a non-vanishing non-real $X_{nl}$
-    $\left|\tilde{W}_{M}\right|^{2}=\left|H(\tau)\right|^{2}=0~{}.$ (51)
-    Substituting back into the scalar potential, we can see that the potential energy at the minimum vanishes.
-Therefore, the presence of a non-trivial $X_{nl}$ sector, whose contributions can uplift the scalar potential, can naturally lead to a flat Minkowski universe in this case.
-    From the expression of $H(\tau)$, the constraint $H(\tau)=0$ requires
-    $j(\tau)^{n/3}[j(\tau)-1728]^{m/2}=0,~{}~{}{\rm or}~{}~{}~{}~{}{\cal P}(j(\tau))=0.$ (52)
-    Obviously, $\tau=i,\omega$ are its solutions for $m\geq 1$ and $n\geq 1$, respectively.
-As ${\cal P}(j(\tau))$ is an arbitrary polynomial function of $j(\tau)$, from the fundamental theorem of algebra, any complex value of $j(\tau)$ can be the root of typical chosen polynomial equation ${\cal P}(j(\tau))=0$, leading to no interesting constraint at all unless we specify the concrete form of ${\cal P}(j(\tau))$.
+    for all values of $\tau$ within the fundamental region.
+The parameter points for local minimum of the scalar potential should satisfy both $H(\tau)=0$ and $X_{nl}=0$.
+As $|H(\tau)|^{2}$ can be effectively factored out from $V_{\tau}$ with $P(j(\tau))$ taking the form (86), solutions of $H(\tau)=0$ indeed satisfy $V_{\tau}=0$ in general (except for typical choices of $m,n\leq 2$ at the fixed points).
+Therefore, in this case, the stabilized modulus VEVs are located at the solutions of $H(\tau)=0$, corresponding to vanishing vacuum energy.
+  • Case $f_{0}\neq 0$, $c_{1}\neq 0$:
+    We now focus on the general case with $f_{0},c_{1}\neq 0$. It can be calculated that the extremum condition for $\theta$ is
+    $\frac{\partial}{\partial\theta}V=-\frac{2f_{0}c_{1}^{3}}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\rho\sin\theta\left[\frac{(2\Im\tau)^{2}}{3}\left|\widetilde{W}_{M}\right|^{2}-2\left|H(\tau)\right|^{2}\right]=0~,$ (90)
+    while the extremum condition for $\rho$ is
+    $\frac{\partial}{\partial\rho}V$ $=$ $\frac{1}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{2\rho\left[\frac{(2\Im\tau)^{2}}{3}\left|\widetilde{W}_{M}\right|^{2}\left(f_{0}^{2}+c_{1}^{6}\right)-2c_{1}^{6}\left|H(\tau)\right|^{2}\right]\right$.
+(91)
+    $+$ $\left.2f_{0}c_{1}^{3}\cos\theta\left[\frac{(2\Im\tau)^{2}}{3}\left|\widetilde{W}_{M}\right|^{2}-2\left|H(\tau)\right|^{2}\right]\right\}=0~.$
+    It should be noted that the point $\tau=i\infty$ does not satisfy the extremum condition for $\theta$ since $(2\Im\tau)|\eta(\tau)|^{4}]^{-1}$ diverges as $\Im\tau\to\infty$.
+    As discussed previously, the $|H(\tau)|^{2}$ factor can be extracted from $\left|\widetilde{W}_{M}\right|^{2}$ for non-vanishing $H(\tau)$.
+Besides, $\left|\widetilde{W}_{M}\right|^{2}$ vanishes when $H(\tau)=0$ for choices of $P(j(\tau))$ in the form (86) and $m,n\neq 1$.
+Therefore, $|H(\tau)|^{2}=0$, which in general satisfies the extremum condition $V_{\tau}=0$ for $\tau$, can also in general satisfy the extremum conditions $V_{\rho}=V_{\theta}=0$ for $X_{nl}$.
+So we can conclude that modulus VEVs that satisfy $|H(\tau)|^{2}=0$ for any values of $X_{nl}$ are always local extrema of the scalar potential.
+    The extremum conditions for $\theta$ and $\rho$ can also be satisfied in the following cases:
+    • From (90), for $\sin\theta\neq 0$ (that corresponds to a non-real $X_{nl}$ with $\theta\neq 0,\pi$), we can get the condition for modulus $\tau$
+      $\frac{(2\Im\tau)^{2}}{3}\left|\widetilde{W}_{M}\right|^{2}-2\left|H(\tau)\right|^{2}=0,$ (92)
+      which is independent of the $\rho$ value. After substituting this condition into (91), we can obtain the extremum condition for the $\rho$ parameter:
+      $\frac{\partial}{\partial\rho}V=4\rho f_{0}^{2}\frac{\left|H(\tau)\right|^{2}}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}=0.$ (93)
+      Such an extremum condition for $\rho$ can be satisfied with
+      $\left|H(\tau)\right|^{2}=0.$ (94)
+      Thus, we have the extremum conditions for the case with $f_{0},c_{1}\neq 0$ and a non-vanishing non-real $X_{nl}$ value:
+      $\left|\widetilde{W}_{M}\right|^{2}=\left|H(\tau)\right|^{2}=0.$ (95)
+      Substituting back into the scalar potential, we can see that the potential energy at this extremum vanishes.
       Then, we would like to focus on the constraint
-    $\tilde{W}_{M}\equiv H^{\prime}(\tau)-\frac{3iH(\tau)}{2\pi}\hat{G}_{2}(\tau,\tau^{*})=0~{},$ (53)
-    which satisfies $H(\tau)=0$. Such a constraint reduces to
-    $H^{\prime}(\tau)=\left\{j(\tau)^{n/3}[j(\tau)-1728]^{m/2}{\cal P}(j(\tau))\right\}^{\prime}=0~{},$ (54)
-    $\Rightarrow$ $j^{\prime}(\tau)\left[\frac{n}{3}j(\tau)^{n/3-1}[j(\tau)-1728]^{m/2}{\cal P}(j(\tau))+\frac{m}{2}j(\tau)^{n/3}[j(\tau)-1728]^{m/2-1}{\cal P}(j(\tau))\right.$
-    $\left.+j(\tau)^{n/3}[j(\tau)-1728]^{m/2}\frac{\partial{\cal P}(j(\tau))}{\partial j(\tau)}\right]=0~{},$
-    $\Rightarrow$ $H(\tau)j^{\prime}(\tau)\left[\frac{n}{3}\frac{1}{j(\tau)}+\frac{m}{2}\frac{1}{j(\tau)-1728}+\frac{1}{{\cal P}(j(\tau))}\frac{\partial{\cal P}(j(\tau))}{\partial j(\tau)}\right]=0~{}.$
-    So $H^{\prime}(\tau)=0$ can trivially hold when $H(\tau)=0$.
-    On the other hand, $H^{\prime}(\tau)=0$ can also be satisfied by the other two conditions
-    $\left[\frac{n}{3}\frac{1}{j(\tau)}+\frac{m}{2}\frac{1}{j(\tau)-1728}+\frac{1}{{\cal P}(j(\tau))}\frac{\partial{\cal P}(j(\tau))}{\partial j(\tau)}\right]=0~{},$ (55)
-    or $j^{\prime}(\tau)=0$.
-    We have the following discussions for the condition (55)
-    • $m+n\neq 0$:
-      When $n\neq 0$ and $m=0$, condition (55) reduces to
-      $\frac{d{\cal P}(j(\tau))}{{\cal P}(j(\tau))}=-\frac{n}{3}\frac{dj(\tau)}{j(\tau)}~{},$ (56)
-      which gives
-      ${\cal P}(j(\tau))=C_{0}\left[j(\tau)\right]^{-n/3}~{}.$ (57)
-      Unless $n\in 3\mathbb{Z}$ is negative, ${\cal P}(j(\tau))$ is not a polynomial of $j(\tau)$.
-Similarly, when $m\neq 0$ and $n=0$, we can deduce that $m\in 2\mathbb{Z}$ should be negative.
-For $m\neq 0$ and $n\neq 0$, similar deduction can give
-      ${\cal P}(j(\tau))=C_{0}^{\prime}\left[j(\tau)\right]^{-n/3}\left[{j(\tau)-1728}\right]^{-m/2}~{},$ (58)
-      which also requires both $m\in 2\mathbb{Z}$ and $n\in 3\mathbb{Z}$ being negative to guarantee ${\cal P}(j(\tau))$ being a polynomial of $j(\tau)$.
-As $m,n$ are defined to be positive integers, this condition requires $m=n=0$.
-    • $m=0$ and $n=0$:
-      With the constraints $m=n=0$, the condition (55) reduces to
-      $\frac{\partial{\cal P}(j(\tau))}{\partial j(\tau)}=0~{},$ (59)
-      which gives
-      ${\cal P}(j(\tau))={\rm Constant}.$ (60)
-      As $H(\tau)=0$, this constant should be zero.
-    Therefore, the condition (55) requires a vanishing $H(\tau)$. For the condition $j^{\prime}(\tau)=0$, from the derivative of the Klein $j$-function
-    $\frac{\partial}{\partial\tau}j(\tau)=-2\pi i\frac{{E_{6}(\tau)}{E_{4}^{2}(\tau)}}{[\eta(\tau)]^{24}}=0~{},$ (61)
-    we can see that its solutions are $\omega$ and $i$, leading to no new additional constraints. Therefore, the extreme conditions (51) can reduce to the requirement of vanishing $H(\tau)$.
-    Before we conclude this scenario, it should be noted that $\theta=0,\pi$, which corresponds to real $X_{nl}=\pm\rho$ with positive $\rho$, are also solutions of the the extreme condition (47) for $\theta$.
+      $\widetilde{W}_{M}\equiv H^{\prime}(\tau)-\frac{3iH(\tau)}{2\pi}\hat{G}_{2}(\tau,\tau^{*})=0,$ (96)
+      which satisfies $H(\tau)=0$.
+Discussions on the solutions of $H^{\prime}(\tau)=0$ are given in Appendix C.
+For the fixed points $\tau=\omega,i$, they are always the local extrema of the scalar potential when $m,n>1$.
+When $m<1$ or $n<1$, whether the fixed points $\tau=\omega,i$ are local extrema can easily be figured out case by case, with the relevant results agreeing with the single modulus case in Leedom:2022zdm .
+The extrema conditions (95) and (66) of the scalar potential can also be satisfied for non-fixed points solutions of $H(\tau)=0$ when $P(j(\tau))$ takes the form (72), and the corresponding vacua are always of Minkowski type.
+The Minkowski vacua determined by $H(\tau)=0$ are independent of the specific value of $X_{nl}$.
+To identify the global minimum of the scalar potential, it is necessary to evaluate the potential energies for all extrema, including the case where $X_{nl}$ is real.
+    • Real values of $X_{nl}$ with $\theta=0$ or $\theta=\pi$.
       In this case, the scalar potential reduces to
       $V(X_{nl},X^{*}_{nl},\tau,\tau^{*})$ (97)
-    $=$ $\frac{1}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}\left(f_{0}^{2}\rho^{2}+c_{1}^{6}\pm 2f_{0}\rho c_{1}^{3}\right)\right.$
-    $+$ $\left|H(\tau)\right|^{2}\left(f^{2}_{0}+2f_{0}^{2}\rho^{2}\pm 2f_{0}c_{1}^{3}\rho+\rho^{2}c_{1}^{6}\frac{}{}\right)~{}$
-    -- $\left.3\left|{H(\tau)}\right|^{2}\left(f_{0}^{2}\rho^{2}+c_{1}^{6}\pm 2f_{0}\rho c_{1}^{3}\right)\right\}~{}$
-    $+$ $\frac{\rho^{2}}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}c_{1}^{6}+\left|H(\tau)\right|^{2}f^{2}_{0}-3c_{1}^{6}\left|{H(\tau)}\right|^{2}\right\}~{}.$
-    The scalar potential, which is a quadratic polynomial of $\rho$, can be written as
-    $V=\frac{1}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left[C_{1}\rho^{2}\pm C_{2}\rho+C_{3}\right]~{},$ (63)
+      $=$ $\frac{1}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}\left(f_{0}^{2}\tilde{X}^{2}+c_{1}^{6}+2f_{0}\tilde{X}c_{1}^{3}\right)\right$.
+      $+$ $\left|H(\tau)\right|^{2}\left(f^{2}_{0}+2f_{0}^{2}\tilde{X}^{2}+2f_{0}c_{1}^{3}\tilde{X}+\tilde{X}^{2}c_{1}^{6}\right)$
+      - $\left.3\left|H(\tau)\right|^{2}\left(f_{0}^{2}\tilde{X}^{2}+c_{1}^{6}+2f_{0}\tilde{X}c_{1}^{3}\right)\right\}$
+      $+$ $\frac{\tilde{X}^{2}}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}c_{1}^{6}+\left|H(\tau)\right|^{2}f^{2}_{0}-3c_{1}^{6}\left|H(\tau)\right|^{2}\right\},$
+      for real values of $X_{nl}$, which we denote simply as $\tilde{X}$. It is a quadratic polynomial of $\tilde{X}$, which can be rewritten as
+      $V(\tilde{X},\tau,\tau^{*})=\frac{1}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left[C_{1}\tilde{X}^{2}+C_{2}\tilde{X}+C_{3}\right]~,$ (98)
       where
@@
       $C_{3}$ $=$ $\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}c_{1}^{6}+\left|H(\tau)\right|^{2}\left(f_{0}^{2}-3c_{1}^{6}\right)~.$ (99)
-    The scalar potential has a minimum with the corresponding $\rho$ parameter
-    $\rho_{0}=\mp\frac{C_{2}}{2C_{1}}~{},$ (65)
-    and the potential energy
-    $\left.V\right|_{\rho=\rho_{0}}=C_{3}-\frac{C_{2}^{2}}{2C_{1}}~{}.$ (66)
-    The modulus VEV is determined by the minimum conditions for the potential (66) with respect to $\tau,\tau^{*}$, and the condition $C_{2}^{2}=2C_{1}C_{3}$ should be established to obtain a vanishing (almost vanishing tiny positive) vacuum energy.
-It can be checked from previous discussions that solutions of ${H(\tau)}=0$ can also satisfy such minimum conditions.
-    Thus, we can conclude that, in the scenario with $k_{X}=0$ for $f_{0}\neq 0$, the extrema of the scalar potential are located at the points determined by $H(\tau)=0$, and the corresponding vacuum is of Minkowski type.
-As noted previously, the constraint $H(\tau)=0$ predicts that the fixed points $z_{0}=\omega,i$ are always the local extrema unless $m=n=0$.
-Other solutions of $\tau$ from ${\cal P}(j(\tau))=0$ are essentially possible because of the arbitrariness of the polynomial form of ${\cal P}(j(\tau))$.
-Therefore, other minima, which are determined by (44) to lie at the boundary of the fundamental domain or at the imaginary axis, can also be possible.
-    In fig.1, we show the energies of the scalar potential at modulus values taken within the fundamental domain with various parameter choices.
-It can be seen that the potential energies are non-negative when $f_{0}\neq 0$.
-    Figure 1: The values of $V(X_{nl},\bar{X}_{nl},\tau,{\tau}^{*})$ within the fundamental domain in the scenario with $k_{X}=0$ for various parameter choices when $f_{0}\neq 0$.
-We adopt $\mathcal{P}(j)=1$ for the left panel and $\mathcal{P}(j)=j-4.5764069879884376\times 10^{-11}$ for the right panel.
-  • Scenario with small $X_{nl}$ (when $|X_{nl}|\rightarrow 0$):
-    In this case, at small $|X_{nl}|$ region, the $|X_{nl}|^{2}$ terms can be neglected in addition to the constraint $X_{nl}^{2}=0$.
-Note that the $|X_{nl}|^{2}$ terms will dominate the scalar potential at very large $|X_{nl}|$ values, ensuring that the potential can be bounded from below.
-    The scalar potential can be simplified to
-    $V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})$ (67)
-    $=$ $\frac{1}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}\left(c_{1}^{6}+2\Re\left[f_{0}X_{nl}c_{1}^{3}\left[\eta(\tau)\right]^{2k_{X}}\right]\right)\right.$
-    $+$ $\left.\frac{(2\Im\tau)^{2}}{3}\left(2\Re\left[k_{X}f_{0}X_{nl}\left[\eta(\tau)\right]^{2k_{X}}\tilde{W}_{E}\right]\right)H(\tau)c_{1}^{3}\tilde{W}_{M}^{*}\right]$
-    $+$ $(2\Im\tau)^{k_{X}}\left|H(\tau)\right|^{2}\left(f_{0}^{2}\left|\eta(\tau)\right|^{4k_{X}}+2\Re\left[f_{0}\eta(\tau)^{2k_{X}}\frac{{X}_{nl}c_{1}^{3}}{(2\Im\tau)^{k_{X}}}\right]\right)$
-    $+$ $2\Re\left[\frac{i(2\Im\tau)k_{X}X_{nl}}{3}\left(f_{0}[\eta(\tau)]^{2k_{X}}H(\tau)\right)^{*}\tilde{W}_{M}c_{1}^{3}\right]$
-    -- $\left.3\left(c_{1}^{6}+2\Re\left[f_{0}X_{nl}c_{1}^{3}\right]\right)\left|{H(\tau)}\right|^{2}\frac{}{}\right\}~{},$
-    with
-    $\tilde{W}_{M}$$\equiv$ $H^{\prime}(\tau)-\frac{3iH(\tau)}{2\pi}\hat{G}_{2}(\tau,\tau^{*})~{},$
-    $\tilde{W}_{E}$$\equiv$ $\frac{i}{2\pi}\hat{G}_{2}(\tau,\tau^{*})+\frac{i}{2\Im\tau}~{}.$ (68)
-    It can be seen easily from the linear terms of $X_{nl}$ that the extrema conditions for $X_{nl}$ will not lead to trivial constraints for $\tau$ with arbitrary choices of $k_{X}$.
-Therefore, other modulus values lying on the boundary of fundamental domain or imaginary axis other than the fixed point, can possibly be the minimum.
-The initially stabilized modulus VEV without contributions from $X_{nl}$ sector can then be destabilized by new contributions from non-vanishing $X_{nl}$.
-We will discuss the extrema conditions for the most general case in our subsequent studies.
-It should be noted that the modulus stabilization mechanism with nilpotent $\textbf{X}_{nl}$ superfield can be fairly generic and ubiquitous, which is independent on the following discussions on the regeneration of small $\mu_{eff}$ term following modulus stabilization.
+      Obviously, the scalar potential is real for all (real) values of $\widetilde{X}$.
+The scalar potential is bounded from below for $\widetilde{X}\in\mathbb{R}$ when the coefficients of the quadratic polynomial in (98) satisfy $C_{1}>0$ (and $C_{2}=0$ when $C_{1}=0$, rendering the scalar potential finite).
+      For $C_{1}>0$, the scalar potential has a minimum with the corresponding $\widetilde{X}$ value and the potential energy
+      $\left.V\right|_{min;\tilde{X}=\frac{-C_{2}}{2C_{1}}}$ $=$ $C_{3}-\frac{C_{2}^{2}}{4C_{1}}~.$~ (100)
+      To determine the extrema of $\left.V\right|_{min}$, we need to determine
+      $\frac{\partial}{\partial\tau}\left[C_{3}-\frac{C_{2}^{2}}{4C_{1}}\right]=\frac{\partial}{\partial\tau^{*}}\left[C_{3}-\frac{C_{2}^{2}}{4C_{1}}\right]=0~.$ (101)
+      These extremum conditions, after factoring out a common non-vanishing $|H(\tau)|^{2}$ term, yield non-trivial equations for $\tau$ that generally require numerical solution.
+      For $C_{1}=0$, it can be checked that $C_{2}=0$ when stationary condition $H(\tau)=0$ for modulus VEV $\tau$ holds, which corresponds to Minkowski vacua for all values of $\tilde{X}$.
+Therefore, in line with earlier discussions, VEVs of modulus field that satisfy $H(\tau)=0$ for all values of ${X}_{nl}$ (whether real or complex) represent local minima of the scalar potential.
+      The pseudo-modulus $X_{nl}$, i.e.
+a continuous space of vacua with degenerate tree-level vacuum energies, can be lifted by SUSY breaking effects.
+Taking into account soft SUSY breaking mass term $m_{X}^{2}|X_{nl}|^{2}$ from additional SUSY breaking sources with proper mediation mechanism (or from the one-loop Coleman-Weinberg potential of $X_{nl}$ after including its couplings to matter fields, for example, of the O’Raifeartaigh type $\lambda_{ij}X_{nl}+M_{ij}]\phi_{i}\phi_{j}$ in the superpotential), the scalar potential for $X_{nl}$ taking the form $V(X_{nl})\approx|\tilde{C}|^{2}f_{0}^{2}+m_{X}^{2}|X_{nl}|^{2}$ can stabilize the pseudo-modulus at $X_{nl}=0$ in the flat SUSY limit, with $\tilde{C}$ some constant determined by the stabilized value of $\tau$.
+[…]
+    For the case $C_{1}>0$, the minimum (100) is always located at some value of $\widetilde{X}$ with $\widetilde{X}\ll M_{Pl}$.
+This minimum can either be the global minimum or merely a local minimum when the bounded-from-below condition is relaxed by allowing $C_{1}\leq 0$.
+    We select some benchmark parameter points with various choices of $f_{0},c_{1}$ and calculate numerically the potential energies for all extrema to identify the global minimum.
+For many benchmark points, the global minimum is determined by the condition $H(\tau)=0$ with vanishing vacuum energy, which are independent of the value of $X_{nl}$  [^2] As noted previously, the pseudo-modulus $X_{nl}$ will be stabilized at $X_{nl}=0$ after including the SUSY breaking contributions.
+So we choose the stabilized value $X_{nl}=0$ for the extrema determined by the condition $H(\tau)=0$..
+In Fig.1, we show the energies of the scalar potential for modulus values within the fundamental domain and $X_{nl}=0$, considering various parameter choices.
+    Figure 1: The values of $V(X_{nl},\bar{X}_{nl},\tau,{\tau}^{*})$ within the fundamental domain in the scenario with $k_{X}=0$ and $X_{nl}=0$ for various parameter choices when $f_{0}\neq 0$ and $f_{0}\gg c_{1}^{2}$.
+We adopt $\mathcal{P}(j(\tau))=1$ for the left panel and $\mathcal{P}(j(\tau))=(j(\tau)-j(\tau_{0}))^{2}$ with $\tau_{0}=e^{i\frac{2\pi}{3}(1-0.006)}$ for the right panel.
+    In the limit $f_{0}\to 0$, this scenario reduces to the one involving only a single modulus field $\tau$, whose vacuum is generally not of Minkowski type and the corresponding vacuum energy can possibly be negative.
+Therefore, for tiny values of $f_{0}$, the vacuum energy of the global minimum can become negative.
+Numerical results on the global minima of some benchmark points with non-vanishing vacuum energy are shown in Table 1.
+    $m$ $n$ $f_{0}/M_{\mathrm{Pl}}^{2}$ $c_{1}/M_{\mathrm{Pl}}$ $X_{nl}/M_{\mathrm{Pl}}$ $\tau$ $P(j(\tau))$ $V_{min}/M_{\mathrm{Pl}}^{4}$
+    0 0 $0.1$ $0.5$ 0.6318-0.6318 $0.0021+1.002i$ 11 0.0525-0.0525
+    11 11 $^{-4} 0.10.1$ $0.9994$ $0.19337+0.98275i$ 11 1.25-1.25
+    22 0 $10^{-7}$ 0.01-0.01 0.9994-0.9994 $0.4892+0.8733i$ 11 $4.466\times 10^{-5}$
+    11 22 $1.128\times 10^{-6}$ 0.01-0.01 0.3322-0.3322 0.1910+0.9830i-0.1910+0.9830i 11 $7.282\times 10^{-5}$
+    22 11 $2.974\times 10^{-7}$ 0.01-0.01 $0.2920$ $0.2865+0.9612i$ 11 $7.434\times 10^{-4}$
+    11 33 $2.337\times 10^{-6}$ 0.01-0.01 0.4126-0.4126 $0.1588+0.9911i$ 11 $7.103\times 10^{-4}$
+    Table 1: The global minima and the corresponding vacuum energies for some benchmark parameter points in the case $k_{X}=0$, $f_{0}\neq 0$ and $c_{1}\neq 0$.
+    From a physical perspective, we are particularly interested in the local minimum with an intermediate/small value of $\widetilde{X}$, rather than a global minimum situated at $\widetilde{X}\sim M_{\text{Pl}}$ or $\widetilde{X}\sim-M_{\text{Pl}}$.
+For typical benchmark parameter choices, an estimation of the lifetime for the local minimum with an intermediate/small value of $\widetilde{X}$ suggests that such a local minimum can be metastable, with a lifetime exceeding the age of the universe.
+Therefore, we will concentrate on the global (or long-lived meta-stable) minimum with $|X_{nl}|\ll M_{Pl}$ in our subsequent discussions.
+    The lifetime of a vacuum can be evaluated semi-classically through a rate of bubble nucleation in unit volume and unit time Coleman ; Coleman1 .
+The decay probability, which sets the decay rate, is given by $\Gamma/V=A\exp(-B)$, where $B$ is the action of a Euclidean $O(4)$-symmetric bounce solution, which smoothly interpolates between a true-vacuum configuration at its center and the false vacuum whose decay is being described.
+The prefactor $A$ depends on the functional determinants and represents the loop correction of the fluctuations of the fields.
+The prefactor $A$ can be approximately given by $A\sim v^{4}$, where $v$ is some characteristic mass scale of the theory under study.
+We adopt a conservative choice $v\sim{\cal O}(M_{Pl})$ for our estimation of lifetime.
+    The generalization to cases involving multiple fields entails determining the bounce configuration by solving a set of coupled ordinary differential equations, subject to the appropriate boundary conditions.
+For general multi-dimensional field spaces, there will be many bounce solutions with different actions, and the tunneling rate will be dominated by the solution with the lowest action.
+The field profile, which solves the Euclidean equations of motion, traverses a one-dimensional path in the field space.
+If one chooses a particular small subset of path in the field space and treats it as a one-field potential, solving the field equations for the one-dimensional potential along this path can give an estimation on the action, which can be close to the actual value of the action 1702-00356 .
+As the tunneling is expected intuitively to happen through the lowest possible barrier and shortest possible path in the field space, the one with the shortest possible path length which is a straight line between the two minima (or the path with the lowest possible barrier in which a broken line goes through the lowest ridge in the barrier), can act as the one-dimensional approximations of potential along such a path parameterized by the path length in the field space.
+    We estimate the decay lifetime with the bounce configuration via the two paths connecting the true and false vacuum for some benchmark parameter choices that can lead to false vacuums at intermediate/small value of $\widetilde{X}$.
+We estimate the one-dimensional action for each case with the approximation of a triangular potential, whose the bounce action is given approximately Duncan by
+    $B\approx\frac{32\pi^{2}}{3}\frac{(\Delta V_{+})^{2}(\Delta\phi)^{4}}{\epsilon^{3}}~.$ (102)
+    in the situation where the difference of potential energy between the true and false vacuum $\epsilon$ is very small.
+Here $\Delta V_{\pm}=V_{Top}-V_{\pm}$ denotes the height for false (and true) vacuum with respect to the top of the barrier and $\Delta\phi=(\phi_{-}-\phi_{+})$ denotes the width of the barrier.
+    For a characteristic benchmark point, such as $f_{0}=10^{-7}M^{2}_{Pl}$ and $c_{1}=-0.01M_{Pl}$ with $m=1,n=1$ and $P(j(\tau))=(j(\tau)-j(\tau_{0}))^{2}$ for $\tau_{0}=1.001\times e^{i\frac{2\pi}{3}(1-\frac{1}{36})}$, we find that the false vacuum lies at $(X_{nl},\tau)=(0,1.001\times e^{i\frac{2\pi}{3}\frac{35}{36}})$ with vanishing vacuum energy while the true minimum vacuum lies at $(X_{nl},\tau)=(-0.99999M_{Pl},-0.42825+0.91479i)$ with $V_{min}=-2.49896M_{Pl}^{4}$.
+The peak of the lowest barrier sits at the parameter point $(X_{nl},\tau)=(-0.46285M_{Pl},-0.40398+0.92017i)$ with $V_{L-peak}=577.87447M_{Pl}^{4}$.
+We find that the bounce action of the broken line path through the lowest possible barrier leads to smaller action.
+It can be calculated with (102) that the corresponding bounce action $B\approx 2.28\times 10^{5}$.
+Therefore, the lifetime of the meta-stable vacuum $\tau=(\Gamma/V)^{-1/4}$ is much larger than the universe.
+In fact, using the space-time volume of our past lightcone Past:LC , $(VT)_{LC}=0.15/H_{0}^{4}=3.4\times 10^{166}{\rm GeV}^{-4}$ with the Hubble constant $H_{0}=1.44\times 10^{-42}{\rm GeV}$, the probability that we should have seen a bubble by now is $(VT)_{LC}](\Gamma/V)\sim 10^{-9670}$, which is exponentially small.
+4.2 The case with $k_{X}=12$
+When $k_{X}=12$, interference terms in the scalar potential obscure the analysis of the extrema conditions.
+Therefore, we leave the most general discussions on this case in our subsequent studies.
+An important observation is that an $|H(\tau)|^{2}$ factor (for $H(\tau)\neq 0$) can be extracted from the scalar potential.
+According to the discussions in the case $k_{X}=0$, for any value of $X_{nl}$, in addition to the fixed points $\tau=i,\omega$ for most choices of $m,n$, non-fixed point values of modulus VEVs satisfying $|H(\tau)|^{2}=0$ are local extrema of the scalar potential with vanishing vacuum energy when $P(j(\tau))$ takes the form (72).
+Other local minima of the scalar potential can be determined numerically from the extrema conditions with respect to $\tau$, $X_{nl}$ and the corresponding Hessian matrix.
+After specifying numerically all other extrema of the scalar potential, we can determined its global minimum or meta-stable minimum by evaluating the corresponding vacuum energies of various local minima (and the lifetime of the meta-stable minimum).
+We show in Fig.2 the vacuum energy of the scalar potential for typical benchmark parameter choices with modulus values taken in the fundamental domain in the scenario with $k_{X}=12$.
+Figure 2: The values of $V(X_{nl},\bar{X}_{nl},\tau,{\tau}^{*})$ within the fundamental domain in the scenario with $k_{X}=12$ and $X_{nl}=0$ for various parameter choices when $f_{0}\neq 0$ and $f_{0}\gg c_{1}^{2}$.
+We adopt $\mathcal{P}(j)=1$ for the left panel and $\mathcal{P}(j(\tau))=(j(\tau)-j(\tau_{0}))^{2}$ with $\tau_{0}=1.05\times e^{i\frac{\pi}{2}(1+\frac{1}{6})}$ for the right panel.
+In our discussion, we assume the existence of modular symmetry in the SUSY breaking sector.
+The modular symmetry can be broken by SUSY breaking effects, depending on the transformation property of SUSY breaking source under the modular symmetry, as discussed in Kikuchi:2022pkd .
+For example, in the moduli mediated SUSY breaking scenario Kikuchi:2022pkd , soft SUSY breaking scalar masses and trilinear terms are invariant under the modular symmetry when the F-term is regarded as a spurion with the modular weight 2-2.
+In our case, non-vanishing F-term VEV of $X_{nl}$ when $k_{X}\neq 0$ can possibly act as the SUSY breaking source and also break the modular symmetry, whose transformation should be compatible with the modular invariance of $K_{X_{nl}^{\dagger}X_{nl}}F_{X_{nl}}$, possibly leading to a soft SUSY breaking spectrum that is non-invariant under modular symmetry.
+The form of the low energy soft SUSY breaking spectrum depends on the concrete mediation mechanism of SUSY breaking that mediates the SUSY breaking effect from the hidden sector to the visible sector.
+It should be noted that the modulus stabilization mechanism with nilpotent $\textbf{X}_{nl}$ superfield can be fairly generic and ubiquitous, which is independent of the following discussions on the regeneration of small $\mu_{eff}$ term following modulus stabilization.
 5 Regeneration of naturally small $\mu_{eff}$ following modulus stabilization
@@
 It is known that the fixed points $\tau_{FP}=i,\omega,i\infty$ for modular theories based on the $\Gamma_{N}$ invariance can preserve $Z_{2}^{S},Z_{3}^{ST},Z_{N}^{T}$ residual symmetries, respectively Novichkov:2021evw .
-We can infer from the modular invariance of scalar potential that the fixed points should be the local extrema of the scalar potential, as they act non-trivially (with residual symmetries) on the derivatives of the potential and must vanish.
+As noted previously, we can infer from the modular invariance of scalar potential that the finite fixed points should in general be the local extrema of the scalar potential, as they act non-trivially (with residual symmetries) on the derivatives of the potential and must vanish.
 Although they may not be the global minima, experiences from modulus stabilization in string theory indicate that the minima should always lie at/near $\tau_{FP}=i,\omega$ for the tree-level scalar potential, as the scalar potential generally diverges for $\tau\rightarrow i\infty$.
-It was found in Novichkov:2021evw ; Feruglio:2021dte that the hierarchies among the charged leptons and quarks mass can follow solely from the properties of the modular forms (strictly speaking, the representation of the mass term bilinear with respect to the residual group) in the vicinity of symmetric points, avoiding the fine-tuning of the constant parameters (or without the need to introducing extra weighton-like scalar fields).
+It was found in Novichkov:2021evw ; Feruglio:2021dte that the hierarchies among the charged leptons and quarks mass can follow solely from the properties of the modular forms (strictly speaking, the representation of the mass term bilinear with respect to the residual group) in the vicinity of symmetric points, avoiding the fine-tuning of the constant parameters (or without the need to introduce extra weighton-like scalar fields).
 Similar ideas, augmented with additional suppression powers of $q^{1/24}$ (or $(2\Im\tau)^{-1}$), can be used to generate the hierarchy between $\Lambda_{SUSY}$ and $\mu_{eff}$.
-Studies on target-space modular symmetry in string theory indicates target:space that all extrema lie on the boundary of the fundamental region or on the imaginary axis.
-We anticipate that, due to SUSY breaking contributions (for example, from $X_{nl}$ sector), the extrema can change with respect to the change of $f_{0}$ so as that the VEV $\tau_{0}$ can be altered slightly away from the fixed points $\tau_{FP}=i,\omega$.
-Such deviations from fixed points by SUSY breaking effects can lead to large suppression factors that regenerate $\mu_{eff}$ of order the electroweak scale.
 Besides, although the scalar potential for modulus field $\tau$ in general diverges when $\tau\rightarrow i\infty$, the presence of $X_{nl}$ contributions can alter the scalar potential, making it possible to stabilize $\tau$ at an intermediate value.
-Consequently, for an appropriate choice of $k_{3}$, the effective $\mu_{eff}$ parameter can naturally be much smaller than the SUSY scale $\Lambda_{SUSY}$, due to a large suppression factor originating from both powers of $q$ [or $(2\Im\tau)^{-1}$] and typical suppression behavior of modular forms away from the fixed point $i\infty$ (at $\tau_{0}$ with medium $\Im\tau_{0}$).
-To regenerate the effective $\mu$-term with a stabilized modulus field, we need to construct a modular weight $-3$ bilinear for $H_{u}$ and $H_{d}$ using proper modular forms.
-For a modular invariant bilinear of the form $\psi_{i}^{c}M_{ij}(\tau)\psi_{j}$, where $M(\tau)_{ij}$ is a modular form of level $N$ and weight $K$, the zero entries of $M(\tau)_{ij}$ at a symmetric point in the mass matrices will generically become non-zero when $\tau$ deviates slightly from the symmetric point.
+Consequently, for an appropriate choice of $k_{3}$, the effective $\mu_{eff}$ parameter can naturally be much smaller than the SUSY scale $\Lambda_{SUSY}$, due to a large suppression factor originating from both powers of $q^{1/24}$ [or $(2\Im\tau)^{-1}$] and typical suppression behavior of modular forms away from the fixed point $i\infty$ (at $\tau_{0}$ with medium $\Im\tau_{0}$).
+To regenerate the effective $\mu$-term with a stabilized modulus field, we need to construct a modular weight 3-3 bilinear for $H_{u}$ and $H_{d}$ using proper modular forms.
+For a modular invariant bilinear of the form $\psi_{i}^{c}M_{ij}(\tau)\psi_{j}$, where $M_{ij}(\tau)$ is a modular form of level $N$ and weight $K$, the zero entries of $M_{ij}(\tau)$ at a symmetric point in the mass matrices will generically become non-zero when $\tau$ deviates slightly from the symmetric point.
 The magnitudes of these residual symmetry breaking entries are controlled by the size of the departure $\epsilon$ from $\tau_{sym}$ and the field transformation properties under the residual symmetry group, which may depend on the modular weights Petcov:2023vws ; Novichkov:2021evw ; Petcov:2022fjf .
@@
 So, the desired suppression factor for the regenerated $\mu_{eff}$ can be directly obtained with proper choice of singlet representations and modular weights.
+We should note that $\mu_{eff}\ll\Lambda_{SUSY}$ can also be naturally realized through a weighton-like mechanism King:2020qaj for $H_{u}H_{d}$ bilinear.
+Specifically, we can introduce a single weighton-like superfield, which is a singlet under both the Standard Model and modular flavor symmetry, with a modular weight $k_{\phi}=1$.
+By appropriately choosing the modular weights and representations under the finite modular group $\Gamma_{N}$ for $H_{u}$ and $H_{d}$, we can ensure that only the combination
+$W\supseteq\frac{Y_{\mathbf{r}}^{(k_{\mu})}}{{\left[\eta(\tau)\right]^{2k_{3}}}}\left(\frac{\phi}{M_{F}}\right)^{k}H_{u}H_{d}~,$ (104)
+is allowed by modular invariance, where $k=k_{\mu}-k_{3}-k_{H_{u}}-k_{H_{d}}+3$ is a positive integer, $2k_{3}\in 24\mathbb{Z}+6$ and $M_{F}$ is a typical cutoff scale.
+When the weighton field $\phi$ acquires a VEV through the relevant superpotential, the electroweak-scale $\mu_{\text{eff}}$ can be naturally regenerated with a large suppression factor of order $\tilde{\phi}^{k}$, where $\tilde{\phi}\equiv\langle\phi\rangle/M_{F}\ll 1$.
+For example, the form of superpotential for weigthon superfield can be chosen to be King:2020qaj
+$W\supseteq X(Y_{\bf 1}^{4}\frac{\phi^{4}}{M_{F}^{2}}-M^{2})~,$ (105)
+with $M$ a dimensional mass parameter and $X$ an $A_{4}$ singlet.
+The F-flat condition $F_{X}^{*}=-W_{X}=0$ drives the weighton VEV $\langle\phi\rangle\sim(MM_{F})^{1/2}$, which leads to the suppression factor $\frac{\phi}{M_{F}}\sim(\frac{M}{M_{F}})^{1/2}\ll 1$ for $M\ll M_{F}$.
+It should be noted that the $H_{u}$ and $H_{d}$ superfields, which transform non-trivially under modular symmetry, can also break the modular invariance after electroweak symmetry breaking.
+As the VEVs of $H_{u}$ and $H_{d}$ are much smaller than the typical scale of modulus stabilization, the extreme conditions for $\tau$ and $X_{nl}$ will receive negligible corrections from $H_{u}$ and $H_{d}$ related terms in the limits $v_{u},v_{d}\rightarrow 0$.
+Electroweak symmetry breaking, triggered by the VEVs of $H_{u}$ and $H_{d}$, are also affected by the modular stabilization mechanism.
+However, additional contributions to the electroweak symmetry breaking sector from supergravity vanish in the $M_{Pl}\rightarrow\infty$ limit, leading to negligible contributions to the low energy MSSM scalar potential for $H_{u}$ and $H_{d}$, consequently negligible shift of the VEVs for $H_{u}$ and $H_{d}$.
 5.1 Natural $\mu_{eff}$ with modulus stabilization near the fixed point $\omega\equiv e^{i2\pi/3}$
@@
 $Y_{\bf 1^{\prime}}^{(8)}$ $=$ $(Y_{1}^{2}+2Y_{2}Y_{3})(Y_{3}^{2}+2Y_{1}Y_{2})~,~~~Y_{\bf 1}^{(8)}=(Y_{1}^{2}+2Y_{2}Y_{3})^{2}~.$ (107)
-From the expansion Petcov:2022fjf of modular functions at the point $z_{0}=\omega+\epsilon$ (with $\epsilon$ depending on the choices of $f_{0}$ such that $|\epsilon|\ll 0$)
+From the expansion Petcov:2022fjf of modular functions at the point $z_{0}=\omega+\epsilon$ (with $\epsilon$ depending on the choices of $f_{0}$ such that $|\epsilon|\ll 1$)
 $\frac{Y_{2}(z_{0})}{Y_{1}(z_{0})}$ $\simeq$ $\omega\left(1+2.235i\epsilon-2.95717\epsilon^{2}+\cdots\right)~,$
@@
 […]
-$W\supseteq\alpha\Lambda_{SUSY}\frac{Y_{\bf r}^{(k_{r})}(\tau)}{\eta(\tau)^{2k_{3}}}H_{u}H_{d}~{},$ (74)
-where $k_{H_{u}}+k_{H_{d}}=k_{r}-k_{3}+3$ and ${\bf r}(H_{u})\otimes{\bf r}(H_{d})\otimes{\bf r}\supset{\bf 1}$.
+$W\supseteq\alpha\Lambda_{SUSY}\frac{Y_{\bf r}^{(k_{r})}(\tau)}{\left[\eta(\tau)\right]^{2k_{3}}}H_{u}H_{d}~,$ (110)
+where $k_{H_{u}}+k_{H_{d}}=k_{r}-k_{3}+3$, $2k_{3}\in 24\mathbb{Z}+6$ and ${\bf r}(H_{u})\otimes{\bf r}(H_{d})\otimes{\bf r}\supset{\bf 1}$.
 When the modulus field is stabilized close to the point $\omega$, the effective $\mu$-term is given by
-$\mu_{eff}$ $=$ $\alpha\Lambda_{SUSY}\eta^{-2k_{3}}(\omega+\epsilon)Y_{\bf r}^{(k_{r})}(\omega+\epsilon)(2\Im\tau)^{(k_{H_{u}}+k_{H_{d}})/2}~{},$ (75)
-$=$ $\alpha\Lambda_{SUSY}\eta^{-2k_{3}}(\omega)Y_{\bf r}^{(k_{r})}(\omega+\epsilon)(\sqrt{3})^{(k_{r}-k_{3}+3)/2}$
+$\mu_{eff}$ $=$ $\alpha\Lambda_{SUSY}\left[\eta(\omega+\epsilon)\right]^{-2k_{3}}Y_{\bf r}^{(k_{r})}(\omega+\epsilon)(2\Im\tau)^{(k_{H_{u}}+k_{H_{d}})/2}~,$ (111)
+$=$ $\alpha\Lambda_{SUSY}\left[\eta(\omega)\right]^{-2k_{3}}Y_{\bf r}^{(k_{r})}(\omega+\epsilon)(\sqrt{3})^{(k_{r}-k_{3}+3)/2}$
 $\approx$ $\alpha\Lambda_{SUSY}\left[0.711506\cdot e^{-0.523599i}\right]^{-k_{3}/2}Y_{\bf r}^{(k_{r})}(\omega+\epsilon)(\sqrt{3})^{k_{r}+3/2}~.$
-To achieve a significantly suppressed $\mu_{eff}$, we either need a very large negative value $k_{3}$ or a tiny $\epsilon$.
-However, calculations show the suppression from $k_{3}$ power is not very efficient for obtaining a tiny number.
+To achieve a significantly suppressed $\mu_{eff}$, we either need a very large negative value $k_{3}$ or a tiny $|\epsilon|$.
+However, calculations show that the suppression from $k_{3}$ power is not very efficient to obtain a tiny number.
 Therefore, we choose to use a small $\epsilon$ to suppress $\mu_{eff}$ with appropriate modular weights and representations for $H_{u},H_{d}$.
@@
 and the modular forms $Y_{\bf 1}^{(8)}$ (or, for example, $Y_{\bf 1}^{(4)}$ with suitable $k_{H_{u}}$ and $k_{H_{d}}$). Thus, we have
-$\mu_{eff}$ $=$ $\frac{\alpha\Lambda_{SUSY}}{|\bar{N}_{Y_{\bf 1}^{(8)}}|}\eta(\omega)^{-6}Y_{\bf 1}^{(8)}(\omega+\epsilon)(\sqrt{3})^{(k_{H_{u}}+k_{H_{d}})/2}~{},$ (77)
+$\mu_{eff}$ $=$ $\frac{\alpha\Lambda_{SUSY}}{|\bar{N}_{Y_{\bf 1}^{(8)}}|}\left[\eta(\omega)\right]^{-6}Y_{\bf 1}^{(8)}(\omega+\epsilon)(\sqrt{3})^{(k_{H_{u}}+k_{H_{d}})/2}~,$ (113)
 $\approx$ $-0.927\times 10^{3}e^{i\frac{\pi}{4}}\alpha\Lambda_{SUSY}\epsilon^{2}~.$
-Here $|\bar{N}_{Y_{\bf 1}^{(8)}}|=1.3423$ denotes the global normalization factor Petcov:2023fwh for $Y_{\bf 1}^{(8)}$, which is independent of the modulus VEV and is defined by
+Here $|\bar{N}_{Y_{\bf 1}^{(8)}}|=1.3423$ denotes the global normalization factor Petcov:2023fwh for $Y_{\bf 1}^{(8)}$, which is independent of the modulus VEV and defined by
 $(\overline{{\rm N}}^{(K)}_{\rm Y})^{2}=\int_{D}\int(N^{(K)}_{\rm Y}(\tau,\tau^{*}))^{2}\,\frac{dxdy}{y^{2}}\equiv\int_{D}\int\sum_{i}|Y^{(K)}_{{\bf r}i}(\tau)|^{2}(2\,y)^{K}\frac{dx\,dy}{y^{2}}\,$, (114)
@@
 for non-cusp modular forms, where $T$ is a real constant, $D_{T}$ is the region below the line $y=T$ in the fundamental domain with $T>1$, and $|a_{0}|^{2}$ is the constant $q$-independent term in the $q$-expansion of $\sum_{i}|Y^{(K)}_{{\bf r}i}(\tau)|^{2}$.
+It should be noted that the normalization can not be fixed in the bottom-up approach to modular symmetry.
 From the decompositions of $\Gamma_{N}$ representations under the residual symmetry group in appendix of Novichkov:2021evw , it can be verified that the suppression factor of $\epsilon$ for the $H_{u}(~{\bf 1^{\prime}},~4)$ and $H_{d}(~{\bf 1^{\prime\prime}},~4)$ bilinear is given by $({k_{H_{u}}+k_{H_{d}}+3})~{\rm modulo}~3=2$, which is consistent with the result from our explicit construction.
-For $\epsilon\sim 10^{-5}$, $\alpha\sim{\cal O}(1)$ and $\Lambda_{SUSY}$ typically of the order $\sqrt{F}\sim\sqrt{m_{3/2}M_{Pl}}\sim 10^{10}$ GeV for $m_{3/2}\sim{\cal O}({\rm 10^{2}~{}GeV})$ in gravity mediation SUSY breaking mechanism, we can obtain a desired $\mu_{eff}$ at electroweak scale.
-If $\Lambda_{SUSY}$ is naively taken to be the Planck scale, smaller $\epsilon\sim 10^{-10}$ or a tiny $\alpha$ is needed to achieve an electroweak-scale $\mu_{eff}$.
+If $\Lambda_{SUSY}$ is naively taken to be the Planck scale, a tiny $\epsilon$ with $|\epsilon|\sim 10^{-10}$ or a tiny $\alpha$ is needed to achieve an electroweak-scale $\mu_{eff}$.
+For a smaller $\Lambda_{\text{SUSY}}$ scale, a smaller suppression factor is required to achieve an effective $\mu_{\text{eff}}$ at the electroweak scale.
+For instance, in gravity mediated SUSY breaking mechanism, where the SUSY breaking scale $\Lambda_{SUSY}$ is typically of the order $\sqrt{F}\sim\sqrt{m_{3/2}M_{Pl}}\sim 10^{10}$ GeV for $m_{3/2}\sim{\cal O}({\rm 10^{2}~GeV})$, a desired $\mu_{eff}$ at electroweak scale can be regenerated for $|\epsilon|\sim 10^{-5}$ when $\alpha\sim{\cal O}(1)$.
+In gauge mediated SUSY breaking mechanism where the SUSY breaking scale is as low as $10^{5}$ GeV, the required suppression factor $\epsilon$ can be of the order $|\epsilon|\sim 10^{-2}$ for $\alpha\sim\mathcal{O}(1)$.
 Note that the smallness of deviation parameter $\epsilon$ (away from $\omega$) can be a consequence of SUSY breaking, which depends on the choice of $f_{0}$.
 For example, if the parameter choice with $f_{0}=0$ leads to the global minimum at $\omega$, SUSY breaking contributions may slightly shift the minimum to some point near $\omega$.
-We present some benchmark points that can lead to small suppression factors in Table 5.1.
-$m$ $n$ $f_{0}/M_{\mathrm{Pl}}^{2}$ $c_{1}/M_{\mathrm{Pl}}$ $\epsilon$ $\mathcal{P}(j(\tau))$ 00 $3$ $0.2$ $0.01$ $4.711\times 10^{-5}$ $1$ 00 $3$ $0.2$ $0.1$ $2.529\times 10^{-5}$ $1$ 00 $2$ $0.2$ $0.1$ $6.758\times 10^{-6}$ $1$ 00 $3$ $0.2$ $0.1$ $4.711\times 10^{-5}$ $1$ $1$ $3$ $0.2$ $0.1$ $4.711\times 10^{-5}$ $1$ 00 $3$ $0.2$ $0.1$ $2.920\times 10^{-5}$ $j-4146.6830+184.0955i$ $1$ $3$ $0.2$ $0.1$ $6.587\times 10^{-5}$ $j-4144.8675+185.3946i$ Table 1: Parameter choices for the scenario with $k_{X}=0$, which can stabilize the modulus at the point $\omega+\epsilon$ (for $|\epsilon\sim 10^{-5}|$).
+The desired condition $|\epsilon|\ll 1$ can be achieved through typical modulus stabilization mechanisms, such as the one proposed in NPP:2201.02020 , which employs large values of $m$ and $n=0$.
+However, for the modulus stabilization mechanism discussed in the previous section, fine-tuning is required to obtain a minimum at $\omega+\epsilon$ with $|\epsilon|\ll 1$.
+This is particularly true when we consider the simplest realization ${\cal P}(j(\tau))=(j(\tau)-c_{p})^{2}$ with $c_{p}\ll 1$ for $H(\tau)=0$.
+We provide in Table 2 some benchmark points that can achieve small suppression factors using the aforementioned modulus stabilization mechanism.
+It is worth noting that the fine-tuning needed here may arise from the specific choices of input parameters for our modulus stabilization mechanism.
+One could explore alternative modulus stabilization mechanisms to achieve a minimum very close to $\omega$ without large fine-tuning.
+Moreover, such fine-tuning is not necessary when employing a weighton-type mechanism for the regeneration of $\mu_{\text{eff}}$.
+$m$ $n$ $f_{0}/M_{\mathrm{Pl}}^{2}$ $c_{1}/M_{\mathrm{Pl}}$ $\epsilon/10^{-5}$ $P(j(\tau))$
+22 22 $0.2$ $0.1$ $(1.659+0.958i)$ $(j-1.765\times 10^{-9})^{2}$
+0 33 $0.2$ $0.1$ $(5.441+3.141i)$ $(j-2.861\times 10^{-9})^{2}$
+11 22 $0.2$ $0.1$ $(8.162+4.172i)$ $(j-3.973\times 10^{-8})^{2}$
+0 22 $0.2$ $0.1$ $(6.802+3.927i)$ $(j-280.469)(j-1.008\times 10^{-8})^{2}$
+22 33 $0.2$ $0.1$ $i$ $(j-1.398\times 10^{-9})^{2}$
+Table 2: Parameter choices for the scenario with $k_{X}=0$, which can stabilize the modulus at the point $\omega+\epsilon$ (for $|\epsilon|\sim 10^{-5}|$) with $X_{nl}=0$.
 5.2 Natural $\mu_{eff}$ with modulus stabilization at medium $\Im\tau$
@@
 Since the coefficient of $\epsilon^{l}$ cannot be simply determined and may involve additional suppression factors, we opt to directly calculate the suppression factor and the coefficient for the bilinear $H_{u}H_{d}$ using an explicit model construction.
-[…]
+Although the scalar potential for the modulus generally diverges as $\tau\rightarrow i\infty$, the introduction of nilpotent Goldstino $\textbf{X}_{nl}$ superfield or multiple moduli/matter fields may alter the shape of the scalar potential at intermediate values of $\tau$, making it possible to stabilize the modulus at $\tau_{0}$ with an intermediate value of $\Im\tau_{0}$ (for example, $\Im\tau_{0}\sim 5-10$).
+Detailed discussions on such a possibility will be given in our subsequent studies  [^4] For example, we can choose the form ${\cal P}(j(\tau))=(j(\tau)-j(i\Im\tau_{0}))^{2}$ (or other polynomials that contain such a factor) so as that $i\Im\tau_{0}$ is a solution of ${\cal P}(j(\tau))={\cal P}^{\prime}(j(\tau))=0$, consequently being an extrema of the scalar potential.
+It is also possible to change the shape of scalar potential near $i\infty$ by introducing non-minimal gravitational couplings of modulus fields to Ricci scalar Nonmimimal:GR .
+After transforming back from the Jordan frame to Einstein frame, the modulus field can be stabilized at intermediate values of $\tau$ more easily with the transformed scalar potential..
+In this work, we just concentrate on the implications of modulus stabilization at an intermediate value of $\Im\tau$ on the SUSY $\mu$-problem.
+Note that the stabilized VEV $\tau_{0}$ (with an intermediate value $\Im\tau_{0}$) can be regarded as lying close to the $i\infty$ fixed point.
 To obtain the suppressed effective $\mu$-parameter, in addition to possible suppression factor from powers of $q^{1/24}$ (from $\eta(\tau)$) or $(2\Im\tau)^{-1}$ (from normalization factor), we would also like to choose the modular function that vanishes at $\Im\tau=i\infty$.
-For example, the modular functions satisfy
+For example, the modular forms satisfy
 $Y_{\bf 1^{\prime}}^{(4)}(i\infty)=Y_{\bf 1^{\prime}}^{(8)}(i\infty)=Y_{\bf 1^{\prime\prime}}^{(8)}(i\infty)=0~,$ (117)
@@
 Let us suppose that the modulus field could be stabilized at the imaginary axis (within the fundamental domain) with $z_{0}=it_{0}$ for medium large $t_{0}$, when proper SUSY breaking contributions with non-vanishing $f_{0}$ are taken into account.
-Naively, if the desired suppression factor $10^{-20}$ to regenerate $\mu_{eff}$ comes solely from $\epsilon^{2}$ when $\Lambda_{SUSY}$ is naively taken to be the Planck scale, very large $t_{0}$ (here $t_{0}\gtrsim 10$) is needed.
+Naively, if the desired suppression factor to regenerate $\mu_{eff}$ comes solely from $\epsilon^{2}$ when $\Lambda_{SUSY}$ is naively taken to be the Planck scale, very large $t_{0}$ (here $t_{0}\gtrsim 10$) is needed.
 Such a large $t_{0}$ possibly causes difficulties for modulus stabilization realizations, as the scalar potential for modulus field generally diverges for $\tau\rightarrow i\infty$.
 Therefore, it is desirable to introduce additional suppression factors other than $\epsilon$ so that medium large $t_{0}$ can lead to desired suppression factors.
-The effective $\mu$-term from the bilinear term for $H_{u}H_{d}$ in (74), which satisfies $k_{H_{u}}+k_{H_{d}}=k_{r}-k_{3}+3$, is given by
-$\mu_{eff}$ $=$ $\alpha\Lambda_{SUSY}\eta(it_{0})^{-2k_{3}}Y_{\bf r}^{(k_{r})}(it_{0})(2t_{0})^{(k_{H_{u}}+k_{H_{d}})/2}~{},$ (84)
+The effective $\mu$-term from the bilinear term for $H_{u}H_{d}$ in (110), which satisfies $k_{H_{u}}+k_{H_{d}}=k_{r}-k_{3}+3$ for $2k_{3}\in 24\mathbb{Z}+6$, is given by
+$\mu_{eff}$ $=$ $\alpha\Lambda_{SUSY}\left[\eta(it_{0})\right]^{-2k_{3}}Y_{\bf r}^{(k_{r})}(it_{0})(2t_{0})^{(k_{H_{u}}+k_{H_{d}})/2}~,$ (120)
 $\approx$ $\alpha\Lambda_{SUSY}e^{\frac{k_{3}\pi t_{0}}{6}}(2t_{0})^{-\frac{k_{3}}{2}}Y_{\bf r}^{(k_{r})}(it_{0})(2t_{0})^{\frac{k_{r}+3}{2}}~,$
-where proper normalization factors for $H_{u}$,$H_{d}$ from the Kahler potential are taken into account.
+where proper normalization factors for $H_{u}$$H_{d}$ from the Kahler potential are taken into account.
 For fixed $k_{r}$, we require that the factor
@@
 where ${\bar{N}_{Y_{\bf 1^{\prime\prime}}^{(8)}}}=20.4152$ is the global normalization factor for the modular form $Y_{\bf 1^{\prime\prime}}^{(8)}$.
-Calculations show that the value $\mu_{eff}\lesssim 10^{-20}M_{SUSY}$ can be achieved for $t_{0}\gtrsim 5.58$, if $\Lambda_{SUSY}$ is naively taken to be the Planck scale.
-For $\Lambda_{SUSY}$ typically of the order $\sqrt{F}\sim\sqrt{m_{3/2}M_{Pl}}\sim 10^{10}$ GeV, it is sufficient to choose $t_{0}\gtrsim 3.2$ to satisfy $\mu_{eff}\lesssim 10^{-8}M_{SUSY}$.
-The tiny value of $\mu_{eff}$ results from the combined effect of suppression from powers of $q$ (or $(2\Im\tau)^{-1}$) and the asymptotic suppression behavior of typical modular forms away from the fixed point $i\infty$.
+Calculations show that the value $\mu_{eff}\lesssim 10^{-17}\Lambda_{SUSY}$ can be achieved for $t_{0}\gtrsim 5.01$, if $\Lambda_{SUSY}$ is naively taken to be the Planck scale.
+For $\Lambda_{SUSY}$ typically of the order $\sqrt{F}\sim\sqrt{m_{3/2}M_{Pl}}\sim 10^{10}$ GeV, it is sufficient to choose $t_{0}\gtrsim 3.2$ to satisfy $\mu_{eff}\lesssim 10^{-8}\Lambda_{SUSY}$.
+For a smaller $\Lambda_{SUSY}$ scale, a smaller suppression factor is needed to achieve an effective $\mu_{eff}$ at the electroweak scale, requiring a smaller value of $t_{0}$.
+The tiny value of $\mu_{eff}$ results from the combined effect of suppression by powers of $q$ (or $(2\Im\tau)^{-1}$) and the asymptotic suppression behavior of typical modular forms away from the fixed point $i\infty$.
 It can also be verified that the suppression power of $\epsilon=\exp[-\frac{2\pi}{3}\Im z_{0}]$ for $H_{u}(~{\bf 1},~16)$ and $H_{d}(~{\bf 1^{\prime}},~16)$ bilinear is given by the power factor $l$ of $(\rho_{i}^{c}\rho_{j})^{*}=\zeta^{l}$ for $\zeta=\exp(2\pi i/3)$, which can be read out from Novichkov:2021evw to be $(1\cdot\zeta)^{*}=\zeta^{2}$ and predict $l=2$.
@@
 Since naturalness requires $\mu_{eff}$ to lie at the electroweak scale or soft SUSY breaking scale, a large suppression factor leading to $\mu_{eff}\ll\Lambda_{SUSY}$ is required when the modulus field is stabilized.
-Such a large suppression factor can naturally arise from a modulus VEV in the vicinity of certain fixed points.
+Such a large suppression factor can naturally arise from a modulus VEV in the vicinity of certain fixed points or by the suppression factor in weighton-like mechanism.
 Ordinary single modulus stabilization mechanism can be affected by SUSY breaking contributions.
@@
 We discuss the stabilization mechanism of a single modulus field in the presence of SUSY breaking contributions, described by the non-linear SUSY realization scheme with a nilpotent Goldstino $\textbf{X}_{nl}$.
-The presence of $\textbf{X}_{nl}$ contributions can slightly alter the original stabilized modulus VEV to a value near the fixed points.
 Natural value of $\mu_{eff}$, which is significantly suppressed relative to the SUSY scale, can be achieved from the expansion of typical modular forms using a tiny deviation parameter near the fixed points.
@@
 The modulus stabilization mechanism with nilpotent $\textbf{X}_{nl}$ superfield can be fairly generic and ubiquitous.
-The regeneration of small $\mu_{eff}$ term following modulus stabilization, which just needs a modulus VEV near some fixed points, is independent on the concrete modulus stabilization mechanism.
+The regeneration of small $\mu_{eff}$ term following modulus stabilization, which just needs a modulus VEV near some fixed points or through a weighton-like mechanism for $H_{u}H_{d}$ bilinear, is independent of the concrete modulus stabilization mechanism.
 Appendix A Properties of Modular $A_{4}$
@@
 $Y_{3}(\tau)=-18q^{2/3}(1+2q+5q^{2}+4q^{3}+8q^{4}+\dots)\,.$ (131)
+Appendix B Eisenstein Series
+The Eisenstein series $G_{k}(\tau)$, defined by
+[…]
+is a modular form of weight $k$ for $\tau\in{\cal H}$ and has the following Fourier decomposition
+$G_{k}(\tau)=2\zeta(k)+2\frac{(2\pi i)^{k}}{\Gamma(k)}\sum\limits_{N=1}^{\infty}\sigma_{k-1}(N)q^{N}~,$ (133)
+with $\sigma_{\alpha}(N)$ the sums of divisors functions, defined for any $\alpha\in\mathcal{C}$ by
+$\sigma_{\alpha}(N)=\sum\limits_{l|N}l^{\alpha}~.$ (134)
+It can be proven that $G_{4}(\rho)=G_{6}(i)=0$.
+Factoring out $2\zeta(k)$, and using the relation between Bernoulli numbers $B_{k}$ and $\zeta(k)$ for even $k$, the normalized Eisenstein series $E_{k}$ can be defined by
+$E_{k}(\tau)=\frac{G_{k}(\tau)}{2\zeta(k)}=1-\frac{2k}{B_{k}}\sum\limits_{N=1}^{\infty}\sigma_{k-1}(N)q^{N}~,$ (135)
+with some of the $\nu_{k}\equiv-{2k}/{B_{k}}$ values given by
+$\nu_{4}=240,~\nu_{6}=504,~\nu_{8}=480.$ (136)
+The quasi-modular transformation property of $E_{2}(\tau)$
+$E_{2}(\gamma\tau)$ $=$ $(c\tau+d)^{2}E_{2}(\tau)+\frac{12}{2\pi i}c(c\tau+d)~,$ (137)
+allows one to evaluate $E_{2}(\tau)$ at the points $\tau=i$ and $\tau=\omega$.
+The values of $E_{k}(\tau)$ are listed below DHoker:2022dxx
+$E_{2}(i)$ $=$ $\frac{3}{\pi},~~E_{2}(\omega)=\frac{2\sqrt{3}}{\pi},~~E_{4}(i)=\frac{48[\Gamma(\frac{5}{4})]^{4}}{\pi^{2}[\Gamma(\frac{3}{4})]^{4}},$~
+$E_{4}(\omega)$ $=$ $0,~~E_{6}(i)=0,~~E_{6}(\omega)=\frac{729}{2\pi^{3}}\frac{[\Gamma(\frac{4}{3})]^{6}}{[\Gamma(\frac{5}{6})]^{6}}.$ (138)
+The second derivative of $j(\tau)$ are given by
+$j^{\prime\prime}(\tau)=\frac{4\pi^{2}}{[\eta(\tau)]^{24}}\left\{\frac{1}{6}E_{2}(\tau)[E_{4}(\tau)]^{2}E_{6}(\tau)-\frac{2}{3}E_{4}(\tau)[E_{6}(\tau)]^{2}-\frac{1}{2}[E_{4}(\tau)]^{4}\right\}~,$ (139)
+from Ramanujan’s formulas for derivatives of Eisenstein series, which gives
+$j^{\prime\prime}(\omega)=0~,~~j^{\prime\prime}(i)=-\frac{2\pi^{2}}{[\eta(i)]^{24}}[E_{4}(i)]^{4}~.$ (140)
+Appendix C Solutions of $H^{\prime}(\tau)=0$
+Such a constraint reduces to
+$H^{\prime}(\tau)=\left\{j(\tau)^{n/3}[j(\tau)-1728]^{m/2}\mathcal{P}(j(\tau))\right\}^{\prime}=0,$ (141)
+or equivalently,
+$H(\tau)j^{\prime}(\tau)\left[\frac{n}{3}\frac{1}{j(\tau)}+\frac{m}{2}\frac{1}{j(\tau)-1728}+\frac{1}{\mathcal{P}(j(\tau))}\frac{\partial\mathcal{P}(j(\tau))}{\partial j(\tau)}\right]=0.$ (142)
+Note that the expression
+$H(\tau)\left[\frac{n}{3}\frac{1}{j(\tau)}+\frac{m}{2}\frac{1}{j(\tau)-1728}+\frac{1}{\mathcal{P}(j(\tau))}\frac{\partial\mathcal{P}(j(\tau))}{\partial j(\tau)}\right],$ (143)
+can be singular at certain solutions of $H(\tau)=0$.
+Such singular behavior would be eliminated by the multiplication of $j^{\prime}(\tau)$ to render $H^{\prime}(\tau)$ finite at such points, which can be seen in (45).
+Firstly, we would like to discuss the case $H(\tau)\neq 0$. From equation (143), it is obvious that either
+$\left[\frac{n}{3}\frac{1}{j(\tau)}+\frac{m}{2}\frac{1}{j(\tau)-1728}+\frac{1}{\mathcal{P}(j(\tau))}\frac{\partial\mathcal{P}(j(\tau))}{\partial j(\tau)}\right]=0~,$ (144)
+or $j^{\prime}(\tau)=0$.
+When $H(\tau)\neq 0$, for non-negative integers $m$ and $n$ (by the definition of $H(\tau)$), we have the following discussions for equation (144) after eliminating the $H(\tau)$ factor:
+  • $m+n>0$:
+    • When $n\neq 0$ and $m=0$, condition (144) reduces to
+      $\frac{d\mathcal{P}(j(\tau))}{\mathcal{P}(j(\tau))}=-\frac{n}{3}\frac{dj(\tau)}{j(\tau)},$ (145)
+      which gives
+      $\mathcal{P}(j(\tau))=C_{0}[j(\tau)]^{-n/3}.$ (146)
+      So, $\mathcal{P}(j(\tau))$ is not a polynomial of $j(\tau)$ unless $n\in 3\mathbb{Z}$ is negative, contradicting the assumption that $n$ is non-negative integers.
+    • When $m\neq 0$ and $n=0$, it can also be deduced that $\mathcal{P}(j(\tau))$ is a polynomial of $j(\tau)$ unless $m\in 2\mathbb{Z}$ is negative, contradicting the assumption that $m$ is non-negative integers.
+    • For $m\neq 0$ and $n\neq 0$, similar deduction gives
+      $\mathcal{P}(j(\tau))=C_{0}^{\prime}[j(\tau)]^{-n/3}[j(\tau)-1728]^{-m/2}.$ (147)
+      So, $\mathcal{P}(j(\tau))$ is not a polynomial of $j(\tau)$ unless both $m\in 2\mathbb{Z}$ and $n\in 3\mathbb{Z}$ are negative.
+As $m,n$ are defined to be non-negative integers, this condition requires $m=n=0$, contradicting the assumptions that $m,n\neq 0$.
+    Therefore, $m+n>0$ can not hold.
+  • $m=0$ and $n=0$:
+    With the constraints $m=n=0$, the condition (143) reduces to
+    $\frac{\partial\mathcal{P}(j(\tau))}{\partial j(\tau)}=0,$ (148)
+    which gives
+    $\mathcal{P}(j(\tau))=\text{Constant}.$ (149)
+    Since $H(\tau)\neq 0$, this constant must therefore be non-zero.
+When $H(\tau)\neq 0$, for the condition $j^{\prime}(\tau)=0$, the derivative of the Klein $j$-function is given by
+$\frac{\partial}{\partial\tau}j(\tau)=-2\pi i\frac{{E_{6}(\tau)}{E_{4}^{2}(\tau)}}{[\eta(\tau)]^{24}}=0~.$ (150)
+It can checked that its solutions are $\omega$ and $i$, contradicting the assumption $H(\tau)\neq 0$ with $m=n=0$.
+Therefore, the condition $H^{\prime}(\tau)=0$ with $H(\tau)\neq 0$ implies that $H(\tau)$ is a non-zero constant.
+The solutions of $H^{\prime}(\tau)=0$ for the case $H(\tau)=0$ are addressed in the paragraphs containing equations (70), (71), and (72).
 Acknowledgements.
-This work was supported by the National Natural Science Foundation of China under grant numbers Nos.12075213 and 12335005, by the Natural Science Foundation for Distinguished Young Scholars of Henan Province under grant number 242300421046.
+These authors contributed equally: Hong Jie Fan, Ying Kai Zhang.
+We are very grateful to the referees for helpful discussions.
+This work was supported by the National Natural Science Foundation of China under grant numbers Nos.12075213 and 12335005, by the Natural Science Foundation of Henan (Distinguished Young Scholars of Henan Province) under grant number 242300421046.